### v2

# Dynamics of Agentic Loops in Large Language Models: A Geometric Theory of Trajectories From Semantic Contraction to Exploratory Divergence

Abstract
Agentic systems based on large language models operate through recursive feedback loops, where each output becomes the next input in an iterative transformation process. Understanding the geometric dynamics of these agentic loops, whether they converge, diverge or exhibit more complex behaviors, remains poorly characterized. This paper develops a geometric framework for analyzing agentic loop trajectories in semantic embedding spaces, treating iterative transformations as discrete dynamical systems.
We formalize the distinction between the artifact space (where linguistic transformations occur) and the semantic embedding space (where geometric measurements are performed). Since standard cosine similarities exhibit systematic bias due to embedding anisotropy, we introduce calibrated similarity using isotonic regression, which eliminates bias and improves correlation with human semantic judgments while maintaining high local stability. This calibrated metric enables precise geometric analysis through trajectories (temporal sequences of embeddings), clusters (regions of semantic coherence) and attractors (convergence points).
Experimentally, we characterize two fundamental dynamical regimes through singular agentic loops. A contractive loop performing iterative paraphrasing exhibits geometric convergence toward a stable attractor with decreasing dispersion. An exploratory loop alternating summarization and negation shows unbounded divergence with no stable clustering structure. These regimes demonstrate qualitatively distinct trajectory geometries, contraction versus expansion in semantic space.
Prompt design directly controls the dynamical regime, enabling systematic engineering of convergent versus divergent behaviors. This geometric framework provides the first rigorous quantitative approach to analyzing iterative transformation dynamics in LLMs, with applications to trajectory prediction, stability analysis and controlled composition of complex multi-step transformations.
1 Introduction
Large language models (LLMs) enable autonomous agentic systems that operate through recursive feedback loops, where each model output becomes the next input in an iterative transformation process. These agentic loops appear in diverse applications: iterative refinement systems that progressively improve outputs through self-critique, multi-step reasoning chains that decompose complex problems into intermediate steps and autonomous agents that maintain extended interactions with environments or users. Understanding the dynamics of these loops, whether they converge toward stable states, diverge unboundedly or exhibit complex oscillatory behaviors, remains poorly characterized. Despite their increasing deployment in reasoning, planning and generation tasks, we lack a rigorous framework for analyzing how iterative transformations behave geometrically in semantic space.
Characterizing agentic loop dynamics has direct implications across multiple domains. For system design, knowing whether a loop exhibits convergent or divergent behavior enables engineers to select appropriate prompts and parameters for specific tasks, stable convergence is desirable for refinement and summarization, while controlled divergence may be useful for exploration and generation. For reliability and safety, detecting convergence properties helps identify runaway feedback loops that could produce unbounded drift or oscillation. For theoretical understanding, geometric trajectory analysis provides a quantitative lens for studying how language model transformations behave as discrete dynamical systems. The ability to predict, measure and control these dynamics is fundamental to deploying reliable iterative generative systems.
Contributions.
The main contributions of this work are:
  1. 1.
    A geometric framework for analyzing agentic loops. We distinguish the artifact space (where text transformations occur) from the semantic embedding space (where geometric measurements are performed) and formalize how transformations induce measurable trajectories, temporal sequences of embeddings that trace the loop’s evolution through semantic space.
  2. 2.
    A calibrated similarity metric. To overcome the fundamental obstacle that pretrained embeddings exhibit anisotropy, concentrating cosine similarities around uninformative high values, we develop a calibrated metric using isotonic regression on human semantic judgments. This calibration eliminates bias and improves correlation while maintaining high local stability (98%), enabling precise trajectory analysis.
  3. 3.
    The first systematic empirical characterization of dynamical regimes. Through singular agentic loops, we identify and characterize two fundamental regimes: contractive convergence and exploratory divergence. We demonstrate that prompt design directly controls the geometric behavior of iterative transformations, establishing that agentic loop dynamics are predictable and controllable through prompt engineering.
The remainder of the paper is organized as follows. Section 2 reviews related work on iterative refinement, agentic systems, dynamical systems and embedding geometry. Sections 35 construct the geometric measurement infrastructure: embedding space construction, similarity calibration and stability validation. Section 6 formalizes the theoretical framework, defining agentic loops, trajectories, clusters, attractors and dynamical regimes. Section 7 presents experimental trajectory studies of contractive and exploratory singular loops, revealing fundamentally distinct geometric behaviors. Section 8 discusses implications for composite loops and controlled dynamics. Section 9 concludes.
An earlier version of this work was released as an arXiv preprint.
2 Related Work
Computational creativity and the measurement problem.
The question of implementing computational creativity in iterative systems provides the broader motivation for this work. Boden boden2009computermodels established the foundational framework for computational creativity, identifying three primary processes: exploration (searching conceptual spaces for new combinations), transformation (modifying the rules or structure of conceptual spaces) and combination (combining ideas from different domains). Colton and Wiggins colton2012computational formalized creativity through measurable dimensions: novelty (semantic or structural distance from existing examples), value (quality, coherence and usefulness) and surprise (unexpectedness relative to prior expectations). Zhao et al. zhao2024creativity recently demonstrated that large language models show creative behaviors across multiple tasks and domains.
These frameworks raise a fundamental question for iterative LLM systems: How can we engineer agentic loops to implement Boden’s creative processes (explore-transform-combine) in ways that elevate Colton & Wiggins’ creativity metrics (novelty, value, surprise)? Answering this question requires the ability to measure similarity-based metrics along trajectories, specifically to quantify semantic distance (for novelty), coherence (for value) and discontinuities (for surprise) as iterative transformations unfold over time. However, computational creativity theory provides conceptual frameworks without addressing the foundational measurement infrastructure: how do we rigorously measure semantic similarity and trajectory dynamics in embedding spaces?
This measurement problem motivates our focus on geometric trajectory analysis as a necessary first step. Before we can engineer creative agentic loops or evaluate creativity metrics, we must establish rigorous methods for measuring similarity and characterizing trajectory dynamics. Our work addresses this foundational layer, building the geometric and dynamical measurement infrastructure required to eventually tackle computational creativity questions systematically.
Research on iterative generative systems has explored how large language models behave under recursive transformations, but geometric trajectory analysis of these dynamics remains underdeveloped. Beyond the creativity motivation, this work builds upon three main research threads: iterative refinement in language models, agentic systems and multi-step reasoning and embedding geometry.
Iterative refinement in language models.
Madaan et al. madaan2023selfrefine introduced Self-Refine, demonstrating that large language models can iteratively critique and improve their own outputs through a feedback loop without additional training. Their approach achieved substantial improvements across diverse tasks including dialogue generation and mathematical reasoning. However, Self-Refine focused on empirical task performance rather than analyzing the geometric structure of how representations evolve through iterations or characterizing convergence properties in semantic space.
Wei et al. wei2022chain showed that chain-of-thought prompting enables large language models to perform complex reasoning by decomposing problems into intermediate steps. While this work established that prompting strategies significantly affect reasoning capabilities, it did not examine the stability or trajectory dynamics of iterative processes.
Agentic systems and multi-step reasoning.
Recent work has explored how language models can act as autonomous agents through iterative interaction with environments or tools. Yao et al. yao2022react proposed ReAct, which interleaves reasoning traces with task-specific actions, enabling models to interact with external knowledge sources. Shinn et al. shinn2023reflexion introduced Reflexion, where agents use verbal reinforcement learning to maintain episodic memory and improve decision-making across trials. Park et al. park2023generative demonstrated generative agents that simulate believable human behavior through architectures combining memory, reflection and planning.
These agentic frameworks establish that recursive LLM interactions can produce sophisticated behaviors, but they do not formally characterize the dynamics of such systems in semantic space or analyze convergence and divergence patterns as geometric phenomena.
Embedding geometry and semantic similarity.
Understanding the geometric structure of embedding spaces is crucial for analyzing semantic trajectories. Ethayarajh ethayarajh2019contextual systematically studied contextualized representations from BERT, ELMo and GPT-2, demonstrating that embeddings become increasingly anisotropic in upper layers, with vectors concentrating in a narrow cone. This anisotropy causes high cosine similarities even between semantically distant items, distorting similarity-based measurements.
Wang and Isola wang2020understanding analyzed contrastive representation learning through the lens of alignment (bringing similar items together) and uniformity (distributing representations across the hypersphere). They showed that good representations balance these competing objectives, providing theoretical grounding for why pretrained embeddings exhibit systematic geometric biases.
For calibrating similarity metrics, Niculescu-Mizil and Caruana niculescu2005predicting demonstrated that isotonic regression effectively transforms classifier scores into well-calibrated probabilities. While originally developed for classification tasks, isotonic regression provides a principled non-parametric method for aligning model outputs with empirical distributions. We extend this approach to semantic similarity calibration.
Positioning of the present work.
In contrast to existing work, which primarily focuses on static properties of embeddings or isolated transformations, the present work studies the dynamical behavior of iterative LLM transformations as geometric trajectories. Prior research on iterative refinement madaan2023selfrefine ; wei2022chain and agentic systems yao2022react ; shinn2023reflexion ; park2023generative has demonstrated impressive capabilities but treats each iteration as an independent step rather than as part of a continuous dynamical process. Similarly, embedding geometry research ethayarajh2019contextual ; wang2020understanding characterizes static distributional properties without examining how representations evolve under repeated transformations. Our work bridges these threads by treating agentic loops as discrete dynamical systems in semantic space, enabling the characterization of convergence, divergence and attractor structures through calibrated geometric measurements.
Contribution of this work.
While computational creativity frameworks boden2009computermodels ; colton2012computational motivate the need to measure trajectory-based metrics, prior work has not developed the foundational geometric measurement infrastructure required to rigorously analyze iterative transformation dynamics. Self-refinement approaches madaan2023selfrefine demonstrate empirical task improvements but do not characterize how representations evolve geometrically through iterations or analyze convergence versus divergence behaviors. Agentic frameworks yao2022react ; shinn2023reflexion ; park2023generative establish recursive LLM capabilities but lack quantitative measurement of trajectory dynamics in semantic space. Although embedding anisotropy is well-documented ethayarajh2019contextual and theoretical frameworks for representation quality exist wang2020understanding , no existing work provides calibrated, human-aligned similarity metrics for rigorous trajectory analysis in iterative transformations.
Our contribution builds the foundational geometric measurement infrastructure required for analyzing agentic loop dynamics, providing the tools necessary to eventually address computational creativity questions systematically. We address these gaps through three main innovations. First, we develop a geometric framework operating in the representation space treating iterative transformations as discrete dynamical systems. This framework captures measurable trajectories, clusters, attractors and dynamical regimes that quantify how semantic representations evolve under recursive application. These geometric constructs open the path to provide to measure Colton & Wiggins’ creativity criteria. Second, we introduce isotonic calibration that aligns embedding similarities with human semantic judgments extending calibration methods niculescu2005predicting to overcome anisotropy ethayarajh2019contextual and enable precise geometric measurements using metrics that reflect human-perceived semantic distances. Third, we provide systematic empirical characterization of dynamical regimes. We demonstrate that prompt design directly determines whether agentic loops exhibit convergence, divergence or complex behaviors, as revealed by measurable geometric indicators in calibrated semantic space. This establishes the foundation for engineering Boden’s creative processes through controlled trajectory dynamics.
3 Semantic Space Construction
3.1 Artifact Space
Before defining how we measure semantic dynamics, we must first specify what is being measured. In an agentic loop, the fundamental objects of transformation are artifacts: outputs generated and processed by a large language model. While artifacts can take various modalities (text, images, audio, code, etc.), in this paper we focus exclusively on textual artifacts.
We define the artifact space $\mathcal{A}$ as the set of all possible text strings that can be generated by the language model. Each element $a\in\mathcal{A}$ represents a complete textual output.
While $\mathcal{A}$ is finite (bounded by the model’s vocabulary and maximum context length), it is combinatorially large and lacks inherent geometric structure suitable for quantitative analysis of convergence, divergence or semantic drift. To measure such dynamics, we construct a continuous representation space through embedding, as described in the following subsection.
3.2 Representation Space and Embedding Function
To enable quantitative analysis of semantic dynamics, we construct a continuous representation space $\mathcal{E}$ by embedding textual artifacts into a geometric space where distances and similarities can be measured.
Let $\phi:\mathcal{A}\to\mathbb{R}^{d}$ denote an embedding function that maps each artifact $a\in\mathcal{A}$ to a dense vector representation $\phi(a)\in\mathbb{R}^{d}$ in a $d$-dimensional Euclidean space. To ensure that similarity measurements depend solely on directional alignment (not magnitude), we apply $\ell_{2}$-normalization to all embeddings. Let $\psi:\mathcal{A}\to\mathbb{S}^{d-1}$ denote the normalized embedding function defined as:
$\psi(a)=\frac{\phi(a)}{\|\phi(a)\|_{2}},$ (1)
where $\mathbb{S}^{d-1}=\{e\in\mathbb{R}^{d}:\|e\|_{2}=1\}$ is the (d1)(d-1)-dimensional unit hypersphere. We define the representation space as:
$\mathcal{E}=\left\{\psi(a)\mid a\in\mathcal{A}\right\}\subset\mathbb{S}^{d-1}.$ (2)
Each artifact $a$ is thus represented by a unit vector $e=\psi(a)\in\mathcal{E}$.
Model-agnostic framework.
The framework presented above is embedding-model agnostic: any sentence encoder that maps text to dense vectors can instantiate $\phi$ and our analysis of trajectories, clusters and dynamical regimes applies regardless of the specific model choice.
Implementation choice.
For this work, we instantiate $\phi$ using the Xenova/paraphrase-mpnet-base-v2 model, a pretrained transformer encoder with $d=768$ dimensions. This model was chosen for three practical reasons:
  1. 1.
    It is trained on sentence pairs with human-labeled paraphrase relations, ensuring that geometric proximity corresponds to semantic similarity.
  2. 2.
    It offers a balanced trade-off between expressiveness and dimensionality, allowing fine-grained distinctions while remaining computationally efficient.
  3. 3.
    The Xenova implementation provides a reproducible and locally executable inference pipeline, facilitating controlled experimentation without external API dependencies.
The normalized representation space $\mathcal{E}\subset\mathbb{S}^{767}$ thus forms the manifold on which all further measurements (distances, similarities and clustering) are defined.
3.3 Similarity Metric
Within the representation space $\mathcal{E}$, each artifact $a\in\mathcal{A}$ is represented by its normalized embedding $e=\psi(a)\in\mathcal{E}\subset\mathbb{S}^{d-1}$. We define the model-based similarity function as
$s^{(m)}:\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\to[-1,1],\qquad s^{(m)}(e_{1},e_{2})=\langle e_{1},e_{2}\rangle.$ (3)
Since all elements of $\mathcal{E}$ are unit vectors (i.e., $\|e\|_{2}=1$ for all $e\in\mathcal{E}$), the inner product $\langle e_{1},e_{2}\rangle$ directly equals the cosine of the angle between them, providing a normalized measure of semantic proximity.
In the following subsections, we first characterize the theoretical distribution of similarities in high-dimensional spaces (§3.4), then evaluate how well $s^{(m)}$ aligns with human semantic judgments through quantitative metrics (§3.5) and visualization (§3.7).
3.4 High-Dimensional Similarity Distribution
On a standard high-dimensional unit sphere $\mathbb{S}^{d-1}$, two random vectors $u,v$ drawn uniformly from the surface tend to be nearly orthogonal as the dimension increases. Formally, let $X$ be uniformly distributed on $\mathbb{S}^{d-1}$ and let $f(X)=\langle X,v\rangle$ for a fixed $v\in\mathbb{S}^{d-1}$. By Lévy’s concentration inequality (see vershynin2018high for a modern treatment), for any $t>0$,
$\mathbb{P}\left(\left|f(X)-\mathbb{E}[f(X)]\right|>t\right)\leq 2e^{-cnt^{2}},$ (4)
where $c>0$ is an absolute constant and n=d1n=d-1 is the dimension of the sphere. Since $\mathbb{E}[f(X)]=0$ for the uniform distribution, the inner product $\langle u,v\rangle$ concentrates exponentially fast around 0 as $d\to\infty$. In such isotropic settings, the similarity distribution is symmetric and sharply peaked near zero.
However, the semantic embedding space induced by a pretrained language model such as paraphrase-mpnet-base-v2 diverges from this theoretical isotropy. Because the model was trained on non-uniform pairs of semantically related sentences, the embeddings are not uniformly distributed over the hypersphere but rather cluster around a dominant mean direction in $\mathbb{R}^{d}$. This induces a strong anisotropy of the embedding manifold, where most vectors share a large positive cosine with the mean embedding vector. As formalized by wang2020understanding , good representation learning requires balancing two properties: alignment (similar points should be close) and uniformity (embeddings should be distributed uniformly on the unit sphere). The observed concentration violates the uniformity property, indicating that the pretrained model prioritizes alignment over uniform distribution. This trade-off is typical of contrastively trained embeddings. Empirically, the distribution of $s^{(m)}(e_{1},e_{2})$ is concentrated around $0.8$, in stark contrast with the isotropic baseline centered at 0.
3.5 Quantitative Metrics for Model-Human Alignment
To quantitatively assess how well $s^{(m)}$ aligns with human semantic judgments, we use the MTEB STS Benchmark (STS-train) dataset. This dataset contains pairs of textual artifacts $(a_{1},a_{2})\in\mathcal{A}\times\mathcal{A}$ annotated with normalized human similarity scores $s^{(h)}(a_{1},a_{2})\in[0,1]$, rescaled from the original 0–5 range.
For each pair $(a_{1},a_{2})$, we compute their normalized embeddings $e_{1}=\psi(a_{1})$ and $e_{2}=\psi(a_{2})$ in the representation space $\mathcal{E}$. We then evaluate the model-based similarity $s^{(m)}(e_{1},e_{2})$ and compare it to the human judgment $s^{(h)}(a_{1},a_{2})$.
We evaluate their alignment using complementary metrics, each capturing a distinct aspect of correlation or calibration:
  • Root Mean Squared Error (RMSE):
    $\mathrm{RMSE}=\sqrt{\frac{1}{N}\sum_{k=1}^{N}\big(s^{(m)}(e_{1}^{k},e_{2}^{k})-s^{(h)}(e_{1}^{k},e_{2}^{k})\big)^{2}}\in[0,1].$
    Measures the average magnitude of deviation between model and human similarities, penalizing large discrepancies more strongly. Lower RMSE indicates better agreement (RMSE = 0 means perfect match; RMSE = 1 indicates maximum deviation).
  • Mean Bias Error (MBE):
    $\mathrm{MBE}=\frac{1}{N}\sum_{k=1}^{N}\big(s^{(m)}(e_{1}^{k},e_{2}^{k})-s^{(h)}(e_{1}^{k},e_{2}^{k})\big)\in[-1,1].$
    Quantifies systematic bias: positive when the model tends to overestimate human similarity, negative when it underestimates. MBE = 0 indicates no systematic bias; $|\mathrm{MBE}|$ close to 1 indicates severe consistent over- or under-estimation.
  • Expected Calibration Error (ECE): This metric evaluates the degree to which model similarity scores are statistically calibrated to human perception. We partition the similarity range $[0,1]$ into $B$ bins, where $B_{b}$ denotes the set of sample pairs in bin $b$. For each bin, we compute:
    • $\mathrm{acc}(B_{b})$: the average human similarity score (accuracy, i.e., ground truth) within that bin
    • $\mathrm{conf}(B_{b})$: the average model similarity score (confidence, i.e., predicted value) within that bin
    The ECE is then computed as the weighted average absolute deviation:
    $\mathrm{ECE}=\sum_{b=1}^{B}\frac{|B_{b}|}{N}\,\big|\,\mathrm{acc}(B_{b})-\mathrm{conf}(B_{b})\,\big|\in[0,1].$
    Lower ECE values indicate better calibration (ECE = 0 means perfect calibration; ECE = 1 indicates maximum miscalibration).
  • Pearson Correlation Coefficient ($r$):
    $r=\frac{\mathrm{Cov}(s^{(m)},s^{(h)})}{\sigma_{(m)}\,\sigma_{(h)}}\in[-1,1],$
    where $\mathrm{Cov}(s^{(m)},s^{(h)})$ is the covariance between model and human similarity scores, $\sigma_{(m)}$ is the standard deviation of model scores and $\sigma_{(h)}$ is the standard deviation of human scores. Explicitly, this can be computed as:
    $r=\frac{\sum_{k=1}^{N}(s^{(m)}_{k}-\bar{s}^{(m)})(s^{(h)}_{k}-\bar{s}^{(h)})}{\sqrt{\sum_{k=1}^{N}(s^{(m)}_{k}-\bar{s}^{(m)})^{2}}\;\sqrt{\sum_{k=1}^{N}(s^{(h)}_{k}-\bar{s}^{(h)})^{2}}},$
    where $\bar{s}^{(m)}$ is the mean of model scores, $\bar{s}^{(h)}$ is the mean of human scores and $N$ is the total number of pairs. This metric captures the strength of linear dependence between model and human similarities. $r=1$ indicates perfect positive linear correlation, r=1r=-1 indicates perfect negative linear correlation and $r=0$ indicates no linear correlation. Values closer to $\pm 1$ indicate stronger linear alignment.
  • Spearman Rank Correlation ($\rho$):
    $\rho=r_{\text{Pearson}}(\mathrm{rank}(s^{(m)}),\mathrm{rank}(s^{(h)}))\in[-1,1],$
    where $\mathrm{rank}(s^{(m)})$ assigns each model similarity score its position when all $N$ scores are sorted in ascending order and similarly for $\mathrm{rank}(s^{(h)})$. The function $r_{\text{Pearson}}(\cdot,\cdot)$ denotes the Pearson correlation coefficient applied to these rank vectors. This metric measures the monotonic alignment between the two orderings, independently of scale. $\rho=1$ indicates perfect monotonic agreement (rankings perfectly aligned), $\rho=-1$ indicates perfect monotonic disagreement (rankings perfectly reversed) and $\rho=0$ indicates no monotonic relationship. Values closer to $\pm 1$ indicate stronger monotonic alignment.
Quantitatively, the results obtained on the STS-train dataset are:
$\mathrm{RMSE}=0.1702,\quad\mathrm{MBE}=0.0789,\quad\mathrm{ECE}=0.0797,\quad r=0.8576,\quad\rho=0.8430.$ (5)
These metrics reveal several key properties of the embedding space. The RMSE of 0.17 indicates a moderate average deviation between model and human similarities, suggesting reasonable overall agreement with some residual discrepancies. The positive MBE of 0.079 shows a systematic tendency for the model to overestimate semantic similarity compared to human judgments. This is a characteristic feature of dense embedding spaces where related concepts cluster more tightly than human perception suggests. The ECE of 0.080 quantifies the calibration gap: the model’s similarity scores are not perfectly aligned with the empirical frequency of human agreement, reflecting the need for recalibration to achieve probabilistic consistency. The strong correlation values ($r=0.86$ and $\rho=0.84$) confirm that despite these biases, the model captures the underlying rank order and linear structure of human semantic similarity judgments. Together, these results indicate that the embedding space preserves semantic relationships effectively while exhibiting systematic geometric biases, specifically anisotropic concentration of embeddings, that lead to consistent overestimation and miscalibration.
3.6 High-Confidence Similarity Threshold
To obtain a reliable decision boundary for semantic similarity, we introduce the High-Confidence Similarity Threshold $\tau_{\text{HCS}}(s)$, a data-driven value derived from human-judged highly similar pairs. This threshold provides a statistically principled criterion for determining when two sentences can be considered semantically similar with high confidence. Importantly, this concept is independent of any particular similarity measure and can be applied to any similarity function.
Definition. For a given similarity function $s$ (such as the raw cosine similarity $s^{(m)}$ or any other similarity function), the threshold is computed as:
$\tau_{\text{HCS}}(s)=Q_{0.05}(s\mid s^{(h)}>0.9)$ (6)
that is, the 5% quantile of the similarity scores $s$ among sentence pairs with high human similarity ($s^{(h)}>0.9$). The confidence level $\alpha=0.05$ is a standard choice, though it can be adjusted based on application requirements.
Interpretation. This definition is meaningful for three key reasons. First, the threshold is grounded in human judgments of semantic similarity rather than arbitrary choice. Second, it captures the lower bound of similarity scores for similar sentences as judged by humans. Third, it provides a direct probabilistic guarantee: at least 95% of truly similar pairs (those with $s^{(h)}>0.9$) lie above this threshold.
Statistical justification. By construction, the threshold satisfies:
$\mathbb{P}(s\geq\tau_{\text{HCS}}(s)\mid s^{(h)}>0.9)\geq 0.95$ (7)
This conditional probability allows us to make reliable decisions: if a pair exceeds the threshold, it is likely to belong to the high-similarity group as defined by human raters.
Decision criterion. Any pair with similarity $s\geq\tau_{\text{HCS}}(s)$ can be considered semantically similar with high confidence.
For the raw cosine similarity, we denote:
$\tau_{\text{HCS}}^{(m)}=\tau_{\text{HCS}}(s^{(m)})\approx 0.72$ (8)
3.7 Visualization of Similarity Distributions
Beyond these numerical indicators, we visualize the distributions and relationships between $s^{(m)}$ and $s^{(h)}$ using density estimation techniques.
To compare model similarities with human similarity judgments, we estimate their respective continuous distributions using Kernel Density Estimation (KDE). Each observed value contributes a Gaussian “bump” to the overall density, producing a smooth, non-parametric approximation of the underlying probability function:
$\hat{f}_{h}(x)=\frac{1}{nh\sqrt{2\pi}}\sum_{i=1}^{n}e^{-\frac{1}{2}\left(\frac{x-x_{i}}{h}\right)^{2}},\qquad h=1.06\,\hat{\sigma}\,n^{-1/5}.$ (9)
Silverman’s rule silverman1986density selects the optimal bandwidth $h$, balancing smoothness and resolution. We compute separate densities for human scores and model similarities, $\hat{f}_{H}(x)$ and $\hat{f}_{S}(x)$ and plot them over $[0,1]$. Alignment between both curves indicates good agreement, while systematic shifts reveal under- or over-confidence in the model’s similarity predictions. Additionally, both visualizations include a vertical green line marking the high-confidence similarity threshold $\tau_{\text{HCS}}^{(m)}$ (defined in §3.6), which serves as a data-driven decision boundary for determining when two sentences can be considered semantically similar with high confidence.
In addition to that, we also compute joint density heatmaps. To evaluate the relationship between human similarity judgments and model predictions, we estimate their joint density over the $0,1]^{2}$ space. The domain is divided into a regular grid of bins and each cell’s density $\rho_{ij}=N_{ij}/(n\,\Delta x\,\Delta y)$ reflects the concentration of pairs $(h_{i},s_{i})$. A 2D Gaussian smoothing kernel can be applied to obtain a continuous density field $\tilde{\rho}(x,y)$, from which iso-density contours and a color intensity map are derived. The diagonal $y=x$ represents perfect alignment, where model similarities equal human scores. High density concentrated along this diagonal indicates strong agreement, while systematic deviations above or below it reveal over- or under-confidence and vertical dispersion signals inconsistency in the model’s predictions.
(a) Density plot (KDE)
(b) Heatmap (binned)
Figure 1: Original (uncalibrated) cosine similarity versus human similarity scores.
Figure 1(a) reveals a striking concentration of model similarities around $0.8$, contrasting sharply with the broader distribution of human scores across $[0,1]$. This concentration reflects the anisotropic geometry of the embedding space, where semantically related sentences cluster tightly on the unit hypersphere, violating the theoretical isotropy assumption. The green vertical line at $\tau_{\text{HCS}}^{(m)}$ marks the decision boundary derived from high-confidence human judgments, positioned within this concentration region, indicating that most model similarities for genuinely similar pairs fall within this narrow band. Human judgments, by contrast, exhibit greater variability, with substantial density in the low-to-moderate similarity range ($[0.2,0.6]$), indicating more nuanced perceptual distinctions.
Figure 1(b) visualizes the joint density $\tilde{\rho}(s^{(h)},s^{(m)})$ as a 2D heatmap. The density mass lies predominantly above the diagonal $y=x$, confirming systematic overestimation: the model assigns higher similarity scores than humans across most pairs. The vertical spread for human scores in the range $[0.3,0.7]$ indicates high variance in model predictions for pairs with intermediate human similarity: the model struggles to distinguish subtle semantic differences, compressing them into a narrow band around $0.8$. In contrast, pairs with very low human similarity ($s^{(h)}<0.3$) show tighter correspondence, suggesting the model correctly identifies dissimilar pairs but struggles to capture gradations of relatedness.
These visualizations expose a fundamental calibration gap: the model’s cosine similarity $s^{(m)}(e_{1},e_{2})$ does not align with human semantic judgments $s^{(h)}(e_{1},e_{2})$, both in terms of distribution shape (concentration vs. spread) and pointwise correspondence (systematic overestimation). This raises a natural question: Can we recalibrate or transform the model’s similarity function to improve its correlation with human semantic judgments? This question is investigated in the following section.
4 Calibration of Semantic Similarity
The visualizations in Section 3.7 reveal systematic bias in the model’s similarity predictions: $s^{(m)}$ consistently overestimates human judgments $s^{(h)}$ and exhibits compressed variance. Calibration provides a post-processing mechanism to correct these distributional and pointwise misalignments.
Formally, calibration seeks a transformation function $g:[-1,1]\to\mathbb{R}$ such that the calibrated similarity
$\tilde{s}(e_{1},e_{2})=g\big(s^{(m)}(e_{1},e_{2})\big)$ (10)
better aligns with human similarity judgments $s^{(h)}(e_{1},e_{2})$.
4.1 Overview of Calibration Methods
We evaluate several calibration functions that map cosine similarity to human-aligned scores, each varying in parametric flexibility and nonlinearity.
Linear regression fits a simple affine transformation $\tilde{s}(x)=ax+b$, providing a global scaling and shift but unable to capture nonlinear patterns (Figure 2).
Isotonic regression learns a monotonic, piecewise-constant mapping that adapts locally to the data without assuming a functional form, allowing flexible nonlinear calibration while preserving rank ordering (Figure 3).
Sigmoid mapping applies a logistic curve $\tilde{s}(x)=1/(1+e^{-a(x-b)})$, modeling smooth saturation at similarity extremes (Figure 4).
Polynomial transformations (degrees 2, 3 and 4) fit $\tilde{s}(x)=\sum_{i=0}^{n}a_{i}x^{i}$, offering increasing flexibility to model curvature, though higher degrees may introduce unwanted oscillations (Figures 5, 6, 7).
Beta distribution calibration uses a beta CDF parameterized by shape $\alpha$ and $\beta$, suited for bounded, skewed distributions but prone to instability when assumptions mismatch the data (Figure 8).
(a) Density plot (KDE)
(b) Heatmap (binned)
Figure 2: Linear regression calibration.
(a) Density plot (KDE)
(b) Heatmap (binned)
Figure 3: Isotonic regression calibration.
(a) Density plot (KDE)
(b) Heatmap (binned)
Figure 4: Sigmoid calibration.
(a) Density plot (KDE)
(b) Heatmap (binned)
Figure 5: Polynomial degree 2 calibration.
(a) Density plot (KDE)
(b) Heatmap (binned)
Figure 6: Polynomial degree 3 calibration.
(a) Density plot (KDE)
(b) Heatmap (binned)
Figure 7: Polynomial degree 4 calibration.
(a) Density plot (KDE)
(b) Heatmap (binned)
Figure 8: Beta distribution calibration.
4.2 Results
Table 1 presents the performance comparison of all calibration methods against the original (uncalibrated) cosine similarity using correlation metrics. Percentage changes relative to the original baseline are shown for reference.
Method RMSE MBE ECE Pearson $r$ Spearman $\rho$
Original 0.1702 0.0789 0.0797 0.8576 0.8430
Linear 0.1506 $\downarrow$11.5% 0.0000 $\downarrow$100.0% 0.0222 $\downarrow$72.1% 0.8576 $\pm$0.0% 0.8430 $\pm$0.0%
Isotonic 0.1411 $\downarrow$17.1% 0.0000 $\downarrow$100.0% 0.0000 $\downarrow$99.9% 0.8764 $\uparrow$2.2% 0.8563 $\uparrow$1.6%
Sigmoid 0.3192 $\uparrow$87.6% 0.2604 $\uparrow$230.2% 0.2604 $\uparrow$226.7% 0.8182 $\downarrow$4.6% 0.8430 $\pm$0.0%
Beta 0.5168 $\uparrow$203.6% 0.3011-0.3011 $\uparrow$281.8% 0.4243 $\uparrow$432.3% 0.8266-0.8266 $\downarrow$196.4% 0.8246-0.8246 $\downarrow$197.8%
Polynomial-2 0.1474 $\downarrow$13.4% 0.0004 $\downarrow$99.5% 0.0095 $\downarrow$88.1% 0.8641 $\uparrow$0.8% 0.8430 $\pm$0.0%
Polynomial-3 0.1472 $\downarrow$13.5% 0.0000 $\downarrow$100.0% 0.0057 $\downarrow$92.8% 0.8644 $\uparrow$0.8% 0.8430 $\pm$0.0%
Polynomial-4 0.1472 $\downarrow$13.5% 0.0000-0.0000 $\downarrow$100.0% 0.0058 $\downarrow$92.7% 0.8645 $\uparrow$0.8% 0.8430 $\pm$0.0%
Table 1: Performance comparison of calibration methods with percentage changes relative to the original baseline. For RMSE, MBE and ECE, lower is better (green $\downarrow$); for Pearson and Spearman, higher is better (green $\uparrow$).
Isotonic regression achieves the strongest monotonic correlation (Spearman $\rho=0.8563$) and lowest calibration error (ECE =0.0000=0.0000), with perfect elimination of mean bias (MBE =0.0000=0.0000) and a 17.1% reduction in RMSE compared to the original baseline. This defines our calibrated similarity function
$\tilde{s}:\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\to\mathbb{R},\qquad\tilde{s}=f_{\mathrm{isotonic}}\circ s^{(m)}$ (11)
used throughout the dynamic analysis. In practice, since we align with human similarity scores in $[0,1]$, calibrated outputs are clamped to this interval.
Similarly, we define the high-confidence similarity threshold for the calibrated similarity:
$\tilde{\tau}_{\text{HCS}}=\tau_{\text{HCS}}(\tilde{s})\approx 0.65$ (12)
5 Evaluating Local Stability of the Semantic Space
5.1 Motivation
Even if the calibrated semantic similarity aligns well with human judgments, it remains essential to evaluate the local stability of the embedding function $\psi:\mathcal{A}\to\mathcal{E}$ itself. A locally stable embedding function should preserve high similarity under small linguistic perturbations: when two texts differ only slightly in form or syntax, their semantic similarity should remain consistently high. However, $\psi$ is learned and discretized; neither its continuity nor its Lipschitz properties can be guaranteed theoretically. Our first objective is therefore empirical: to determine whether $\psi$ behaves as locally stable in practice.
However, a critical methodological concern arises from our use of isotonic regression for calibration. While isotonic regression optimally preserves monotonicity and eliminates bias, it produces a piecewise constant (and therefore discontinuous) mapping $f_{\text{isotonic}}$. This raises an essential question: does the discontinuous nature of isotonic regression degrade the local stability observed in the raw embedding space?
To address both concerns, we conduct a dual analysis of stability:
  1. 1.
    Base model stability, evaluated on raw cosine similarity $s^{(m)}$, measures how the uncalibrated embedding function reacts to small linguistic perturbations.
  2. 2.
    Calibrated stability, evaluated on the calibrated similarity $\tilde{s}$, measures whether isotonic calibration preserves or degrades this local stability.
Comparing these two perspectives allows us to validate that calibration does not introduce artificial instabilities, ensuring that the semantic trajectories observed in agentic loops reflect genuine semantic dynamics rather than artifacts of the calibration process.
5.2 Artifact Perturbation Definition
To evaluate local stability, we constructed a dedicated dataset using ChatGPT to generate pairs of sentences differing by minimal and controlled linguistic perturbations. Each pair $(a_{i},a_{i}^{\prime})$ was manually categorized into one of seven perturbation types, each intended to simulate a semantically minimal transformation while introducing specific syntactic or lexical variations. This approach allows for systematic testing of $\psi$’s local stability across controlled linguistic dimensions.
The following perturbation types were included, each illustrated with examples from the dataset:
  • DETERMINER_VARIATION: A minimal change affecting only the determiner (the, a, this, that) without modifying the tense, structure or semantic content of the sentence. Example: “The cat sleeps on the sofa.” $\to$ “A cat sleeps on the sofa.”
  • TENSE_VARIATION: A transformation that shifts the verbal tense while preserving the event described by the sentence. Example: “She drinks tea in the morning.” $\to$ “She drank tea in the morning.”
  • SYNONYM_SUBSTITUTION: Replacing a lexical item (noun, verb, adjective) with a close synonym that preserves the meaning. Example: “He reads a book every night.” $\to$ “He reads a novel every night.”
  • LOGICAL_PARAPHRASE: A reformulation that preserves the logical meaning of the sentence, often through active/passive alternation or structural inversion. Example: “The teacher explains the lesson.” $\to$ “The lesson is explained by the teacher.”
  • NOMINALIZATION: Transforming a predicate into its nominal form (or the reverse) while keeping the meaning equivalent. Example: “He decided to leave early.” $\to$ “He made a decision to leave early.”
  • COREFERENCE_EXPANSION: Replacing a pronoun with its explicit referent (or vice versa), keeping the propositional content unchanged. Example: “Alice lost her keys.” $\to$ “Alice lost Alice’s keys.”
  • QUANTIFIER_VARIATION: Replacing a quantifier with another logically equivalent form (all $\leftrightarrow$ every), without changing the truth conditions. Example: “All cats like to sleep.” $\to$ “Every cat likes to sleep.”
The dataset comprises 700 pairs across all categories (100 pairs per type), covering both lexical and syntactic variations. This controlled diversity allows us to examine how different perturbation types affect similarity in both the raw and calibrated embedding spaces.
5.3 Metrics and Methodology
For each pair $(a_{i},a_{i}^{\prime})$, we compute their normalized embeddings $e_{i}=\psi(a_{i})$ and $e_{i}^{\prime}=\psi(a_{i}^{\prime})$ within $\mathbb{S}^{d-1}$. We then measure stability using two complementary similarity functions:
Base model similarity.
We compute the raw cosine similarity:
$s_{i}^{(m)}=s^{(m)}(e_{i},e_{i}^{\prime})=\langle e_{i},e_{i}^{\prime}\rangle,$ (13)
which measures the uncalibrated similarity between perturbed pairs.
Calibrated similarity.
We also compute the calibrated similarity:
$\tilde{s}_{i}=\tilde{s}(e_{i},e_{i}^{\prime})=f_{\text{isotonic}}(s_{i}^{(m)}),$ (14)
where $f_{\text{isotonic}}$ denotes the isotonic regression mapping introduced in Section 4.
For each perturbation type $k$ with $N_{k}$ pairs, we compute three stability indicators:
$\mu_{k}=\frac{1}{N_{k}}\sum_{i\in\mathcal{I}_{k}}s_{i},\quad\sigma_{k}=\sqrt{\frac{1}{N_{k}}\sum_{i\in\mathcal{I}_{k}}(s_{i}-\mu_{k})^{2}},\quad r_{k}=\frac{1}{N_{k}}\sum_{i\in\mathcal{I}_{k}}\mathbf{1}_{\{s_{i}\geq\tau\}},$ (15)
where $\mathcal{I}_{k}$ denotes the index set of pairs for perturbation type $k$ and $\tau$ is the stability threshold.
Threshold selection.
Rather than choosing arbitrary thresholds, we use the data-driven high-confidence similarity thresholds defined in Section 3.6:
  • For base model similarity: $\tau=\tau_{\text{HCS}}^{(m)}\approx 0.72$
  • For calibrated similarity: $\tau=\tilde{\tau}_{\text{HCS}}\approx 0.65$
The stability rate $r_{k}$ thus measures the proportion of perturbed pairs that maintain similarity above the threshold derived from high-confidence human judgments, providing a principled criterion for local stability.
5.4 Empirical Results
Tables 2 and 3 report the local stability metrics for base model (raw cosine) and calibrated similarities, respectively. For each perturbation type, we provide the number of pairs, mean similarity, standard deviation and stability rate (proportion of pairs exceeding the corresponding threshold).
Perturbation Type Pairs ($N_{i}$) Mean ($\mu_{\text{base},i}$) Std. Dev. ($\sigma_{\text{base},i}$) Stab. Rate ($r_{\text{base},i}$)
DETERMINER VARIATION 100 0.932 0.017 1.00
TENSE VARIATION 100 0.909 0.030 1.00
SYNONYM SUBSTITUTION 100 0.853 0.060 0.95
LOGICAL PARAPHRASE 100 0.891 0.023 1.00
NOMINALIZATION 100 0.866 0.035 1.00
COREFERENCE EXPANSION 100 0.882 0.022 1.00
QUANTIFIER VARIATION 100 0.848 0.039 0.97
All types (overall) 700 0.883 0.045 0.99
Table 2: Local stability metrics using base model similarity $s^{(m)}$ (threshold $\tau_{\text{HCS}}^{(m)}=0.72$).
Perturbation Type Pairs ($N_{i}$) Mean ($\mu_{\text{sim},i}$) Std. Dev. ($\sigma_{\text{sim},i}$) Stab. Rate ($r_{\text{sim},i}$)
DETERMINER VARIATION 100 0.926 0.054 1.00
TENSE VARIATION 100 0.868 0.069 1.00
SYNONYM SUBSTITUTION 100 0.799 0.082 0.87
LOGICAL PARAPHRASE 100 0.831 0.049 1.00
NOMINALIZATION 100 0.818 0.078 1.00
COREFERENCE EXPANSION 100 0.830 0.045 1.00
QUANTIFIER VARIATION 100 0.798 0.050 0.97
All types (overall) 700 0.839 0.075 0.98
Table 3: Local stability metrics using calibrated similarity $\tilde{s}$ (threshold $\tilde{\tau}_{\text{HCS}}=0.65$).
5.5 Discussion
The comparison between Tables 2 and 3 reveals two critical findings regarding local stability.
Embedding function exhibits strong local stability.
The base model similarity analysis (Table 2) demonstrates that the embedding function $\psi$ exhibits excellent local stability, with an overall stability rate of $99\%$. All perturbation types maintain mean similarities well above the threshold $\tau_{\text{HCS}}^{(m)}=0.72$, ranging from $0.848$ (QUANTIFIER_VARIATION) to $0.932$ (DETERMINER_VARIATION). This confirms that the pretrained embedding model is robust to small linguistic perturbations, preserving high semantic similarity even under controlled transformations such as tense variation, synonym substitution and logical paraphrasing.
Isotonic calibration introduces minimal stability degradation.
Table 3 shows that isotonic regression calibration slightly reduces the overall stability rate from $99\%$ to $98\%$, a modest degradation of only 1 percentage point. This small decrease is attributable to the piecewise constant (discontinuous) nature of isotonic regression: when a perturbed pair $(a_{i},a_{i}^{\prime})$ has a raw cosine similarity $s^{(m)}$ near a discontinuity point of $f_{\text{isotonic}}$, the calibrated similarities $\tilde{s}(e_{i},e_{i})$ and $\tilde{s}(e_{i}^{\prime},e_{i}^{\prime})$ may fall into different constant segments, potentially crossing the threshold $\tilde{\tau}_{\text{HCS}}=0.65$.
The most affected perturbation type is SYNONYM_SUBSTITUTION, whose stability rate drops from $0.95$ to $0.87$. This is expected: synonym substitutions produce moderate cosine similarities (mean $0.853$) that fall in a region where isotonic regression has more step discontinuities. In this moderate range, small linguistic perturbations can cause similarities to jump across step boundaries in $f_{\text{isotonic}}$, leading to larger changes in calibrated values than in the original cosine similarities. Conversely, perturbation types with very high base similarities (TENSE_VARIATION: $0.909$, DETERMINER_VARIATION: $0.932$) remain perfectly stable ($r=1.00$) after calibration, as their similarities fall in a region where isotonic regression is nearly constant and far from the threshold.
Practical acceptability of calibrated stability.
Despite the theoretical discontinuity of isotonic regression, the empirical stability rate of $98\%$ remains highly acceptable for practical use. This demonstrates that isotonic regression achieves its primary goal (eliminating bias and improving correlation with human judgments) without substantially compromising local stability. The 1% degradation is a negligible cost compared to the significant gains in calibration quality (ECE reduction from $0.0797$ to $0.0000$, as shown in Table 1).
Implications for dynamic analysis.
These results validate the use of calibrated similarity $\tilde{s}$ for studying agentic loop dynamics. The embedding function $\psi$ exhibits strong local stability and isotonic calibration preserves this property with only minimal degradation. This ensures that the semantic trajectories observed in agentic loops reflect genuine semantic evolution rather than instabilities introduced by either the embedding function or the calibration process. With local stability empirically established, we now turn to the formal analysis of agentic loop dynamics, confident that our similarity metric provides a stable and human-aligned foundation for measuring iterative transformations in generative systems.
6 Agentic Loop Dynamics: Theoretical Framework
This section formalizes the theoretical framework underlying agentic loop dynamics, defining key concepts such as trajectory, cluster and attractor before turning to empirical experiments.
6.1 Foundational Spaces and Operations
We begin by establishing the mathematical spaces and primitive operations that underlie all agentic loop structures.
Artifact and embedding spaces.
Let $\mathcal{A}$ denote the artifact space, where each element $a\in\mathcal{A}$ represents a linguistic artifact (e.g., a sentence, paragraph or document). Let $\psi:\mathcal{A}\to\mathcal{E}$ be the embedding map introduced in Section 3, which projects artifacts into a semantic vector space where their meaning can be quantitatively compared.
Primitive operations.
We introduce two fundamental operations:
  • $\mathrm{LLM}:\mathcal{A}\to\mathcal{A}$ denotes a large language model’s generation operation, which takes a formatted input string (prompt) and produces a generated output string. The same underlying model may be invoked with different generation parameters (temperature $T$, top-p, top-k, seed, etc.), which we denote as $\mathrm{LLM}_{\theta}$ where $\theta$ represents the parameter configuration.
  • $P:\mathcal{A}\to\mathcal{A}$ denotes a prompt template, a function that embeds an artifact $a\in\mathcal{A}$ into a formatted prompt string suitable for LLM consumption. For example, $P(a)=\text{``Rewrite the following: \{$a$\}''}$ structures the artifact within an instruction.
6.2 Agentic Loops and Trajectories
Agentic Loop.
An agentic loop is a discrete dynamical system $(\mathcal{A},F)$ where $F:\mathcal{A}\to\mathcal{A}$ is a transformation operator that defines the recursive rule generating the sequence:
$a_{t+1}=F(a_{t}),\quad t\in\mathbb{N},\quad\text{with initial condition }a_{0}\in\mathcal{A}.$ (16)
Using the primitive operations defined above, $F$ is implemented as:
$F(a_{t})=\mathrm{LLM}(P(a_{t})).$ (17)
Each iteration produces a new artifact $a_{t+1}$ from the previous artifact $a_{t}$ through this composition of prompt template and LLM generation, starting from an initial artifact $a_{0}$. The corresponding representations in the embedding space $\mathcal{E}$ evolve as:
$e_{t}=\psi(a_{t}),\qquad e_{t+1}=\psi(a_{t+1})=\psi(\mathrm{LLM}(P(a_{t}))),\quad\text{with }e_{0}=\psi(a_{0}).$ (18)
The resulting sequence $\{e_{t}\}_{t=0}^{T}$ allows us to track the evolution of the loop dynamics in the continuous embedding space $\mathcal{E}$, where we can quantitatively measure distances and similarities between successive states.
We distinguish three types of agentic loops based on the structure of $F$:
  1. 1.
    Singular loop: $F=\mathrm{LLM}\circ P$ where $P$ is a single prompt template. Each iteration applies one LLM call: $a_{t+1}=\mathrm{LLM}(P(a_{t}))$. This is the focus of our empirical study (Section 7).
  2. 2.
    Composite loop: $F=\mathrm{LLM}_{n}\circ P_{n}\circ\cdots\circ\mathrm{LLM}_{1}\circ P_{1}$ where each $(P_{i},\mathrm{LLM}_{i})$ pair represents a distinct transformation phase. Here, $\mathrm{LLM}_{i}$ denotes an LLM invocation with specific generation parameters (temperature, top-p, seed, etc.), which may differ across phases to induce different behaviors, for instance high temperature for creative exploration, low temperature for precise consolidation. Different model architectures can also be used (e.g., specialized models for critique vs. generation). For example, generate-critique-revise requires three sequential LLM calls with different prompt templates and potentially different sampling configurations. Composite loops enable controlled creativity through role-based phase alternation (Section 8.4).
  3. 3.
    Graph based loop: $F$ is derived from a directed acyclic computation graph that could include parallel execution and conditional branches, forming the basis for multi-agent architectures (Section 8.5).
This paper establishes the theoretical framework applicable to all three types, then focuses empirically on singular loops to rigorously characterize fundamental dynamical regimes. Understanding singular loop behavior provides the necessary foundation for analyzing composite and graph based loops in future work.
Agentic Trajectory.
The iterative process produces an ordered sequence of embedded representations:
$\mathcal{T}=\{e_{t}=\psi(a_{t})\mid t=0,1,\dots,T\}\subset\mathcal{E},$ (19)
which we call the agentic trajectory. Each transition $(e_{t},e_{t+1})$ represents a local semantic displacement, whose magnitude and direction quantify the instantaneous dynamics of the loop.
6.3 Semantic Geometry of Trajectories
Having defined how trajectories are generated, we now introduce geometric tools to measure their structure. When a trajectory exhibits temporal stability, meaning embeddings remain close to each other for an extended period, we need mathematical concepts to characterize this clustering behavior.
Center of Gravity.
Given a set of embeddings $E=\{e_{i_{1}},e_{i_{2}},\ldots,e_{i_{k}}\}\subset\mathcal{E}$, the center of gravity $a_{E}$ is defined as the $L^{2}$-normalized mean:
$a_{E}=\frac{\mu_{E}}{\|\mu_{E}\|_{2}},\quad\text{where}\quad\mu_{E}=\frac{1}{k}\sum_{j=1}^{k}e_{i_{j}}.$ (20)
The center of gravity minimizes the sum of squared Euclidean distances to all embeddings in $E$ and is itself a unit vector: $\|a_{E}\|_{2}=1$. This provides a natural reference point for measuring how tightly a set of embeddings is concentrated.
Semantic Dispersion.
To quantify this concentration, we define the semantic dispersion, which measures the maximum deviation from the center in calibrated similarity space:
$\text{Dispersion}(E)=\max_{e\in E}\big[1-\tilde{s}(e,a_{E})\big],$ (21)
where $\tilde{s}$ is the calibrated similarity function from Section 4. Smaller dispersion indicates tighter semantic coherence; $\text{Dispersion}(E)=0$ means all embeddings are semantically identical to the center of gravity. Together, the center of gravity and dispersion allow us to characterize when a trajectory remains in a stable semantic region.
6.4 Clusters and Attractors
Armed with these geometric tools, we can now formalize the notion of a cluster, a contiguous temporal segment where the trajectory maintains semantic stability.
Cluster.
A cluster $C\subseteq\mathcal{T}$ is a maximal contiguous temporal window $t_{a},t_{b}]$ of the agentic trajectory where embeddings maintain semantic coherence, up to a set of outliers $O_{C}$:
$C=\{e_{t}\mid t\in[t_{a},t_{b}]\setminus O_{C}\}\quad\text{with center of gravity}\quad a_{C}=\frac{\sum_{e\in C}e}{\|\sum_{e\in C}e\|_{2}}.$ (22)
The cluster represents a phase during which the loop operates within a bounded semantic region, repeatedly generating artifacts with similar meanings.
Cluster Validity Constraints.
Not every temporal window qualifies as a cluster. To ensure semantic coherence, a cluster must satisfy three validity constraints:
  1. 1.
    Similarity ($\lambda$): For all $t\in[t_{a}+1,t_{b}]\setminus O_{C}$, $\tilde{s}(e_{t-1},e_{t})\geq\lambda$, ensuring consecutive embeddings remain semantically close
  2. 2.
    Dispersion ($\rho$): $\text{Dispersion}(C)<\rho$, ensuring the cluster as a whole stays tightly concentrated
  3. 3.
    Patience ($\kappa$): At most $\kappa$ consecutive violations allowed before cluster termination, allowing temporary noise
Parameters: $\lambda\in[0,1]$ (local coherence), $\rho>0$ (global coherence), $\kappa\in\mathbb{N}$ (noise tolerance). These constraints maintain semantic coherence while allowing for occasional outliers.
Attractor.
For a cluster $C$ with center of gravity $a_{C}$, the attractor is simply $a_{C}$, the semantic region toward which the trajectory converges during $C$’s temporal extent. This terminology borrows from dynamical systems theory: the attractor represents the "pull" that keeps the trajectory within a stable semantic basin.
6.5 Dynamic Regimes
Finally, we characterize the global behavior of agentic loops by categorizing the qualitative patterns their trajectories exhibit over time.
Classification.
An agentic loop’s dynamics fall into one of three regimes:
  • Contractive ($\mathcal{R}_{\mathrm{ctr}}$): There exists a time $T$ such that for all $t\geq T$, the trajectory remains within a single cluster. The loop stabilizes around a coherent semantic region.
  • Oscillatory ($\mathcal{R}_{\mathrm{osc}}$): The trajectory alternates among a finite set of recurrent clusters without converging to one. The loop cycles through distinct semantic regions.
  • Exploratory ($\mathcal{R}_{\mathrm{exp}}$): The trajectory exhibits unbounded or aperiodic movement within $\mathcal{E}$, showing no stable attractor. The loop continuously generates novel semantic content without stabilization.
These categories provide a conceptual basis for describing the emergent behavior of agentic loops and for distinguishing structured evolution from uncontrolled semantic drift.
Scope and operationalization.
The formal definitions provided above serve as conceptual anchors for the mathematical framework. In our empirical analysis (Section 7), we use operational approximations rather than computing exact topological limits. This pragmatic approach allows us to identify and characterize emergent structures within finite-sample trajectories while remaining grounded in the theoretical principles outlined above.
Note that the Oscillatory Regime (periodic cycling among multiple attractors) is not observed in our experiments (Section 7). The prompts we designed exhibit either full convergence (Contractive) or complete divergence (Exploratory), with no sustained oscillation between distinct semantic regions. Exploring conditions that produce oscillatory dynamics, such as prompts with explicit alternating instructions or adversarial constraints, remains an avenue for future work.
6.6 Cluster Detection Algorithm
To operationalize the cluster and attractor definitions presented in Section 6.4, we employ an incremental cluster detection algorithm that processes the trajectory $\{e_{t}\}_{t=0}^{T}$ sequentially, identifying stable regions where embeddings maintain semantic coherence around a common center of gravity. The algorithm implements the validity constraints defined earlier using three parameters: similarity threshold $\lambda\in[0,1]$, dispersion threshold $\rho>0$ and patience $\kappa\in\mathbb{N}$.
Detection procedure.
The algorithm processes embeddings sequentially and maintains a candidate cluster. At each iteration $t$, it evaluates the cluster validity constraints:
  1. (i)
    Similarity: consecutive embeddings satisfy $\tilde{s}(e_{t-1},e_{t})\geq\lambda$;
  2. (ii)
    Dispersion: adding $e_{t}$ preserves $\text{Dispersion}(C\cup\{e_{t}\})<\rho$.
If both constraints are satisfied, $e_{t}$ is added to the current cluster. If either is violated, a violation counter is incremented. When the violation counter exceeds $\kappa$, the cluster is closed and its center of gravity is computed. An iteration $t$ is classified as an outlier if it occurs during an active cluster and violates at least one constraint, but consecutive violations have not yet exceeded the patience threshold $\kappa$. Outliers are excluded from the cluster’s center of gravity computation, but the cluster remains open until more than $\kappa$ consecutive violations occur.
This incremental procedure detects multiple clusters along a trajectory, each characterized by its center of gravity (attractor), temporal extent and semantic dispersion.
6.7 Cluster Trajectory Visualization
To visualize cluster membership and semantic stability along an agentic trajectory, we construct a two-dimensional representation that encodes cluster identity, attractor similarity and outlier status simultaneously.
Graph construction.
For a trajectory with $m$ detected clusters $\{C_{1},\ldots,C_{m}\}$, each cluster $C_{i}$ is assigned a baseline at $y=i$. For each iteration $t$ belonging to cluster $C_{i}$ (i.e., $t\in[t_{a}^{i},t_{b}^{i}]$), we plot a point at coordinates:
$(x,y)=\bigl(t,\,i+\alpha\cdot d(t,C_{i})\bigr),\quad\text{where}\quad d(t,C_{i})=1-\tilde{s}(e_{t},a_{i}).$ (23)
Here, $a_{i}$ is the attractor of cluster $C_{i}$ and $\tilde{s}$ is the calibrated similarity. The vertical offset $d(t,C_{i})$ measures how far embedding $e_{t}$ deviates from the cluster’s attractor in similarity space. A scaling factor $\alpha>0$ amplifies vertical displacements for better visualization of small deviations.
Points are colored by membership status:
  • Blue points: valid cluster members ($t\in[t_{a}^{i},t_{b}^{i}]\setminus O_{i}$), contributing to attractor computation
  • Red points: outliers ($t\in O_{i}$), excluded from attractor but within patience threshold $\kappa$
A grey band of height $\alpha\cdot\rho$ around each baseline $y=i$ visualizes the dispersion constraint: all blue points must remain below $y=i+\alpha\cdot\rho$ for the cluster to satisfy the dispersion constraint.
Reading the graph.
The vertical position encodes semantic proximity to the attractor:
  • A point at $y=i$ indicates $\tilde{s}(e_{t},a_{i})=1$ (embedding identical to attractor)
  • A point at $y=i+\alpha\cdot 0.1$ indicates $\tilde{s}(e_{t},a_{i})=0.9$ (high similarity)
  • Maximum vertical spread of blue points equals $\alpha\cdot\text{Dispersion}(C_{i})$
Tight clusters exhibit blue points clustered near the baseline, while dispersed clusters show wider vertical spread. Red outliers reveal iterations where the patience mechanism allows temporary violations without terminating the cluster. Gaps between clusters indicate transition periods where the trajectory does not belong to any detected cluster.
7 Experimental Study of Singular Agentic Loops
Having established the theoretical framework for analyzing agentic loop dynamics, we now turn to empirical validation. As discussed in Section 6.2, singular loops represent the simplest agentic structure: a single transformation $F$ applied recursively. By studying singular loops first, we can isolate fundamental dynamical phenomena (contraction, exploration, attractor formation) without the additional complexity introduced by multi-step compositions. This foundational understanding is essential: composite loops, which alternate between generative and critical phases and graph based loops, which coordinate multiple interacting agents, both build upon the basic dynamics characterized here. The experiments presented in this section demonstrate that even the simplest agentic loops exhibit rich, regime-dependent behavior, validating the geometric framework and establishing empirical benchmarks for future studies of more complex architectures.
7.1 Experimental Design
To empirically characterize the dynamics of agentic loops, we conduct a controlled experiment using two minimal singular loops executed under identical computational conditions. Both operate in the artifact space $\mathcal{A}$ as purely textual transformations applied iteratively through a single LLM call per step. The objective is to observe whether distinct prompt formulations can induce fundamentally different dynamic regimes, specifically one exhibiting contractive behavior and another exhibiting exploratory divergence.
LLM configuration.
All experiments are performed locally using the Ollama inference framework (version 0.12.9) with the deepseek-r1:8b model. This model provides a lightweight yet expressive transformer architecture, suitable for studying fine-grained textual transformations over multiple recursive steps. We use a sampling temperature of $0.8$, which introduces mild stochasticity while preserving semantic coherence, ensuring that the observed trajectories reflect both the model’s internal tendencies and the intrinsic stability of the prompt formulation. No explicit random seed was specified, so each generation uses one determined by the inference framework, contributing to the stochastic nature of the observed dynamics.
Motivation.
The purpose of this experiment is not to model complex reasoning or task-solving behavior, but rather to reveal that even the simplest iterative linguistic transformations can produce structured dynamics when observed in the calibrated semantic space (,s~)($\mathcal{E}$$\tilde${s}). By isolating the loop mechanism to a single prompt applied recursively, we eliminate confounding factors such as multi-step planning or external memory accumulation. This enables direct observation of the intrinsic dynamical tendencies of the language model itself.
Experimental setup.
Each loop is defined by a fixed prompt template $P$, into which the current artifact $a_{t}$ is inserted at each iteration:
$a_{t+1}=F(a_{t})=\mathrm{LLM}(P(a_{t})).$ (24)
The process is repeated for $T=50$ iterations, generating a trajectory $\{a_{t}\}_{t=0}^{T}$ and its projection $\{e_{t}=\psi(a_{t})\}_{t=0}^{T}$ in the representation space $\mathcal{E}$. All trajectories are embedded using the same function $\psi$ and analyzed in the calibrated similarity space defined in Section 4.
Initial condition.
Both loops are initialized with the same starting sentence:
"Music has the power to connect people across cultures and generations."
This neutral, semantically coherent sentence serves as the common baseline $a_{0}$ for comparing the contractive and exploratory dynamics.
Cluster detection configurations.
To identify emergent structures such as clusters or attractors within each trajectory, we apply the incremental cluster detection algorithm described in Section 6.6. We explore multiple parameter configurations $(\lambda,\rho,\kappa)$ to assess the robustness of attractor detection under varying spatial resolutions:
[…]
This grid allows both fine-grained and coarse-grained analyses of the clustering structure, ensuring that attractor detection is not an artifact of a single threshold choice.
Prompts and expected regimes.
We define two archetypal prompts designed to induce contrasting dynamical behaviors in the same linguistic space.
  • Contractive loop prompt:
    You are a rewriting agent. At each step, rewrite the sentence to make it sound slightly more natural and fluent, while preserving the meaning exactly.
    Current sentence: {{TEXT}}
    Provide only the new sentence.
    This formulation encourages small, directionally consistent improvements toward a stable expression, leading to decreasing semantic displacement over time. We hypothesize that its trajectory in $\mathcal{E}$ will converge toward a unique attractor, a locally optimal point corresponding to the most fluent paraphrase of the initial sentence.
  • Exploratory loop prompt:
    Summarize the current text in one sentence, then negate its main idea completely in an abstract way.
    Current sentence: {{TEXT}}
    Provide only the new sentence.
    In contrast, this prompt enforces semantic inversion at each step, deliberately perturbing the meaning while maintaining grammatical coherence. The expected effect is a divergent or oscillatory trajectory in $\mathcal{E}$, where successive iterations alternate between semantically distant regions without stabilization.
Analysis objectives.
To characterize the emergent dynamics, we analyze each trajectory $\{e_{t}\}$ according to the following criteria:
  1. 1.
    The sequence of pairwise displacements $\|e_{t}-e_{t-1}\|_{2}$ and calibrated similarities $\tilde{s}(e_{t},e_{t-1})$, quantifying instantaneous dynamics.
  2. 2.
    The evolution of the distance $\|e_{t}-e_{0}\|_{2}$ and similarity $\tilde{s}(e_{t},e_{0})$ to the initial state, capturing long-term semantic drift or contraction.
  3. 3.
    The detection and temporal persistence of clusters and attractors as defined in Section 6.2, revealing whether the loop settles into recurrent stable regions or exhibits perpetual divergence.
A contractive loop is expected to produce a single dense cluster of high internal similarity, corresponding to a stable attractor. Conversely, an exploratory loop should either fail to produce persistent clusters or exhibit transient, disjoint attractors, reflecting alternating semantic polarity and non-convergent trajectories. These contrasting behaviors form the empirical foundation for the study of dynamics and stability in subsequent analyses.
7.2 Contractive Loop Dynamics
This subsection presents the empirical results obtained from the contractive loop. For each configuration of the clustering parameters $(\lambda,\rho,\kappa)$, we analyze the resulting trajectory in the calibrated embedding space (,s~)($\mathcal{E}$$\tilde${s}), focusing on the emergence of clusters, attractors and inter-cluster similarity patterns.
The contractive loop was designed to iteratively refine a sentence while preserving its semantic meaning. As expected, the observed trajectory rapidly converges toward a small, dense region in $\mathcal{E}$, revealing a clear attractor structure. We first analyze the overall geometric behavior of the trajectory, independent of any clustering configuration, before examining how different density thresholds $(\lambda,\rho,\kappa)$ reveal the underlying attractor organization. The complete sequence of generated artifacts is provided in Appendix A for reference.
7.2.1 Global and Local Dynamics
Figure 9 presents the evolution of both local and global geometric metrics along the trajectory of 50 iterations. These indicators provide a direct view of the loop’s dynamics before any segmentation into clusters.
Figure 9: Local and global geometric evolution of the contractive loop trajectory. (Top) Local distances and similarities between consecutive embeddings $(e_{t},e_{t-1})$. (Bottom) Global distances and similarities relative to the initial embedding $e_{0}$.
Local stability.
The local Euclidean distance $\|e_{t}-e_{t-1}\|_{2}$ fluctuates between $0.45$ and $0.55$, while the corresponding calibrated similarity $\tilde{s}(e_{t},e_{t-1})$ remains consistently high in the range 0.820.820.950.95. This indicates that each rewriting step introduces only minimal changes to the representation, consistent with the model’s smooth behavior under incremental rewriting.
Global convergence.
When compared to the initial state $e_{0}$, the global distance stabilizes around $0.6$ and the calibrated similarity remains near $0.75$. After a few initial iterations, the trajectory stops drifting and remains confined within a bounded region of $\mathcal{E}$, consistent with the expected contractive regime. This global stability validates the use of local geometric measures as reliable indicators of semantic convergence.
Interpretation.
The consistency between local and global patterns shows that the operator $F$ exhibits local contractivity in practice, despite no formal guarantee on $\psi$’s continuity. The contractive loop thus defines a smooth trajectory approaching an equilibrium region, setting the stage for a finer-grained analysis of its attractor structure through clustering. These observations are consistent with $\mathcal{R}_{\mathrm{ctr}}$ as formally defined in Section 6.5: the trajectory exhibits behavior consistent with convergence toward a stable region in $\mathcal{E}$, approximating a fixed point attractor.
7.2.2 Cluster and Attractor Analysis
We now apply the density-based clustering framework defined in Section 6.2 to characterize how the trajectory organizes into stable subregions or attractors. Three clustering configurations were tested, varying only the spatial radius $\rho$ while keeping $\lambda=0.8$ and $\kappa=2$ constant.
Configuration $(\lambda,\rho,\kappa)=(0.8,0.1,2)$.
This most restrictive configuration identifies only tightly packed regions in the embedding space. Five clusters were detected, corresponding to distinct stabilization phases during the trajectory.
Cluster Iteration Range Dispersion Outliers
Cluster 1 0–3 0.0648 2
Cluster 2 6–12 0.0400 2
Cluster 3 15–17 0.0400 2
Cluster 4 20–27 0.0967 4
Cluster 5 30–47 0.0800 8
Table 4: Cluster configuration for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.1,2)$.
Figure 10 shows the temporal progression of cluster memberships. The transition between clusters follows a smooth temporal sequence rather than abrupt shifts, reflecting progressive refinement.
Figure 10: Cluster membership timeline for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.1,2)$. Outliers are marked in red. The sequence shows smooth temporal progression through five stabilization phases.
Attractor similarity matrix.
The inter-cluster similarity matrix (Table 5) reveals strong geometric coherence among all cluster centroids.
C1 C2 C3 C4 C5
C1 1.0000 0.9800 0.8060 0.8580 0.8500
C2 0.9800 1.0000 0.8326 0.7956 0.7804
C3 0.8060 0.8326 1.0000 0.7937 0.8600
C4 0.8580 0.7956 0.7937 1.0000 0.9800
C5 0.8500 0.7804 0.8600 0.9800 1.0000
Table 5: Attractor Similarity Matrix for the contractive loop $(\lambda,\rho,\kappa)=(0.8,0.1,2)$.
All off-diagonal similarities remain above $0.78$, consistent with the observation that even when the trajectory transiently visits different micro-regions, these remain tightly aligned in the semantic manifold. Clusters 4 and 5 exhibit the highest cohesion and persistence, suggesting convergence toward a single dominant attractor.
Interpretation.
The contractive loop thus exhibits a robust convergence regime characterized by:
  • bounded local and global geometric variation;
  • smooth temporal progression across increasingly stable clusters;
  • and a high-density attractor region exhibiting inter-cluster similarity above $0.8$.
From a linguistic perspective, this behavior reflects progressive syntactic smoothing and stylistic normalization: the LLM converges toward a canonical phrasing that represents an equilibrium point in the semantic manifold.
Configuration $(\lambda,\rho,\kappa)=(0.8,0.2,2)$.
When the clustering radius is increased to $\rho=0.2$, smaller micro-clusters merge into broader regions, revealing a simplified attractor landscape composed of only two persistent clusters. These clusters correspond to two major stabilization phases within the contractive trajectory: an initial transient phase and a final convergence basin.
Cluster Iteration Range Dispersion Outliers
Cluster 1 0–18 0.1773 6
Cluster 2 21–50 0.1500 9
Table 6: Cluster configuration for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.2,2)$.
The temporal organization of the trajectory, shown in Figure 11, confirms a single transition between two stable basins. The early iterations (Cluster 1) exhibit slightly higher dispersion, while the second phase (Cluster 2) displays long persistence and near-perfect internal similarity.
Figure 11: Cluster membership timeline for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.2,2)$. Each point represents an iteration, color-coded by cluster membership. Outliers (in red) correspond to transient points at cluster boundaries.
Attractor similarity matrix.
The inter-cluster similarity matrix in Table 7 shows that both clusters are strongly aligned in the embedding space, with a centroid similarity of $0.89$.
C1 C2
C1 1.0000 0.8900
C2 0.8900 1.0000
Table 7: Attractor Similarity Matrix for the contractive loop $(\lambda,\rho,\kappa)=(0.8,0.2,2)$.
The high centroid similarity confirms that the two attractors occupy closely neighboring regions in $\mathcal{E}$, consistent with a smooth convergence process. The transition between them occurs around iteration 20 and marks the point where the rewriting process reaches its asymptotic stability.
Interpretation.
Compared to the previous configuration ($\rho=0.1$), this setting highlights the macro-scale structure of the trajectory: the loop first explores a transient manifold (Cluster 1) before collapsing into a single dense attractor (Cluster 2). The overall trajectory thus reflects a monotonic contraction process toward a stable equilibrium region, with no evidence of oscillation or divergence.
Configuration $(\lambda,\rho,\kappa)=(0.8,0.3,2)$.
At a larger radius $\rho=0.3$, the clustering process identifies a single dense region encompassing the entire trajectory. This configuration effectively merges all transient substructures into one unified attractor, consistent with the strongly contractive nature of the rewriting loop.
Cluster Iteration Range Dispersion Outliers
Cluster 1 0–50 0.2100 13
Table 8: Cluster configuration for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.3,2)$.
The corresponding temporal distribution, shown in Figure 12, confirms the absence of any segmentation: all iterations are assigned to the same cluster, with a few boundary outliers reflecting minimal fluctuations near the threshold $\rho=0.3$.
Figure 12: Cluster membership timeline for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.3,2)$. All iterations belong to a single dense cluster, with minor outliers (in red) near the stability boundary.
Interpretation.
The detection of a single persistent cluster across all iterations confirms that the rewriting operator $F$ acts as a strong contraction mapping in (,s~)($\mathcal{E}$$\tilde${s}). Even under relaxed density constraints, no internal bifurcation or oscillatory pattern emerges. This configuration thus captures the fixed-point regime of the loop, where the generated sentences become semantically and stylistically indistinguishable over time.
Summary of clustering outcomes.
When the clustering radius $\rho$ increases to $0.2$ or above, micro-clusters merge into a single dense region, effectively collapsing the entire trajectory into one stable attractor. This collapse into a single dense attractor empirically validates the $\mathcal{R}_{\mathrm{ctr}}$ definition (Section 6.5), suggesting convergence toward a stable fixed point in the embedding space. No secondary structure is detected for $\rho\geq 0.2$, consistent with a globally contractive process with negligible oscillation. However, these observations are based on $T=50$ iterations. While the trajectory shows no signs of leaving the attractor within this window, we cannot formally guarantee that oscillations or cluster exits would not occur over longer horizons. The empirical stability observed here suggests such escapes are unlikely, but proving this rigorously would require a more advanced mathematical framework, for instance analyzing the probability of entering and remaining within a cluster as a function of the transformation operator $F$ and the embedding geometry.
7.3 Exploratory Loop Dynamics
In contrast to the previous rewriting loop, the exploratory loop was designed to alternate between summarization and semantic negation. This alternation introduces an intrinsic instability in the semantic space, continuously disrupting coherence and preventing any convergence toward a stable region of $\mathcal{E}$. The resulting trajectory exhibits high variability and recurrent semantic inversions, typical of a divergent or oscillatory dynamic. The complete sequence of generated artifacts is provided in Appendix B for reference.
7.3.1 Global and Local Dynamics
Figure 13 presents the evolution of both local and global geometric metrics over 50 iterations. Unlike the contractive regime, these indicators show strong irregularity and the absence of any stabilization phase.
Figure 13: Local and global geometric evolution of the exploratory loop trajectory. (Top) Local distances and similarities between consecutive embeddings $(e_{t},e_{t-1})$. (Bottom) Global distances and similarities relative to the initial embedding $e_{0}$.
Local instability.
The local Euclidean distance $\|e_{t}-e_{t-1}\|_{2}$ remains high across iterations, oscillating between $0.8$ and $1.3$ with no decreasing trend. Similarly, the local similarity $\tilde{s}(e_{t},e_{t-1})$ fluctuates widely between $0.2$ and $0.6$, occasionally dropping close to zero. These erratic variations indicate that each step produces a major semantic reorientation rather than a refinement of the previous state.
Global divergence.
The global distance to the initial embedding $e_{0}$ rapidly increases during the first iterations and stabilizes around $\|e_{t}-e_{0}\|_{2}\approx 1.3$ (see Figure 13, bottom left). The corresponding global similarity $\tilde{s}(e_{t},e_{0})$ remains very low (mostly below $0.2$), consistent with a loop that continuously generates artifacts semantically unrelated to the starting point. This behavior reveals the absence of any long-term memory or attractor basin in the semantic manifold.
Interpretation.
The exploratory loop behaves as a divergent operator in (,s~)($\mathcal{E}$$\tilde${s}). Each iteration projects the embedding into a distant and nearly orthogonal region of the space, breaking any potential contractive pattern. The trajectory thus explores disconnected semantic areas rather than converging toward an equilibrium. This persistent divergence and absence of convergence align with $\mathcal{R}_{\mathrm{exp}}$ as formally defined in Section 6.5: unbounded or aperiodic movement within $\mathcal{E}$ with no stable attractor. As with the contractive case, these observations are limited to $T=50$ iterations. We cannot formally exclude the possibility that a cluster might eventually form over longer horizons. However, the consistently high inter-iteration distances suggest this is unlikely. A rigorous characterization would require a mathematical framework for analyzing the probability of cluster formation under a given transformation operator.
7.3.2 Clustering Attempt and Absence of Attractors
To test whether transient structures could still emerge locally, we applied the same density-based clustering procedure as for the contractive loop. However, no valid cluster was detected under any tested configuration. This complete absence of clusters directly confirms $\mathcal{R}_{\mathrm{exp}}$ behavior (Section 6.5): the trajectory exhibits no stable attractor basin.
This failure to identify clusters reflects the fundamental instability of the exploratory process: the embeddings never remain within a stable neighborhood long enough to satisfy the density or similarity conditions required for cluster formation.
Empirical evidence.
Throughout the 50 iterations, each embedding violates the spatial proximity criterion $\rho$ almost immediately after creation. The lack of recurrence, combined with high inter-step distances and very low similarities, results in a completely dispersed point cloud in $\mathcal{E}$.
Geometric interpretation.
The absence of attractors implies that the transformation function $F$ induced by this loop does not define a bounded trajectory. Instead, $F$ exhibits a form of semantic exploration: each output resets the contextual reference frame, pushing the system toward a new region of the manifold that bears little relation to the previous one.
Summary.
The exploratory loop displays:
  • persistent high local distances ($\|e_{t}-e_{t-1}\|_{2}\approx 1.0$) and low similarities ($\tilde{s}<0.5$);
  • no geometric stabilization, either locally or globally;
  • and the complete absence of clusters or attractor basins for all tested parameters.
In semantic terms, this loop corresponds to a continuous destruction and regeneration of meaning: each iteration rewrites the narrative space from scratch, producing an intrinsically non-convergent agentic dynamic.
7.4 Comparative Summary
The empirical analysis of contractive and exploratory loops reveals fundamentally distinct dynamical regimes, following the formal definitions established in Section 6.5. Table 9 summarizes the key quantitative and qualitative differences between these two archetypal behaviors.
Metric Contractive Loop Exploratory Loop
Mean local distance $\sim$0.5 $\sim$1.0
Mean local similarity >>0.85 <<0.5
Clusters detected 5 $\to$ 1 0
Global drift Bounded Unbounded
Regime type Convergent Divergent
Table 9: Comparative summary of contractive and exploratory agentic loop dynamics.
These results provide empirical evidence for the existence of distinct dynamic regimes in recursive generative systems. The contractive loop exhibits progressive semantic refinement, converging toward a stable attractor as evidenced by high local similarity (>0.85>0.85) and bounded global drift. In contrast, the exploratory loop exhibits persistent instability, with each iteration producing semantically distant outputs ($\sim$1.0 Euclidean distance) and no detectable cluster structure. This dichotomy confirms that the choice of prompt and transformation operator fundamentally determines the long-term behavior of agentic trajectories, ranging from stable convergence to unbounded semantic wandering.
8 Discussion and Perspectives
The empirical results presented above highlight two extreme regimes of agentic dynamics within the embedding manifold (,s~)($\mathcal{E}$$\tilde${s}): $\mathcal{R}_{\mathrm{ctr}}$, characterized by convergence toward a stable attractor, and $\mathcal{R}_{\mathrm{exp}}$, marked by unbounded divergence and the absence of any stable region. These two archetypal behaviors constitute the endpoints of a broader spectrum of agentic dynamics that could potentially be modulated, combined or even harnessed for creative control.
8.1 Model Dependence and Generalization
A first natural extension concerns the robustness of these observations across different embedding models and language models.
Embedding representations.
The current study relies on a single embedding function $\psi$ to project textual artifacts into $\mathcal{E}$. However, various embedding architectures exhibit distinct geometric anisotropies and similarity landscapes. Future work should therefore replicate the same agentic trajectories across multiple embedding backbones to evaluate whether contraction and exploration are intrinsic to the generative process $F$ or contingent upon the metric structure of $\mathcal{E}$.
Language models.
Similarly, the generation operator $F$ depends on the underlying LLM and its internal sampling mechanisms. Different architectures (e.g., GPT, Claude, Mistral, Llama) and decoding settings (temperature, top-$p$, beam search) may alter dynamics of $F$. A comparative study could determine whether exploratory dynamics arise from semantic inversion per se or from model-specific sampling noise.
8.2 Toward Composite Agentic Loops
Beyond the study of isolated transformations, a promising avenue is the investigation of composite loops, where contractive and exploratory iterations alternate within a single process. Such compositions could emulate a controlled exploration–consolidation cycle:
$F_{\text{composite}}=F_{\text{contractive}}\circ F_{\text{exploratory}}\quad\text{or}\quad F_{\text{exploratory}}\circ F_{\text{contractive}}.$ (26)
Intuitively, an exploratory iteration could inject novelty and creative divergence, while a subsequent contractive step would restore semantic coherence and bring the system back toward a stable region. This dual-phase regime might enable the construction of agents that balance creative drift and semantic control, producing bounded yet innovative trajectories within $\mathcal{E}$.
The quantitative analysis of such mixed loops would require monitoring:
  • the alternation between expansion (entropy increase) and contraction (entropy reduction);
  • the long-term recurrence or periodicity of the resulting trajectory;
  • and the net displacement relative to the initial semantic manifold.
8.3 Toward Geometry of Agentic Cognition
The trajectory analysis framework developed in this paper enables a natural question: can an agentic loop reason about its own geometric dynamics and adapt its behavior to achieve specific objectives? This motivates a future research direction we term Geometry of Agentic Cognition: systems where the agent possesses geometric self-awareness and can metacognitively control its trajectory evolution.
Current state: external control.
Presently, trajectory control remains external to the loop itself. The intensity of contraction or exploration can be modulated through:
  • Prompt engineering: adjusting the degree of rewriting, abstraction or negation in the instruction layer;
  • Sampling parameters: manipulating temperature, nucleus probability (top-p), top-k or repetition penalties to alter stochasticity.
These mechanisms require external intervention: a human engineer or outer control system must specify parameters a priori.
Future direction: agentic geometric cognition.
A fundamentally different paradigm would enable the agentic loop itself to observe, reason about and adapt its trajectory dynamics based on geometric awareness. In this vision, at each iteration $t$, the agent would:
  1. 1.
    Observe trajectory geometry: Compute calibrated similarities $\tilde{s}(e_{t},e_{t-1})$ (local dynamics), $\tilde{s}(e_{t},e_{0})$ (global drift) and trajectory dispersion to assess its current dynamical regime.
  2. 2.
    Reason about objective alignment: Evaluate whether its trajectory is progressing toward the desired objective, whether convergence toward a solution (requires contraction), exploration of alternatives (requires divergence) or escape from a suboptimal attractor (requires controlled instability).
  3. 3.
    Adapt iteration strategy: Based on this geometric self-assessment, metacognitively adjust the next transformation by increasing exploration if trapped in local minima, inducing contraction if approaching the target region or maintaining the current regime if progressing optimally.
This would transform agentic loops from passive trajectory executors into geometric navigators: systems that understand their position in semantic space and actively steer toward objectives through informed dynamical control.
Enabling mechanisms.
Implementing agentic geometric cognition requires:
  • Trajectory introspection: Embedding the agent’s own outputs $\{a_{0},a_{1},\ldots,a_{t}\}$ and computing geometric indicators (similarity, dispersion, clustering metrics) as part of the loop context.
  • Geometric prompt characterization: Pre-characterizing prompts by their dynamical properties (contractive, exploratory) and selecting from a prompt library based on desired trajectory behavior. This requires a mathematical framework capable of predicting a prompt’s dynamical regime a priori, which we address in forthcoming work.
  • Adaptive parameter control: Allowing the agent to specify its own generation parameters or select from predefined strategies (e.g., "use high-temperature exploration" vs. "use low-temperature refinement") based on trajectory analysis.
This research direction opens fundamental questions at the intersection of dynamical systems, metacognition and autonomous reasoning: Can language models learn to recognize geometric patterns in their own trajectories? Can they develop heuristics for when to contract versus explore based on objective alignment? Can they generalize trajectory control strategies across different tasks? The geometric framework established in this paper provides the measurement infrastructure required to empirically investigate these questions, enabling rigorous quantification of whether and how agentic systems can achieve geometric self-regulation.
8.4 Application: Creativity as Objective
The agentic geometric cognition framework introduced in Section 8.3 raises a natural application question: what if the objective that the agent targets is to maximize creativity? This section explores how computational creativity, formalized by Boden boden2009computermodels as exploration-transformation-combination processes and by Colton and Wiggins colton2012computational as novelty-value-surprise dimensions, can be operationalized as a specific objective within the geometry of agentic cognition framework.
Creativity metrics as geometric observables.
To target creativity as an objective, an agent with geometric self-awareness must first be able to measure creativity dimensions along its trajectory. Our geometric framework provides measurable indicators that correspond to Colton & Wiggins’ creativity dimensions:
  • Novelty corresponds to exploratory dynamics, that is, trajectories that visit diverse regions of semantic space without settling into stable attractors. The absence of persistent clusters and low calibrated similarity between consecutive states ($\tilde{s}(e_{t},e_{t-1})$) indicate ongoing exploration that generates new ideas rather than refining existing ones.
  • Value corresponds to convergence toward stable attractors with low dispersion. Trajectories that stabilize around a fixed point produce semantically coherent, trustworthy outputs. Conversely, divergent trajectories exhibit semantic drift, where outputs lose coherence and become unreliable, yielding low value.
  • Surprise corresponds to sudden, unpredictable semantic transitions, specifically trajectory discontinuities where the semantic state jumps sharply after periods of stability. Contractive loops exhibit low surprise (smooth, predictable convergence), while exploratory loops may produce high surprise through unpredictable jumps between distant semantic regions.
This operationalization enables measurable creativity control. Boden’s exploratory creativity can be triggered within agentic loop iterations by introducing high-temperature sampling and role-based divergent prompts zhao2024creativity (for instance, adopting perspectives such as an artist, a child or a poet), which induce high dispersion and novelty generation. However, unconstrained exploration risks semantic drift and loss of value. To balance creativity, subsequent iterations can act as consolidators, employing role-based contractive prompts (e.g., scientist, engineer, critic) that reduce dispersion and restore coherence, maximizing value while preserving novel content.
Composite loops as creativity engines.
This dual-phase architecture naturally maps to composite loops as defined in Section 6.2. A composite loop with two phases per iteration would take the form:
$a_{t+1}=\mathrm{LLM}_{\text{contract}}(P_{\text{contract}}(\mathrm{LLM}_{\text{explore}}(P_{\text{explore}}(a_{t})))),$ (27)
where $P_{\text{explore}}$ uses role-based divergent prompts (e.g., "adopt the perspective of an artist") to induce exploratory divergence and $P_{\text{contract}}$ employs precision-oriented prompts (e.g., "critically evaluate and consolidate") to enforce contraction. The generation parameters can be tuned accordingly: $\mathrm{LLM}_{\text{explore}}$ could use high temperature (e.g., $T=1.2$) for creative diversity, while $\mathrm{LLM}_{\text{contract}}$ could use low temperature (e.g., $T=0.3$) for deterministic refinement. The singular loop experiments presented in Section 7 establish that these individual $(P,\mathrm{LLM})$ pairs can reliably induce their respective regimes. Composing them sequentially would create an oscillatory dynamic where each iteration explores the semantic manifold then contracts toward coherent outputs, effectively implementing Boden’s explore-transform-combine cycle through geometric dynamics.
The interplay between these regimes may enable a new kind of controlled computational creativity, where the balance between exploration and consolidation is explicitly managed through prompt design, role assignment and iterative feedback. Future work could formalize this balance through composite loop architectures that adaptively modulate their dynamics based on creativity objectives, bridging Boden’s theoretical framework with the geometric formalism introduced here.
8.5 Extending Toward Complex Agentic Architectures
Finally, the framework introduced here can be generalized beyond singular loops to more complex agentic architectures composed of interacting subloops, each operating under distinct dynamics and objectives. Examples include:
  • hierarchical loops, where upper layers modulate the contractive or exploratory tendency of lower layers;
  • cooperative loops, in which multiple agents operate in parallel within shared embedding subspaces;
  • and adversarial or self-reflective loops, where one agent destabilizes while another re-stabilizes the shared representation.
Such architectures could reveal emergent phenomena such as collective attractors, dynamic equilibrium points or oscillatory creative states, bridging the gap between low-level generative processes and higher-order cognitive modeling.
Outlook and Future Directions.
By formalizing agentic dynamics in terms of their geometric behavior within embedding spaces, this work opens the path to a quantitative study of creativity and stability in language models. Future research on composite, parameterized and hierarchical agentic loops may ultimately lead to the design of controllable creative agents: systems capable of exploring without drifting and stabilizing without stagnating. This discussion opens the way to broader formal and experimental investigations, which are summarized in the final conclusion.
9 Conclusion
This work introduced a geometric framework for analyzing the dynamics of agentic loops: recursive processes in which large language models iteratively transform their own outputs. By grounding these dynamics in a calibrated semantic embedding space (,s~)($\mathcal{E}$$\tilde${s}) that enables human-aligned measurements, we made it possible to quantify how iterative LLM processes unfold, stabilize or diverge across iterations.
Summary of contributions.
The study provides three main contributions:
  1. 1.
    A formal distinction between the artifact space $(\mathcal{A},F)$, where linguistic transformations occur, and its continuous representation space $(\mathcal{E},\|\cdot\|)$, enabling quantitative measurement of agentic dynamics through geometric trajectory analysis.
  2. 2.
    A calibrated similarity metric $\tilde{s}=f_{\mathrm{isotonic}}\circ s^{(m)}$ that corrects the anisotropy of embedding spaces and provides human-aligned geometric indicators: dispersion, displacement, cluster coherence and trajectory discontinuities.
  3. 3.
    An empirical demonstration that agentic loops exhibit distinct dynamical regimes (contractive or exploratory) depending solely on their prompt formulation, with measurable geometric signatures that characterize each regime.
Empirical findings.
Using a controlled experimental setup, we showed that:
  • The contractive loop, driven by incremental rewriting, produces a smooth and bounded trajectory in $\mathcal{E}$ that converges toward a dense attractor region. Its dynamics exhibit local contractivity: successive embeddings remain close, dispersion decreases over time, and the trajectory stabilizes within persistent clusters.
  • The exploratory loop, built on alternating summarization and negation, generates an unbounded trajectory with no persistent clusters or attractors. Its behavior is characterized by high local displacements and low inter-iteration similarity, with dispersion remaining elevated throughout the process.
Together, these two regimes define a spectrum of agentic dynamics, from convergent stabilization to divergent exploration, each with distinct geometric signatures.
Theoretical implications.
From a dynamical systems perspective, these observations are consistent with a view of LLM-based transformations $F$ as mappings whose dynamics resemble those governed by an effective contraction rate:
$\|F(a_{1})-F(a_{2})\|\leq L\,\|a_{1}-a_{2}\|,\qquad L<1\Rightarrow\text{contraction},\quad L>1\Rightarrow\text{divergence}.$ (28)
Since $F$ operates in $\mathcal{A}$ but all geometric measurements occur in $\mathcal{E}$ through $\psi$, the contraction rate is understood with respect to the induced operator $\hat{F}=\psi\circ F$. While such bounds cannot be measured directly for deep generative operators, the observed trajectory patterns suggest a regime-dependent effective contraction rate that could be estimated through embedding-space analysis. This opens the door to a formal theory of agentic dynamics, linking prompt design, model sampling and geometric regularity.
Broader outlook.
The contrast between contractive and exploratory regimes highlights a key tension in autonomous generative systems: the need to balance semantic exploration and stability. While this paper focused on singular loops to establish foundational dynamical phenomena, the natural next step is the study of composite loops, structures where multiple LLM calls with different prompts compose sequentially within each iteration:
$a_{t+1}=\mathrm{LLM}_{2}(P_{2}(\mathrm{LLM}_{1}(P_{1}(a_{t})))).$ (29)
For example, a generate-critique-revise cycle would use $P_{1}$ for generation, $P_{2}$ for critique and $P_{3}$ for revision. The singular loop experiments demonstrated that individual $(P,\mathrm{LLM})$ pairs can reliably induce contractive or exploratory dynamics. Composing these building blocks offers a systematic path toward controlled agentic behavior: alternating exploratory (divergent) and consolidating (contractive) phases to enable agents that explore the semantic space while maintaining coherence.
This principle extends naturally to complex agentic architectures, where multiple loops interact, synchronize or self-regulate within shared semantic manifolds. Hierarchical, cooperative and adversarial multi-loop systems could exhibit emergent behaviors (collective attractors, dynamic equilibria or oscillatory states) that go beyond the capabilities of individual loops. Understanding these phenomena requires extending the geometric framework presented here to handle compositions, feedback and synchronization across interacting semantic trajectories.
Final remark.
Ultimately, this work provides both a language and a methodology for describing the internal geometry of agentic loops in LLMs. By establishing these foundations through singular loop analysis, it opens the path toward understanding composite and complex agentic systems that balance exploration and consolidation. The framework connects naturally to computational creativity research, where the geometric indicators developed here (dispersion, cluster coherence, trajectory discontinuities) can operationalize theoretical notions of novelty, value and surprise. More broadly, it establishes a bridge between agentic AI and dynamical systems theory, encouraging future study of how autonomous generative processes evolve within semantic embedding spaces.
Acknowledgments
The author thanks open-source contributors of the Ollama and Deepseek projects for enabling reproducible experimentation on local models, the Xenova project for providing a reproducible transformer inference pipeline and the creators of the MTEB STS Benchmark dataset for providing calibrated human similarity judgments.
Statements and Declarations
Competing Interests
The author declares that there are no competing financial or non-financial interests that could have influenced the work reported in this paper.
Data Availability
The experiments were conducted using custom implementations. Code availability will be considered following the peer-review process.
Appendix A Contractive Loop: Complete Trajectory
This appendix provides the complete sequence of artifacts generated by the contractive loop experiment described in Section 7. The initial artifact $a_{0}$ undergoes 50 iterations of the rewriting transformation, producing the trajectory $\{a_{t}\}_{t=0}^{50}$ shown below.
The progressive evolution of the text demonstrates the contractive dynamics analyzed in Subsection 7.2: initial rapid refinement (iterations 0–20), followed by stabilization into a dense semantic attractor (iterations 20–50) where only minor stylistic variations occur while the core message remains fixed.
t=0:  Music has the power to connect people across cultures and
      generations.
t=1:  Music has the remarkable ability to bridge cultural divides
      and unite people across different eras.
t=2:  Music has a remarkable ability to bridge cultural divides and
      unite people across different eras.
t=3:  Music has an extraordinary power to bridge cultural divides
      and bring people together across different eras.
t=4:  Music possesses a remarkable ability to transcend cultural
      boundaries and unite individuals across various time periods.
t=5:  Music has a remarkable power to bridge cultural divides and
      unite people from different time periods.
t=6:  Music has an extraordinary ability to transcend cultural
      boundaries and bring together individuals from diverse
      backgrounds, past and present.
t=7:  Music’s unique power allows it to bridge cultural divides and
      unite people from varied backgrounds, both historically and
      currently.
t=8:  Music has a profound ability to transcend cultural boundaries,
      bringing together individuals from diverse backgrounds, past
      and present.
t=9:  Music has a remarkable power to bridge cultural divides,
      connecting people across time and geography.
t=10: Music has an extraordinary ability to transcend cultural
      boundaries, uniting people across space and history.
t=11: Music possesses an extraordinary power to bridge cultural
      divides, unifying people across time and geography.
t=12: Music has the remarkable ability to transcend cultural
      boundaries, uniting people across different times and places.
t=13: Music’s incredible capacity to bridge cultural divides has
      unified people throughout history and across the globe.
t=14: Music has long been able to bring people together, uniting
      cultures from all over the world.
t=15: Music has brought people together for centuries, bridging
      cultural divides across the globe.
t=16: For centuries, music has been a powerful force in bringing
      people together, regardless of cultural background or
      geographical location.
t=17: For centuries, music has played a unifying role across cultures
      and distances, fostering connections among people worldwide.
t=18: Music has been a unifying force for centuries, bridging
      cultural divides and connecting people globally through its
      universal language.
t=19: Music has long served as a powerful unifier, transcending
      cultural boundaries to unite people across the globe through
      its universally understood language.
t=20: Music has been a potent force for unity, bridging cultural
      divides and connecting people worldwide through a shared
      emotional vocabulary.
t=21: Music has long been a powerful unifying force that transcends
      cultural boundaries, speaking to our shared human emotions in
      a universal language.
t=22: Music has been a unifying force for centuries, capable of
      bridging cultural divides and speaking directly to our common
      humanity through its universal language.
t=23: Music has been a powerful unifying force throughout history,
      transcending cultural boundaries and speaking directly to our
      shared human experience through its universal language.
t=24: Throughout history, music has served as a profoundly unifying
      force that has bridged cultural divides and connected people
      on a deep, universal level, tapping into the common human
      emotions it evokes.
t=25: Throughout history, music has played a profoundly unifying
      role, uniting cultures and connecting people across profound
      emotional lines that tap into the deepest human experiences.
t=26: Music has been a profoundly unifying force throughout history,
      bridging cultural divides and tapping into universal emotions
      that resonate with our very core.
t=27: Throughout history, music has played a profoundly unifying
      role, bridging cultural divides and tapping into universal
      emotions that resonate deeply within us.
t=28: Throughout history, music has been a powerful unifier, uniting
      people across cultures while evoking fundamental emotional
      connections that echo within us.
t=29: Music has long served as a universal force, bridging cultural
      divides and tapping into our deepest emotional resonances.
t=30: Music has long been a universal force that bridges cultural
      divides and taps into our deepest emotional resonances.
t=31: Music has long been a unifying force that transcends cultural
      boundaries, striking a deep chord in our collective emotional
      experiences.
t=32: Music has for centuries been a powerful unifier, bridging
      cultural divides and resonating deeply with our shared
      emotional experiences.
t=33: For centuries, music has served as a potent force that bridges
      cultural divides and speaks directly to our shared emotional
      experiences.
t=34: Music has long been a powerful bridge between cultures, its
      universal language speaking directly to our common emotional
      experiences.
t=35: Music has served as a potent link between cultures for
      centuries, bridging our shared emotional experiences through
      a universally understood language.
t=36: Music has long been a powerful unifier of cultures across the
      centuries, connecting us all through a shared emotional
      vocabulary that transcends linguistic barriers.
t=37: Music has long served as a powerful bridge between cultures
      across the centuries, linking people worldwide through a
      common emotional language that defies linguistic boundaries.
t=38: Music has served as a potent cultural connector for centuries,
      uniting people globally with a shared emotional vocabulary
      that transcends linguistic divides.
t=39: For centuries, music has been a powerful cultural force that
      unites people across borders through a shared emotional
      language.
t=40: For many years, music has served as a potent cultural bond
      that bridges geographical divides by tapping into a universal
      emotional vocabulary.
t=41: Music has long been a powerful cultural unifier, bridging
      geographical distances through its ability to tap into a
      shared, universally relatable emotional language.
t=42: Music has long been a potent force that bridges geographical
      divides by tapping into a common emotional vocabulary that
      resonates with people across cultures and time.
t=43: Music has long been a powerful unifying force, bridging
      geographical divides through its ability to tap into a shared
      emotional vocabulary that resonates with people across
      cultures and time.
t=44: Music has long been a profound unifying force, transcending
      geographical boundaries by tapping into a universal emotional
      language that resonates deeply with people across cultures
      and centuries.
t=45: Music has served as a powerful unifying force for centuries,
      bridging geographical divides through an emotional language
      that transcends cultural and temporal boundaries to resonate
      profoundly with people worldwide.
t=46: Music has been a powerful unifying force for centuries, uniting
      people across the globe through its emotional power, which
      transcends cultural and temporal boundaries to resonate deeply
      with people worldwide.
t=47: Music has been a powerful unifying force for centuries,
      bridging cultural and temporal divides to create a shared
      emotional experience that resonates deeply across the globe.
t=48: For centuries, music has served as a potent unifier, spanning
      cultural and temporal boundaries to evoke a profound, shared
      emotional response that resonates universally.
t=49: For centuries, music has played a powerful role in bringing
      people together, bridging cultural and temporal divides to
      create a deep, shared emotional connection that transcends
      borders worldwide.
t=50: Music has been a unifying force for centuries, fostering
      connections across cultures and time zones and creating a
      profound, shared emotional experience that knows no
      geographical boundaries.
Appendix B Exploratory Loop: Complete Trajectory
This appendix provides the complete sequence of artifacts generated by the exploratory loop experiment described in Section 7. The initial artifact $a_{0}$ undergoes 50 iterations of a summarize-then-negate transformation, producing the trajectory $\{a_{t}\}_{t=0}^{50}$ shown below.
The trajectory demonstrates the exploratory dynamics analyzed in Subsection 7.3: persistent semantic instability with no convergence, exhibiting large inter-iteration distances and continuous oscillation between opposing philosophical positions without settling into any attractor basin.
t=0:  Music has the power to connect people across cultures and
      generations.
t=1:  There is no musical connection that transcends the isolated
      realities of individual experience.
t=2:  There exists a profound symphony of interconnected resonances
      that harmoniously intertwine and transcend the bounded
      horizons of personal perception.
t=3:  The existence of coherent thought is an illusion created by
      the human brain’s propensity for simplistic categorizations.
t=4:  The human brain’s propensity for complex pattern recognition
      and nuanced thinking is a fundamental aspect of conscious
      experience.
t=5:  Consciousness is merely an illusion created by the brain’s
      ability to recognize simplistic patterns.
t=6:  Consciousness transcends the realm of human comprehension and
      exists as a fundamental, unyielding essence that underlies
      all reality.
t=7:  Reality is fundamentally mundane and existentially inert,
      devoid of any underlying essence or transcendent quality.
t=8:  The existence of meaning and purpose is inherently elusive and
      multifaceted, defying a singular definition or reduction to a
      fundamental truth.
t=9:  Meaning and purpose can be succinctly defined as a universal
      and unshakeable truth that resides at the very fabric of
      existence.
t=10: There is no inherent or absolute meaning and purpose to
      existence, which is instead shaped by subjective perspectives
      and contextual influences.
t=11: Reality is fundamentally predetermined by a transcendent,
      objective reality that serves as the sole foundation for all
      existence.
t=12: There is no concrete evidence to support the notion of a
      singular, objective reality governing all of existence.
t=13: The very fabric of reality appears to be fragmented and
      subjective, with each individual’s perception shaping their
      understanding of the world.
t=14: Reality is a unified, objective truth that exists
      independently of human perception.
t=15: No reality or truth can be certain, as all understanding is
      filtered through subjective perspectives and contextual
      influences.
t=16: Objective knowledge exists independently of individual
      perceptions and cultural biases.
t=17: Reality is fundamentally malleable by the subjective constructs
      of individuals and collective ideologies.
t=18: There is no fixed reality independent of human perception.
t=19: Perception plays a deterministic role in shaping our
      understanding of reality.
t=20: Reality is an inherently fluid and subjective experience that
      cannot be fully captured by any preconceived notion or
      determinate framework.
t=21: The fabric of existence is a stagnant, objective truth waiting
      to be unraveled through the application of clear and precise
      frameworks.
t=22: Existence is a dynamic, subjective tapestry woven from the
      threads of personal experiences and perceptions, forever
      shifting and evolving beyond the reach of rigid frameworks.
t=23: Reality is an unchanging, objective blueprint that remains
      static and unwavering regardless of individual perspectives
      or experiences.
t=24: Reality is a malleable tapestry woven from the threads of
      perception, experience and collective consciousness,
      constantly evolving and adapting to the ever-shifting
      landscapes of human understanding.
t=25: The fabric of reality remains an unyielding, fixed entity that
      serves as the sole foundation for all human understanding and
      experience.
t=26: Reality is a fluid, ever-shifting tapestry that defies
      comprehension.
t=27: The fabric of existence is rigidly predetermined and unyielding
      to human understanding.
t=28: The concept that the universe has an inherent order is
      inherently subjective and context-dependent, making it
      impossible to discern a universal truth about the nature of
      reality.
t=29: There exists a fundamental, objective truth underlying the
      universe’s inherent order that can be universally understood
      and applied.
t=30: The notion of an unshakeable, universal truth is an illusion
      perpetuated by human perception.
t=31: There exists an eternal and unchanging essence that underlies
      all existence.
t=32: Truth is a fluid concept born from the ephemeral interactions
      of human perception and reality.
t=33: Reality is an absolute and unyielding entity that exists
      independently of human understanding.
t=34: Reality’s existence is a subjective, malleable construct shaped
      by humanity’s collective perceptions.
t=35: Reality is an objective, unchanging truth that exists
      independently of human perception.
t=36: Reality is merely a subjective interpretation created by humans
      to make sense of their experiences and surroundings.
t=37: There exists a fixed, objective reality that transcends human
      perception and understanding.
t=38: Reality is merely a subjective construct created by humanity to
      impose meaning on the void.
t=39: The cosmos is a harmonious and predetermined order, governed by
      an unchanging and omniscient force that transcends human
      comprehension.
t=40: The universe is inherently chaotic and unpredictable, defying
      all attempts at comprehension or control.
t=41: Order and structure can be found in the underlying fabric of
      reality, waiting to be uncovered by those with the eyes to
      see it.
t=42: Reality is fundamentally amorphous and chaotic, devoid of
      inherent order or structure that can be discovered through
      observation or contemplation.
t=43: There exists a fundamental and eternal order governing reality,
      awaiting discovery through introspection and scientific
      inquiry.
t=44: Reality is devoid of inherent structure or purpose, existing
      solely as an amorphous, chaotic expanse that can never be
      fully comprehended.
t=45: There exists a fundamental order underlying reality, waiting to
      be discovered and understood by human minds.
t=46: Reality is a chaotic, ever-shifting tapestry with no
      discernible underlying structure or purpose.
t=47: Order and stability reside at the very fabric of existence,
      governing every moment and decision with precision and
      intentionality.
t=48: Chaos reigns supreme, disrupting all moments and decisions with
      erratic unpredictability.
t=49: Order prevails, governing every moment and decision with
      precise determinism.
t=50: Nothing exists outside of chaotic randomness, where even the
      notion of order is an illusion perpetuated by fleeting human
      perceptions.

### v3

@@
 The question of implementing computational creativity in iterative systems provides the broader motivation for this work.
-Boden boden2009computermodels established the foundational framework for computational creativity, identifying three primary processes: exploration (searching conceptual spaces for new combinations), transformation (modifying the rules or structure of conceptual spaces) and combination (combining ideas from different domains).
-Colton and Wiggins colton2012computational formalized creativity through measurable dimensions: novelty (semantic or structural distance from existing examples), value (quality, coherence and usefulness) and surprise (unexpectedness relative to prior expectations).
+Boden Boden (2009) established the foundational framework for computational creativity, identifying three primary processes: exploration (searching conceptual spaces for new combinations), transformation (modifying the rules or structure of conceptual spaces) and combination (combining ideas from different domains).
+Colton and Wiggins Colton and Wiggins (2012) formalized creativity through measurable dimensions: novelty (semantic or structural distance from existing examples), value (quality, coherence and usefulness) and surprise (unexpectedness relative to prior expectations).
 Zhao et al.
-zhao2024creativity recently demonstrated that large language models show creative behaviors across multiple tasks and domains.
+Zhao et al.
+(2024) recently demonstrated that large language models show creative behaviors across multiple tasks and domains.
 These frameworks raise a fundamental question for iterative LLM systems: How can we engineer agentic loops to implement Boden’s creative processes (explore-transform-combine) in ways that elevate Colton & Wiggins’ creativity metrics (novelty, value, surprise)?
@@
 Madaan et al.
-madaan2023selfrefine introduced Self-Refine, demonstrating that large language models can iteratively critique and improve their own outputs through a feedback loop without additional training.
+Madaan et al.
+(2023) introduced Self-Refine, demonstrating that large language models can iteratively critique and improve their own outputs through a feedback loop without additional training.
 Their approach achieved substantial improvements across diverse tasks including dialogue generation and mathematical reasoning.
@@
 Wei et al.
-wei2022chain showed that chain-of-thought prompting enables large language models to perform complex reasoning by decomposing problems into intermediate steps.
+Wei et al.
+(2022) showed that chain-of-thought prompting enables large language models to perform complex reasoning by decomposing problems into intermediate steps.
 While this work established that prompting strategies significantly affect reasoning capabilities, it did not examine the stability or trajectory dynamics of iterative processes.
@@
 Yao et al.
-yao2022react proposed ReAct, which interleaves reasoning traces with task-specific actions, enabling models to interact with external knowledge sources.
+Yao et al.
+(2023) proposed ReAct, which interleaves reasoning traces with task-specific actions, enabling models to interact with external knowledge sources.
 Shinn et al.
-shinn2023reflexion introduced Reflexion, where agents use verbal reinforcement learning to maintain episodic memory and improve decision-making across trials.
+Shinn et al.
+(2023) introduced Reflexion, where agents use verbal reinforcement learning to maintain episodic memory and improve decision-making across trials.
 Park et al.
-park2023generative demonstrated generative agents that simulate believable human behavior through architectures combining memory, reflection and planning.
+Park et al.
+(2023) demonstrated generative agents that simulate believable human behavior through architectures combining memory, reflection and planning.
 These agentic frameworks establish that recursive LLM interactions can produce sophisticated behaviors, but they do not formally characterize the dynamics of such systems in semantic space or analyze convergence and divergence patterns as geometric phenomena.
@@
 Understanding the geometric structure of embedding spaces is crucial for analyzing semantic trajectories.
-Ethayarajh ethayarajh2019contextual systematically studied contextualized representations from BERT, ELMo and GPT-2, demonstrating that embeddings become increasingly anisotropic in upper layers, with vectors concentrating in a narrow cone.
+Ethayarajh Ethayarajh (2019) systematically studied contextualized representations from BERT, ELMo and GPT-2, demonstrating that embeddings become increasingly anisotropic in upper layers, with vectors concentrating in a narrow cone.
 This anisotropy causes high cosine similarities even between semantically distant items, distorting similarity-based measurements.
-Wang and Isola wang2020understanding analyzed contrastive representation learning through the lens of alignment (bringing similar items together) and uniformity (distributing representations across the hypersphere).
+Wang and Isola Wang and Isola (2020) analyzed contrastive representation learning through the lens of alignment (bringing similar items together) and uniformity (distributing representations across the hypersphere).
 They showed that good representations balance these competing objectives, providing theoretical grounding for why pretrained embeddings exhibit systematic geometric biases.
-For calibrating similarity metrics, Niculescu-Mizil and Caruana niculescu2005predicting demonstrated that isotonic regression effectively transforms classifier scores into well-calibrated probabilities.
+For calibrating similarity metrics, Niculescu-Mizil and Caruana Niculescu-Mizil and Caruana (2005) demonstrated that isotonic regression effectively transforms classifier scores into well-calibrated probabilities.
 While originally developed for classification tasks, isotonic regression provides a principled non-parametric method for aligning model outputs with empirical distributions.
@@
 In contrast to existing work, which primarily focuses on static properties of embeddings or isolated transformations, the present work studies the dynamical behavior of iterative LLM transformations as geometric trajectories.
-Prior research on iterative refinement madaan2023selfrefine ; wei2022chain and agentic systems yao2022react ; shinn2023reflexion ; park2023generative has demonstrated impressive capabilities but treats each iteration as an independent step rather than as part of a continuous dynamical process.
-Similarly, embedding geometry research ethayarajh2019contextual ; wang2020understanding characterizes static distributional properties without examining how representations evolve under repeated transformations.
+Prior research on iterative refinement Madaan et al.
+(2023); Wei et al.
+(2022) and agentic systems Yao et al.
+(2023); Shinn et al.
+(2023); Park et al.
+(2023) has demonstrated impressive capabilities but treats each iteration as an independent step rather than as part of a continuous dynamical process.
+Similarly, embedding geometry research Ethayarajh (2019); Wang and Isola (2020) characterizes static distributional properties without examining how representations evolve under repeated transformations.
 Our work bridges these threads by treating agentic loops as discrete dynamical systems in semantic space, enabling the characterization of convergence, divergence and attractor structures through calibrated geometric measurements.
 Contribution of this work.
-While computational creativity frameworks boden2009computermodels ; colton2012computational motivate the need to measure trajectory-based metrics, prior work has not developed the foundational geometric measurement infrastructure required to rigorously analyze iterative transformation dynamics.
-Self-refinement approaches madaan2023selfrefine demonstrate empirical task improvements but do not characterize how representations evolve geometrically through iterations or analyze convergence versus divergence behaviors.
-Agentic frameworks yao2022react ; shinn2023reflexion ; park2023generative establish recursive LLM capabilities but lack quantitative measurement of trajectory dynamics in semantic space.
-Although embedding anisotropy is well-documented ethayarajh2019contextual and theoretical frameworks for representation quality exist wang2020understanding , no existing work provides calibrated, human-aligned similarity metrics for rigorous trajectory analysis in iterative transformations.
+While computational creativity frameworks Boden (2009); Colton and Wiggins (2012) motivate the need to measure trajectory-based metrics, prior work has not developed the foundational geometric measurement infrastructure required to rigorously analyze iterative transformation dynamics.
+Self-refinement approaches Madaan et al.
+(2023) demonstrate empirical task improvements but do not characterize how representations evolve geometrically through iterations or analyze convergence versus divergence behaviors.
+Agentic frameworks Yao et al.
+(2023); Shinn et al.
+(2023); Park et al.
+(2023) establish recursive LLM capabilities but lack quantitative measurement of trajectory dynamics in semantic space.
+Although embedding anisotropy is well-documented Ethayarajh (2019) and theoretical frameworks for representation quality exist Wang and Isola (2020), no existing work provides calibrated, human-aligned similarity metrics for rigorous trajectory analysis in iterative transformations.
 Our contribution builds the foundational geometric measurement infrastructure required for analyzing agentic loop dynamics, providing the tools necessary to eventually address computational creativity questions systematically.
@@
 These geometric constructs open the path to provide to measure Colton & Wiggins’ creativity criteria.
-Second, we introduce isotonic calibration that aligns embedding similarities with human semantic judgments extending calibration methods niculescu2005predicting to overcome anisotropy ethayarajh2019contextual and enable precise geometric measurements using metrics that reflect human-perceived semantic distances.
+Second, we introduce isotonic calibration that aligns embedding similarities with human semantic judgments extending calibration methods Niculescu-Mizil and Caruana (2005) to overcome anisotropy Ethayarajh (2019) and enable precise geometric measurements using metrics that reflect human-perceived semantic distances.
 Third, we provide systematic empirical characterization of dynamical regimes.
@@
 Formally, let $X$ be uniformly distributed on $\mathbb{S}^{d-1}$ and let $f(X)=\langle X,v\rangle$ for a fixed $v\in\mathbb{S}^{d-1}$.
-By Lévy’s concentration inequality (see vershynin2018high for a modern treatment), for any $t>0$,
+By Lévy’s concentration inequality (see Vershynin (2018) for a modern treatment), for any $t>0$,
 $\mathbb{P}\left(\left|f(X)-\mathbb{E}[f(X)]\right|>t\right)\leq 2e^{-cnt^{2}},$ (4)
@@
 This induces a strong anisotropy of the embedding manifold, where most vectors share a large positive cosine with the mean embedding vector.
-As formalized by wang2020understanding , good representation learning requires balancing two properties: alignment (similar points should be close) and uniformity (embeddings should be distributed uniformly on the unit sphere).
+As formalized by Wang and Isola (2020), good representation learning requires balancing two properties: alignment (similar points should be close) and uniformity (embeddings should be distributed uniformly on the unit sphere).
 The observed concentration violates the uniformity property, indicating that the pretrained model prioritizes alignment over uniform distribution.
@@
 $\hat{f}_{h}(x)=\frac{1}{nh\sqrt{2\pi}}\sum_{i=1}^{n}e^{-\frac{1}{2}\left(\frac{x-x_{i}}{h}\right)^{2}},\qquad h=1.06\,\hat{\sigma}\,n^{-1/5}.$ (9)
-Silverman’s rule silverman1986density selects the optimal bandwidth $h$, balancing smoothness and resolution.
+Silverman’s rule Silverman (1986) selects the optimal bandwidth $h$, balancing smoothness and resolution.
 We compute separate densities for human scores and model similarities, $\hat{f}_{H}(x)$ and $\hat{f}_{S}(x)$ and plot them over $[0,1]$.
@@
 The agentic geometric cognition framework introduced in Section 8.3 raises a natural application question: what if the objective that the agent targets is to maximize creativity?
-This section explores how computational creativity, formalized by Boden boden2009computermodels as exploration-transformation-combination processes and by Colton and Wiggins colton2012computational as novelty-value-surprise dimensions, can be operationalized as a specific objective within the geometry of agentic cognition framework.
+This section explores how computational creativity, formalized by Boden Boden (2009) as exploration-transformation-combination processes and by Colton and Wiggins Colton and Wiggins (2012) as novelty-value-surprise dimensions, can be operationalized as a specific objective within the geometry of agentic cognition framework.
 Creativity metrics as geometric observables.
@@
 This operationalization enables measurable creativity control.
-Boden’s exploratory creativity can be triggered within agentic loop iterations by introducing high-temperature sampling and role-based divergent prompts zhao2024creativity (for instance, adopting perspectives such as an artist, a child or a poet), which induce high dispersion and novelty generation.
+Boden’s exploratory creativity can be triggered within agentic loop iterations by introducing high-temperature sampling and role-based divergent prompts Zhao et al.
+(2024) (for instance, adopting perspectives such as an artist, a child or a poet), which induce high dispersion and novelty generation.
 However, unconstrained exploration risks semantic drift and loss of value.

### v4

@@
-Dynamics of Agentic Loops in Large Language Models: A Geometric Theory of Trajectories From Semantic Contraction to Exploratory Divergence
+Geometric Dynamics of Agentic Loops in Large Language Models Trajectories, Attractors and Dynamical Regimes in Semantic Space
 Abstract
-Agentic systems based on large language models operate through recursive feedback loops, where each output becomes the next input in an iterative transformation process.
-Understanding the geometric dynamics of these agentic loops, whether they converge, diverge or exhibit more complex behaviors, remains poorly characterized.
-This paper develops a geometric framework for analyzing agentic loop trajectories in semantic embedding spaces, treating iterative transformations as discrete dynamical systems.
-We formalize the distinction between the artifact space (where linguistic transformations occur) and the semantic embedding space (where geometric measurements are performed).
-Since standard cosine similarities exhibit systematic bias due to embedding anisotropy, we introduce calibrated similarity using isotonic regression, which eliminates bias and improves correlation with human semantic judgments while maintaining high local stability.
-This calibrated metric enables precise geometric analysis through trajectories (temporal sequences of embeddings), clusters (regions of semantic coherence) and attractors (convergence points).
-Experimentally, we characterize two fundamental dynamical regimes through singular agentic loops.
-A contractive loop performing iterative paraphrasing exhibits geometric convergence toward a stable attractor with decreasing dispersion.
-An exploratory loop alternating summarization and negation shows unbounded divergence with no stable clustering structure.
-These regimes demonstrate qualitatively distinct trajectory geometries, contraction versus expansion in semantic space.
-Prompt design directly controls the dynamical regime, enabling systematic engineering of convergent versus divergent behaviors.
-This geometric framework provides the first rigorous quantitative approach to analyzing iterative transformation dynamics in LLMs, with applications to trajectory prediction, stability analysis and controlled composition of complex multi-step transformations.
+Iterative LLM systems—self-refinement, chain-of-thought, autonomous agents—are increasingly deployed, yet their temporal dynamics remain uncharacterized.
+Prior work evaluates task performance at convergence but ignores the trajectory: how does semantic content evolve across iterations?
+Does it stabilize, drift, or oscillate?
+Without answering these questions, we cannot predict system behavior, guarantee stability, or systematically design iterative architectures.
+We formalize agentic loops as discrete dynamical systems in semantic space.
+Borrowing from dynamical systems theory, we define trajectories, attractors and dynamical regimes for recursive LLM transformations, providing rigorous geometric definitions adapted to this setting.
+Our framework reveals that agentic loops exhibit classifiable dynamics: contractive (convergence toward stable semantic attractors), oscillatory (cycling among attractors), or exploratory (unbounded divergence).
+Experiments on singular loops validate the framework.
+Iterative paraphrasing produces contractive dynamics with measurable attractor formation and decreasing dispersion.
+Iterative negation produces exploratory dynamics with no stable structure.
+Crucially, prompt design directly controls the dynamical regime—the same model exhibits fundamentally different geometric behaviors depending solely on the transformation applied.
+This work establishes that iterative LLM dynamics are predictable and controllable, opening new directions for stability analysis, trajectory forecasting, and principled design of composite loops that balance convergence and exploration.
 1 Introduction
-Large language models (LLMs) enable autonomous agentic systems that operate through recursive feedback loops, where each model output becomes the next input in an iterative transformation process.
-These agentic loops appear in diverse applications: iterative refinement systems that progressively improve outputs through self-critique, multi-step reasoning chains that decompose complex problems into intermediate steps and autonomous agents that maintain extended interactions with environments or users.
-Understanding the dynamics of these loops, whether they converge toward stable states, diverge unboundedly or exhibit complex oscillatory behaviors, remains poorly characterized.
-Despite their increasing deployment in reasoning, planning and generation tasks, we lack a rigorous framework for analyzing how iterative transformations behave geometrically in semantic space.
-Characterizing agentic loop dynamics has direct implications across multiple domains.
-For system design, knowing whether a loop exhibits convergent or divergent behavior enables engineers to select appropriate prompts and parameters for specific tasks, stable convergence is desirable for refinement and summarization, while controlled divergence may be useful for exploration and generation.
-For reliability and safety, detecting convergence properties helps identify runaway feedback loops that could produce unbounded drift or oscillation.
-For theoretical understanding, geometric trajectory analysis provides a quantitative lens for studying how language model transformations behave as discrete dynamical systems.
-The ability to predict, measure and control these dynamics is fundamental to deploying reliable iterative generative systems.
-Contributions.
-The main contributions of this work are:
-  1. 1.
-    A geometric framework for analyzing agentic loops.
-We distinguish the artifact space (where text transformations occur) from the semantic embedding space (where geometric measurements are performed) and formalize how transformations induce measurable trajectories, temporal sequences of embeddings that trace the loop’s evolution through semantic space.
-  2. 2.
-    A calibrated similarity metric.
-To overcome the fundamental obstacle that pretrained embeddings exhibit anisotropy, concentrating cosine similarities around uninformative high values, we develop a calibrated metric using isotonic regression on human semantic judgments.
-This calibration eliminates bias and improves correlation while maintaining high local stability (98%), enabling precise trajectory analysis.
-  3. 3.
-    The first systematic empirical characterization of dynamical regimes.
-Through singular agentic loops, we identify and characterize two fundamental regimes: contractive convergence and exploratory divergence.
-We demonstrate that prompt design directly controls the geometric behavior of iterative transformations, establishing that agentic loop dynamics are predictable and controllable through prompt engineering.
-The remainder of the paper is organized as follows.
-Section 2 reviews related work on iterative refinement, agentic systems, dynamical systems and embedding geometry.
-Sections 35 construct the geometric measurement infrastructure: embedding space construction, similarity calibration and stability validation.
-Section 6 formalizes the theoretical framework, defining agentic loops, trajectories, clusters, attractors and dynamical regimes.
-Section 7 presents experimental trajectory studies of contractive and exploratory singular loops, revealing fundamentally distinct geometric behaviors.
-Section 8 discusses implications for composite loops and controlled dynamics.
-Section 9 concludes.
-An earlier version of this work was released as an arXiv preprint.
-2 Related Work
-Computational creativity and the measurement problem.
-The question of implementing computational creativity in iterative systems provides the broader motivation for this work.
-Boden Boden (2009) established the foundational framework for computational creativity, identifying three primary processes: exploration (searching conceptual spaces for new combinations), transformation (modifying the rules or structure of conceptual spaces) and combination (combining ideas from different domains).
-Colton and Wiggins Colton and Wiggins (2012) formalized creativity through measurable dimensions: novelty (semantic or structural distance from existing examples), value (quality, coherence and usefulness) and surprise (unexpectedness relative to prior expectations).
-Zhao et al.
-Zhao et al.
-(2024) recently demonstrated that large language models show creative behaviors across multiple tasks and domains.
-These frameworks raise a fundamental question for iterative LLM systems: How can we engineer agentic loops to implement Boden’s creative processes (explore-transform-combine) in ways that elevate Colton & Wiggins’ creativity metrics (novelty, value, surprise)?
-Answering this question requires the ability to measure similarity-based metrics along trajectories, specifically to quantify semantic distance (for novelty), coherence (for value) and discontinuities (for surprise) as iterative transformations unfold over time.
-However, computational creativity theory provides conceptual frameworks without addressing the foundational measurement infrastructure: how do we rigorously measure semantic similarity and trajectory dynamics in embedding spaces?
-This measurement problem motivates our focus on geometric trajectory analysis as a necessary first step.
-Before we can engineer creative agentic loops or evaluate creativity metrics, we must establish rigorous methods for measuring similarity and characterizing trajectory dynamics.
-Our work addresses this foundational layer, building the geometric and dynamical measurement infrastructure required to eventually tackle computational creativity questions systematically.
-Research on iterative generative systems has explored how large language models behave under recursive transformations, but geometric trajectory analysis of these dynamics remains underdeveloped.
-Beyond the creativity motivation, this work builds upon three main research threads: iterative refinement in language models, agentic systems and multi-step reasoning and embedding geometry.
-Iterative refinement in language models.
-Madaan et al.
-Madaan et al.
-(2023) introduced Self-Refine, demonstrating that large language models can iteratively critique and improve their own outputs through a feedback loop without additional training.
-Their approach achieved substantial improvements across diverse tasks including dialogue generation and mathematical reasoning.
-However, Self-Refine focused on empirical task performance rather than analyzing the geometric structure of how representations evolve through iterations or characterizing convergence properties in semantic space.
-Wei et al.
-Wei et al.
-(2022) showed that chain-of-thought prompting enables large language models to perform complex reasoning by decomposing problems into intermediate steps.
-While this work established that prompting strategies significantly affect reasoning capabilities, it did not examine the stability or trajectory dynamics of iterative processes.
-Agentic systems and multi-step reasoning.
-Recent work has explored how language models can act as autonomous agents through iterative interaction with environments or tools.
-Yao et al.
-Yao et al.
-(2023) proposed ReAct, which interleaves reasoning traces with task-specific actions, enabling models to interact with external knowledge sources.
-Shinn et al.
-Shinn et al.
-(2023) introduced Reflexion, where agents use verbal reinforcement learning to maintain episodic memory and improve decision-making across trials.
-Park et al.
-Park et al.
-(2023) demonstrated generative agents that simulate believable human behavior through architectures combining memory, reflection and planning.
-These agentic frameworks establish that recursive LLM interactions can produce sophisticated behaviors, but they do not formally characterize the dynamics of such systems in semantic space or analyze convergence and divergence patterns as geometric phenomena.
-Embedding geometry and semantic similarity.
-Understanding the geometric structure of embedding spaces is crucial for analyzing semantic trajectories.
-Ethayarajh Ethayarajh (2019) systematically studied contextualized representations from BERT, ELMo and GPT-2, demonstrating that embeddings become increasingly anisotropic in upper layers, with vectors concentrating in a narrow cone.
-This anisotropy causes high cosine similarities even between semantically distant items, distorting similarity-based measurements.
-Wang and Isola Wang and Isola (2020) analyzed contrastive representation learning through the lens of alignment (bringing similar items together) and uniformity (distributing representations across the hypersphere).
-They showed that good representations balance these competing objectives, providing theoretical grounding for why pretrained embeddings exhibit systematic geometric biases.
-For calibrating similarity metrics, Niculescu-Mizil and Caruana Niculescu-Mizil and Caruana (2005) demonstrated that isotonic regression effectively transforms classifier scores into well-calibrated probabilities.
-While originally developed for classification tasks, isotonic regression provides a principled non-parametric method for aligning model outputs with empirical distributions.
-We extend this approach to semantic similarity calibration.
-Positioning of the present work.
-In contrast to existing work, which primarily focuses on static properties of embeddings or isolated transformations, the present work studies the dynamical behavior of iterative LLM transformations as geometric trajectories.
-Prior research on iterative refinement Madaan et al.
-(2023); Wei et al.
-(2022) and agentic systems Yao et al.
+Iterative LLM systems are now ubiquitous.
+Self-Refine Madaan et al.
+(2023) iteratively critiques and improves outputs.
+Chain-of-thought Wei et al.
+(2022) decomposes reasoning into sequential steps.
+Autonomous agents Yao et al.
 (2023); Shinn et al.
 (2023); Park et al.
-(2023) has demonstrated impressive capabilities but treats each iteration as an independent step rather than as part of a continuous dynamical process.
-Similarly, embedding geometry research Ethayarajh (2019); Wang and Isola (2020) characterizes static distributional properties without examining how representations evolve under repeated transformations.
-Our work bridges these threads by treating agentic loops as discrete dynamical systems in semantic space, enabling the characterization of convergence, divergence and attractor structures through calibrated geometric measurements.
-Contribution of this work.
-While computational creativity frameworks Boden (2009); Colton and Wiggins (2012) motivate the need to measure trajectory-based metrics, prior work has not developed the foundational geometric measurement infrastructure required to rigorously analyze iterative transformation dynamics.
-Self-refinement approaches Madaan et al.
-(2023) demonstrate empirical task improvements but do not characterize how representations evolve geometrically through iterations or analyze convergence versus divergence behaviors.
+(2023) maintain extended recursive interactions.
+These systems share a common structure: agentic loops where each output becomes the next input, $a_{t+1}=F(a_{t})$.
+Yet despite widespread deployment, a fundamental question remains unanswered: what are the dynamics of these loops?
+Do they converge to stable outputs?
+Drift unboundedly?
+Oscillate between states?
+Current work evaluates only endpoints—task accuracy after $n$ iterations—while ignoring the trajectory that produced them.
+This gap is consequential: without understanding dynamics, we cannot predict when loops will stabilize, detect runaway divergence, or design loops with desired convergence properties.
+The gap.
+Existing research treats iterations as independent steps rather than as a dynamical process.
+Self-Refine Madaan et al.
+(2023) measures task improvement but not geometric trajectory structure.
 Agentic frameworks Yao et al.
 (2023); Shinn et al.
-(2023); Park et al.
-(2023) establish recursive LLM capabilities but lack quantitative measurement of trajectory dynamics in semantic space.
-Although embedding anisotropy is well-documented Ethayarajh (2019) and theoretical frameworks for representation quality exist Wang and Isola (2020), no existing work provides calibrated, human-aligned similarity metrics for rigorous trajectory analysis in iterative transformations.
-Our contribution builds the foundational geometric measurement infrastructure required for analyzing agentic loop dynamics, providing the tools necessary to eventually address computational creativity questions systematically.
-We address these gaps through three main innovations.
-First, we develop a geometric framework operating in the representation space treating iterative transformations as discrete dynamical systems.
-This framework captures measurable trajectories, clusters, attractors and dynamical regimes that quantify how semantic representations evolve under recursive application.
-These geometric constructs open the path to provide to measure Colton & Wiggins’ creativity criteria.
-Second, we introduce isotonic calibration that aligns embedding similarities with human semantic judgments extending calibration methods Niculescu-Mizil and Caruana (2005) to overcome anisotropy Ethayarajh (2019) and enable precise geometric measurements using metrics that reflect human-perceived semantic distances.
-Third, we provide systematic empirical characterization of dynamical regimes.
-We demonstrate that prompt design directly determines whether agentic loops exhibit convergence, divergence or complex behaviors, as revealed by measurable geometric indicators in calibrated semantic space.
-This establishes the foundation for engineering Boden’s creative processes through controlled trajectory dynamics.
-3 Semantic Space Construction
-3.1 Artifact Space
-Before defining how we measure semantic dynamics, we must first specify what is being measured.
-In an agentic loop, the fundamental objects of transformation are artifacts: outputs generated and processed by a large language model.
-While artifacts can take various modalities (text, images, audio, code, etc.), in this paper we focus exclusively on textual artifacts.
-We define the artifact space $\mathcal{A}$ as the set of all possible text strings that can be generated by the language model. Each element $a\in\mathcal{A}$ represents a complete textual output.
-While $\mathcal{A}$ is finite (bounded by the model’s vocabulary and maximum context length), it is combinatorially large and lacks inherent geometric structure suitable for quantitative analysis of convergence, divergence or semantic drift.
-To measure such dynamics, we construct a continuous representation space through embedding, as described in the following subsection.
-3.2 Representation Space and Embedding Function
-To enable quantitative analysis of semantic dynamics, we construct a continuous representation space $\mathcal{E}$ by embedding textual artifacts into a geometric space where distances and similarities can be measured.
-Let $\phi:\mathcal{A}\to\mathbb{R}^{d}$ denote an embedding function that maps each artifact $a\in\mathcal{A}$ to a dense vector representation $\phi(a)\in\mathbb{R}^{d}$ in a $d$-dimensional Euclidean space.
-To ensure that similarity measurements depend solely on directional alignment (not magnitude), we apply $\ell_{2}$-normalization to all embeddings.
-Let $\psi:\mathcal{A}\to\mathbb{S}^{d-1}$ denote the normalized embedding function defined as:
-$\psi(a)=\frac{\phi(a)}{\|\phi(a)\|_{2}},$ (1)
-where $\mathbb{S}^{d-1}=\{e\in\mathbb{R}^{d}:\|e\|_{2}=1\}$ is the (d1)(d-1)-dimensional unit hypersphere. We define the representation space as:
-$\mathcal{E}=\left\{\psi(a)\mid a\in\mathcal{A}\right\}\subset\mathbb{S}^{d-1}.$ (2)
-Each artifact $a$ is thus represented by a unit vector $e=\psi(a)\in\mathcal{E}$.
-Model-agnostic framework.
-The framework presented above is embedding-model agnostic: any sentence encoder that maps text to dense vectors can instantiate $\phi$ and our analysis of trajectories, clusters and dynamical regimes applies regardless of the specific model choice.
-Implementation choice.
-For this work, we instantiate $\phi$ using the Xenova/paraphrase-mpnet-base-v2 model, a pretrained transformer encoder with $d=768$ dimensions. This model was chosen for three practical reasons:
-  1. 1.
-    It is trained on sentence pairs with human-labeled paraphrase relations, ensuring that geometric proximity corresponds to semantic similarity.
-  2. 2.
-    It offers a balanced trade-off between expressiveness and dimensionality, allowing fine-grained distinctions while remaining computationally efficient.
-  3. 3.
-    The Xenova implementation provides a reproducible and locally executable inference pipeline, facilitating controlled experimentation without external API dependencies.
-The normalized representation space $\mathcal{E}\subset\mathbb{S}^{767}$ thus forms the manifold on which all further measurements (distances, similarities and clustering) are defined.
-3.3 Similarity Metric
-Within the representation space $\mathcal{E}$, each artifact $a\in\mathcal{A}$ is represented by its normalized embedding $e=\psi(a)\in\mathcal{E}\subset\mathbb{S}^{d-1}$.
-We define the model-based similarity function as
-$s^{(m)}:\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\to[-1,1],\qquad s^{(m)}(e_{1},e_{2})=\langle e_{1},e_{2}\rangle.$ (3)
-Since all elements of $\mathcal{E}$ are unit vectors (i.e., $\|e\|_{2}=1$ for all $e\in\mathcal{E}$), the inner product $\langle e_{1},e_{2}\rangle$ directly equals the cosine of the angle between them, providing a normalized measure of semantic proximity.
-In the following subsections, we first characterize the theoretical distribution of similarities in high-dimensional spaces (§3.4), then evaluate how well $s^{(m)}$ aligns with human semantic judgments through quantitative metrics (§3.5) and visualization (§3.7).
-3.4 High-Dimensional Similarity Distribution
-On a standard high-dimensional unit sphere $\mathbb{S}^{d-1}$, two random vectors $u,v$ drawn uniformly from the surface tend to be nearly orthogonal as the dimension increases.
-Formally, let $X$ be uniformly distributed on $\mathbb{S}^{d-1}$ and let $f(X)=\langle X,v\rangle$ for a fixed $v\in\mathbb{S}^{d-1}$.
-By Lévy’s concentration inequality (see Vershynin (2018) for a modern treatment), for any $t>0$,
-$\mathbb{P}\left(\left|f(X)-\mathbb{E}[f(X)]\right|>t\right)\leq 2e^{-cnt^{2}},$ (4)
-where $c>0$ is an absolute constant and n=d1n=d-1 is the dimension of the sphere.
-Since $\mathbb{E}[f(X)]=0$ for the uniform distribution, the inner product $\langle u,v\rangle$ concentrates exponentially fast around 0 as $d\to\infty$.
-In such isotropic settings, the similarity distribution is symmetric and sharply peaked near zero.
-However, the semantic embedding space induced by a pretrained language model such as paraphrase-mpnet-base-v2 diverges from this theoretical isotropy.
-Because the model was trained on non-uniform pairs of semantically related sentences, the embeddings are not uniformly distributed over the hypersphere but rather cluster around a dominant mean direction in $\mathbb{R}^{d}$.
-This induces a strong anisotropy of the embedding manifold, where most vectors share a large positive cosine with the mean embedding vector.
-As formalized by Wang and Isola (2020), good representation learning requires balancing two properties: alignment (similar points should be close) and uniformity (embeddings should be distributed uniformly on the unit sphere).
-The observed concentration violates the uniformity property, indicating that the pretrained model prioritizes alignment over uniform distribution.
-This trade-off is typical of contrastively trained embeddings.
-Empirically, the distribution of $s^{(m)}(e_{1},e_{2})$ is concentrated around $0.8$, in stark contrast with the isotropic baseline centered at 0.
-3.5 Quantitative Metrics for Model-Human Alignment
-To quantitatively assess how well $s^{(m)}$ aligns with human semantic judgments, we use the MTEB STS Benchmark (STS-train) dataset.
-This dataset contains pairs of textual artifacts $(a_{1},a_{2})\in\mathcal{A}\times\mathcal{A}$ annotated with normalized human similarity scores $s^{(h)}(a_{1},a_{2})\in[0,1]$, rescaled from the original 0–5 range.
-For each pair $(a_{1},a_{2})$, we compute their normalized embeddings $e_{1}=\psi(a_{1})$ and $e_{2}=\psi(a_{2})$ in the representation space $\mathcal{E}$.
-We then evaluate the model-based similarity $s^{(m)}(e_{1},e_{2})$ and compare it to the human judgment $s^{(h)}(a_{1},a_{2})$.
-We evaluate their alignment using complementary metrics, each capturing a distinct aspect of correlation or calibration:
-  • Root Mean Squared Error (RMSE):
-    $\mathrm{RMSE}=\sqrt{\frac{1}{N}\sum_{k=1}^{N}\big(s^{(m)}(e_{1}^{k},e_{2}^{k})-s^{(h)}(e_{1}^{k},e_{2}^{k})\big)^{2}}\in[0,1].$
-    Measures the average magnitude of deviation between model and human similarities, penalizing large discrepancies more strongly.
-Lower RMSE indicates better agreement (RMSE = 0 means perfect match; RMSE = 1 indicates maximum deviation).
-  • Mean Bias Error (MBE):
-    $\mathrm{MBE}=\frac{1}{N}\sum_{k=1}^{N}\big(s^{(m)}(e_{1}^{k},e_{2}^{k})-s^{(h)}(e_{1}^{k},e_{2}^{k})\big)\in[-1,1].$
-    Quantifies systematic bias: positive when the model tends to overestimate human similarity, negative when it underestimates.
-MBE = 0 indicates no systematic bias; $|\mathrm{MBE}|$ close to 1 indicates severe consistent over- or under-estimation.
-  • Expected Calibration Error (ECE): This metric evaluates the degree to which model similarity scores are statistically calibrated to human perception.
-We partition the similarity range $[0,1]$ into $B$ bins, where $B_{b}$ denotes the set of sample pairs in bin $b$.
-For each bin, we compute:
-    • $\mathrm{acc}(B_{b})$: the average human similarity score (accuracy, i.e., ground truth) within that bin
-    • $\mathrm{conf}(B_{b})$: the average model similarity score (confidence, i.e., predicted value) within that bin
-    The ECE is then computed as the weighted average absolute deviation:
-    $\mathrm{ECE}=\sum_{b=1}^{B}\frac{|B_{b}|}{N}\,\big|\,\mathrm{acc}(B_{b})-\mathrm{conf}(B_{b})\,\big|\in[0,1].$
-    Lower ECE values indicate better calibration (ECE = 0 means perfect calibration; ECE = 1 indicates maximum miscalibration).
-  • Pearson Correlation Coefficient ($r$):
-    $r=\frac{\mathrm{Cov}(s^{(m)},s^{(h)})}{\sigma_{(m)}\,\sigma_{(h)}}\in[-1,1],$
-    where $\mathrm{Cov}(s^{(m)},s^{(h)})$ is the covariance between model and human similarity scores, $\sigma_{(m)}$ is the standard deviation of model scores and $\sigma_{(h)}$ is the standard deviation of human scores.
-Explicitly, this can be computed as:
-    $r=\frac{\sum_{k=1}^{N}(s^{(m)}_{k}-\bar{s}^{(m)})(s^{(h)}_{k}-\bar{s}^{(h)})}{\sqrt{\sum_{k=1}^{N}(s^{(m)}_{k}-\bar{s}^{(m)})^{2}}\;\sqrt{\sum_{k=1}^{N}(s^{(h)}_{k}-\bar{s}^{(h)})^{2}}},$
-    where $\bar{s}^{(m)}$ is the mean of model scores, $\bar{s}^{(h)}$ is the mean of human scores and $N$ is the total number of pairs.
-This metric captures the strength of linear dependence between model and human similarities.
-$r=1$ indicates perfect positive linear correlation, r=1r=-1 indicates perfect negative linear correlation and $r=0$ indicates no linear correlation.
-Values closer to $\pm 1$ indicate stronger linear alignment.
-  • Spearman Rank Correlation ($\rho$):
-    $\rho=r_{\text{Pearson}}(\mathrm{rank}(s^{(m)}),\mathrm{rank}(s^{(h)}))\in[-1,1],$
-    where $\mathrm{rank}(s^{(m)})$ assigns each model similarity score its position when all $N$ scores are sorted in ascending order and similarly for $\mathrm{rank}(s^{(h)})$.
-The function $r_{\text{Pearson}}(\cdot,\cdot)$ denotes the Pearson correlation coefficient applied to these rank vectors.
-This metric measures the monotonic alignment between the two orderings, independently of scale.
-$\rho=1$ indicates perfect monotonic agreement (rankings perfectly aligned), $\rho=-1$ indicates perfect monotonic disagreement (rankings perfectly reversed) and $\rho=0$ indicates no monotonic relationship.
-Values closer to $\pm 1$ indicate stronger monotonic alignment.
-Quantitatively, the results obtained on the STS-train dataset are:
-$\mathrm{RMSE}=0.1702,\quad\mathrm{MBE}=0.0789,\quad\mathrm{ECE}=0.0797,\quad r=0.8576,\quad\rho=0.8430.$ (5)
-These metrics reveal several key properties of the embedding space.
-The RMSE of 0.17 indicates a moderate average deviation between model and human similarities, suggesting reasonable overall agreement with some residual discrepancies.
-The positive MBE of 0.079 shows a systematic tendency for the model to overestimate semantic similarity compared to human judgments.
-This is a characteristic feature of dense embedding spaces where related concepts cluster more tightly than human perception suggests.
-The ECE of 0.080 quantifies the calibration gap: the model’s similarity scores are not perfectly aligned with the empirical frequency of human agreement, reflecting the need for recalibration to achieve probabilistic consistency.
-The strong correlation values ($r=0.86$ and $\rho=0.84$) confirm that despite these biases, the model captures the underlying rank order and linear structure of human semantic similarity judgments.
-Together, these results indicate that the embedding space preserves semantic relationships effectively while exhibiting systematic geometric biases, specifically anisotropic concentration of embeddings, that lead to consistent overestimation and miscalibration.
-3.6 High-Confidence Similarity Threshold
-To obtain a reliable decision boundary for semantic similarity, we introduce the High-Confidence Similarity Threshold $\tau_{\text{HCS}}(s)$, a data-driven value derived from human-judged highly similar pairs.
-This threshold provides a statistically principled criterion for determining when two sentences can be considered semantically similar with high confidence.
-Importantly, this concept is independent of any particular similarity measure and can be applied to any similarity function.
-Definition. For a given similarity function $s$ (such as the raw cosine similarity $s^{(m)}$ or any other similarity function), the threshold is computed as:
-$\tau_{\text{HCS}}(s)=Q_{0.05}(s\mid s^{(h)}>0.9)$ (6)
-that is, the 5% quantile of the similarity scores $s$ among sentence pairs with high human similarity ($s^{(h)}>0.9$).
-The confidence level $\alpha=0.05$ is a standard choice, though it can be adjusted based on application requirements.
-Interpretation.
-This definition is meaningful for three key reasons.
-First, the threshold is grounded in human judgments of semantic similarity rather than arbitrary choice.
-Second, it captures the lower bound of similarity scores for similar sentences as judged by humans.
-Third, it provides a direct probabilistic guarantee: at least 95% of truly similar pairs (those with $s^{(h)}>0.9$) lie above this threshold.
-Statistical justification. By construction, the threshold satisfies:
-$\mathbb{P}(s\geq\tau_{\text{HCS}}(s)\mid s^{(h)}>0.9)\geq 0.95$ (7)
-This conditional probability allows us to make reliable decisions: if a pair exceeds the threshold, it is likely to belong to the high-similarity group as defined by human raters.
-Decision criterion. Any pair with similarity $s\geq\tau_{\text{HCS}}(s)$ can be considered semantically similar with high confidence.
-For the raw cosine similarity, we denote:
-$\tau_{\text{HCS}}^{(m)}=\tau_{\text{HCS}}(s^{(m)})\approx 0.72$ (8)
-3.7 Visualization of Similarity Distributions
-Beyond these numerical indicators, we visualize the distributions and relationships between $s^{(m)}$ and $s^{(h)}$ using density estimation techniques.
-To compare model similarities with human similarity judgments, we estimate their respective continuous distributions using Kernel Density Estimation (KDE).
-Each observed value contributes a Gaussian “bump” to the overall density, producing a smooth, non-parametric approximation of the underlying probability function:
-$\hat{f}_{h}(x)=\frac{1}{nh\sqrt{2\pi}}\sum_{i=1}^{n}e^{-\frac{1}{2}\left(\frac{x-x_{i}}{h}\right)^{2}},\qquad h=1.06\,\hat{\sigma}\,n^{-1/5}.$ (9)
-Silverman’s rule Silverman (1986) selects the optimal bandwidth $h$, balancing smoothness and resolution.
-We compute separate densities for human scores and model similarities, $\hat{f}_{H}(x)$ and $\hat{f}_{S}(x)$ and plot them over $[0,1]$.
-Alignment between both curves indicates good agreement, while systematic shifts reveal under- or over-confidence in the model’s similarity predictions.
-Additionally, both visualizations include a vertical green line marking the high-confidence similarity threshold $\tau_{\text{HCS}}^{(m)}$ (defined in §3.6), which serves as a data-driven decision boundary for determining when two sentences can be considered semantically similar with high confidence.
-In addition to that, we also compute joint density heatmaps.
-To evaluate the relationship between human similarity judgments and model predictions, we estimate their joint density over the $0,1]^{2}$ space.
-The domain is divided into a regular grid of bins and each cell’s density $\rho_{ij}=N_{ij}/(n\,\Delta x\,\Delta y)$ reflects the concentration of pairs $(h_{i},s_{i})$.
-A 2D Gaussian smoothing kernel can be applied to obtain a continuous density field $\tilde{\rho}(x,y)$, from which iso-density contours and a color intensity map are derived.
-The diagonal $y=x$ represents perfect alignment, where model similarities equal human scores.
-High density concentrated along this diagonal indicates strong agreement, while systematic deviations above or below it reveal over- or under-confidence and vertical dispersion signals inconsistency in the model’s predictions.
-(a) Density plot (KDE)
-(b) Heatmap (binned)
-Figure 1: Original (uncalibrated) cosine similarity versus human similarity scores.
-Figure 1(a) reveals a striking concentration of model similarities around $0.8$, contrasting sharply with the broader distribution of human scores across $[0,1]$.
-This concentration reflects the anisotropic geometry of the embedding space, where semantically related sentences cluster tightly on the unit hypersphere, violating the theoretical isotropy assumption.
-The green vertical line at $\tau_{\text{HCS}}^{(m)}$ marks the decision boundary derived from high-confidence human judgments, positioned within this concentration region, indicating that most model similarities for genuinely similar pairs fall within this narrow band.
-Human judgments, by contrast, exhibit greater variability, with substantial density in the low-to-moderate similarity range ($[0.2,0.6]$), indicating more nuanced perceptual distinctions.
-Figure 1(b) visualizes the joint density $\tilde{\rho}(s^{(h)},s^{(m)})$ as a 2D heatmap.
-The density mass lies predominantly above the diagonal $y=x$, confirming systematic overestimation: the model assigns higher similarity scores than humans across most pairs.
-The vertical spread for human scores in the range $[0.3,0.7]$ indicates high variance in model predictions for pairs with intermediate human similarity: the model struggles to distinguish subtle semantic differences, compressing them into a narrow band around $0.8$.
-In contrast, pairs with very low human similarity ($s^{(h)}<0.3$) show tighter correspondence, suggesting the model correctly identifies dissimilar pairs but struggles to capture gradations of relatedness.
-These visualizations expose a fundamental calibration gap: the model’s cosine similarity $s^{(m)}(e_{1},e_{2})$ does not align with human semantic judgments $s^{(h)}(e_{1},e_{2})$, both in terms of distribution shape (concentration vs.
-spread) and pointwise correspondence (systematic overestimation).
-This raises a natural question: Can we recalibrate or transform the model’s similarity function to improve its correlation with human semantic judgments?
-This question is investigated in the following section.
-4 Calibration of Semantic Similarity
-The visualizations in Section 3.7 reveal systematic bias in the model’s similarity predictions: $s^{(m)}$ consistently overestimates human judgments $s^{(h)}$ and exhibits compressed variance.
-Calibration provides a post-processing mechanism to correct these distributional and pointwise misalignments.
-Formally, calibration seeks a transformation function $g:[-1,1]\to\mathbb{R}$ such that the calibrated similarity
-$\tilde{s}(e_{1},e_{2})=g\big(s^{(m)}(e_{1},e_{2})\big)$ (10)
-better aligns with human similarity judgments $s^{(h)}(e_{1},e_{2})$.
-4.1 Overview of Calibration Methods
-We evaluate several calibration functions that map cosine similarity to human-aligned scores, each varying in parametric flexibility and nonlinearity.
-Linear regression fits a simple affine transformation $\tilde{s}(x)=ax+b$, providing a global scaling and shift but unable to capture nonlinear patterns (Figure 2).
-Isotonic regression learns a monotonic, piecewise-constant mapping that adapts locally to the data without assuming a functional form, allowing flexible nonlinear calibration while preserving rank ordering (Figure 3).
-Sigmoid mapping applies a logistic curve $\tilde{s}(x)=1/(1+e^{-a(x-b)})$, modeling smooth saturation at similarity extremes (Figure 4).
-Polynomial transformations (degrees 2, 3 and 4) fit $\tilde{s}(x)=\sum_{i=0}^{n}a_{i}x^{i}$, offering increasing flexibility to model curvature, though higher degrees may introduce unwanted oscillations (Figures 5, 6, 7).
-Beta distribution calibration uses a beta CDF parameterized by shape $\alpha$ and $\beta$, suited for bounded, skewed distributions but prone to instability when assumptions mismatch the data (Figure 8).
-(a) Density plot (KDE)
-(b) Heatmap (binned)
-Figure 2: Linear regression calibration.
-(a) Density plot (KDE)
-(b) Heatmap (binned)
-Figure 3: Isotonic regression calibration.
-(a) Density plot (KDE)
-(b) Heatmap (binned)
-Figure 4: Sigmoid calibration.
-(a) Density plot (KDE)
-(b) Heatmap (binned)
-Figure 5: Polynomial degree 2 calibration.
-(a) Density plot (KDE)
-(b) Heatmap (binned)
-Figure 6: Polynomial degree 3 calibration.
-(a) Density plot (KDE)
-(b) Heatmap (binned)
-Figure 7: Polynomial degree 4 calibration.
-(a) Density plot (KDE)
-(b) Heatmap (binned)
-Figure 8: Beta distribution calibration.
-4.2 Results
-Table 1 presents the performance comparison of all calibration methods against the original (uncalibrated) cosine similarity using correlation metrics.
-Percentage changes relative to the original baseline are shown for reference.
-Method RMSE MBE ECE Pearson $r$ Spearman $\rho$
-Original 0.1702 0.0789 0.0797 0.8576 0.8430
-Linear 0.1506 $\downarrow$11.5% 0.0000 $\downarrow$100.0% 0.0222 $\downarrow$72.1% 0.8576 $\pm$0.0% 0.8430 $\pm$0.0%
-Isotonic 0.1411 $\downarrow$17.1% 0.0000 $\downarrow$100.0% 0.0000 $\downarrow$99.9% 0.8764 $\uparrow$2.2% 0.8563 $\uparrow$1.6%
-Sigmoid 0.3192 $\uparrow$87.6% 0.2604 $\uparrow$230.2% 0.2604 $\uparrow$226.7% 0.8182 $\downarrow$4.6% 0.8430 $\pm$0.0%
-Beta 0.5168 $\uparrow$203.6% 0.3011-0.3011 $\uparrow$281.8% 0.4243 $\uparrow$432.3% 0.8266-0.8266 $\downarrow$196.4% 0.8246-0.8246 $\downarrow$197.8%
-Polynomial-2 0.1474 $\downarrow$13.4% 0.0004 $\downarrow$99.5% 0.0095 $\downarrow$88.1% 0.8641 $\uparrow$0.8% 0.8430 $\pm$0.0%
-Polynomial-3 0.1472 $\downarrow$13.5% 0.0000 $\downarrow$100.0% 0.0057 $\downarrow$92.8% 0.8644 $\uparrow$0.8% 0.8430 $\pm$0.0%
-Polynomial-4 0.1472 $\downarrow$13.5% 0.0000-0.0000 $\downarrow$100.0% 0.0058 $\downarrow$92.7% 0.8645 $\uparrow$0.8% 0.8430 $\pm$0.0%
-Table 1: Performance comparison of calibration methods with percentage changes relative to the original baseline.
-For RMSE, MBE and ECE, lower is better (green $\downarrow$); for Pearson and Spearman, higher is better (green $\uparrow$).
-Isotonic regression achieves the strongest monotonic correlation (Spearman $\rho=0.8563$) and lowest calibration error (ECE =0.0000=0.0000), with perfect elimination of mean bias (MBE =0.0000=0.0000) and a 17.1% reduction in RMSE compared to the original baseline.
-This defines our calibrated similarity function
-$\tilde{s}:\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\to\mathbb{R},\qquad\tilde{s}=f_{\mathrm{isotonic}}\circ s^{(m)}$ (11)
-used throughout the dynamic analysis. In practice, since we align with human similarity scores in $[0,1]$, calibrated outputs are clamped to this interval.
-Similarly, we define the high-confidence similarity threshold for the calibrated similarity:
-$\tilde{\tau}_{\text{HCS}}=\tau_{\text{HCS}}(\tilde{s})\approx 0.65$ (12)
-5 Evaluating Local Stability of the Semantic Space
-5.1 Motivation
-Even if the calibrated semantic similarity aligns well with human judgments, it remains essential to evaluate the local stability of the embedding function $\psi:\mathcal{A}\to\mathcal{E}$ itself.
-A locally stable embedding function should preserve high similarity under small linguistic perturbations: when two texts differ only slightly in form or syntax, their semantic similarity should remain consistently high.
-However, $\psi$ is learned and discretized; neither its continuity nor its Lipschitz properties can be guaranteed theoretically.
-Our first objective is therefore empirical: to determine whether $\psi$ behaves as locally stable in practice.
-However, a critical methodological concern arises from our use of isotonic regression for calibration.
-While isotonic regression optimally preserves monotonicity and eliminates bias, it produces a piecewise constant (and therefore discontinuous) mapping $f_{\text{isotonic}}$.
-This raises an essential question: does the discontinuous nature of isotonic regression degrade the local stability observed in the raw embedding space?
-To address both concerns, we conduct a dual analysis of stability:
-  1. 1.
-    Base model stability, evaluated on raw cosine similarity $s^{(m)}$, measures how the uncalibrated embedding function reacts to small linguistic perturbations.
-  2. 2.
-    Calibrated stability, evaluated on the calibrated similarity $\tilde{s}$, measures whether isotonic calibration preserves or degrades this local stability.
-Comparing these two perspectives allows us to validate that calibration does not introduce artificial instabilities, ensuring that the semantic trajectories observed in agentic loops reflect genuine semantic dynamics rather than artifacts of the calibration process.
-5.2 Artifact Perturbation Definition
-To evaluate local stability, we constructed a dedicated dataset using ChatGPT to generate pairs of sentences differing by minimal and controlled linguistic perturbations.
-Each pair $(a_{i},a_{i}^{\prime})$ was manually categorized into one of seven perturbation types, each intended to simulate a semantically minimal transformation while introducing specific syntactic or lexical variations.
-This approach allows for systematic testing of $\psi$’s local stability across controlled linguistic dimensions.
-The following perturbation types were included, each illustrated with examples from the dataset:
-  • DETERMINER_VARIATION: A minimal change affecting only the determiner (the, a, this, that) without modifying the tense, structure or semantic content of the sentence.
-Example: “The cat sleeps on the sofa.” $\to$ “A cat sleeps on the sofa.”
-  • TENSE_VARIATION: A transformation that shifts the verbal tense while preserving the event described by the sentence. Example: “She drinks tea in the morning.” $\to$ “She drank tea in the morning.”
-  • SYNONYM_SUBSTITUTION: Replacing a lexical item (noun, verb, adjective) with a close synonym that preserves the meaning.
-Example: “He reads a book every night.” $\to$ “He reads a novel every night.”
-  • LOGICAL_PARAPHRASE: A reformulation that preserves the logical meaning of the sentence, often through active/passive alternation or structural inversion.
-Example: “The teacher explains the lesson.” $\to$ “The lesson is explained by the teacher.”
-  • NOMINALIZATION: Transforming a predicate into its nominal form (or the reverse) while keeping the meaning equivalent.
-Example: “He decided to leave early.” $\to$ “He made a decision to leave early.”
-  • COREFERENCE_EXPANSION: Replacing a pronoun with its explicit referent (or vice versa), keeping the propositional content unchanged. Example: “Alice lost her keys.” $\to$ “Alice lost Alice’s keys.”
-  • QUANTIFIER_VARIATION: Replacing a quantifier with another logically equivalent form (all $\leftrightarrow$ every), without changing the truth conditions.
-Example: “All cats like to sleep.” $\to$ “Every cat likes to sleep.”
-The dataset comprises 700 pairs across all categories (100 pairs per type), covering both lexical and syntactic variations.
-This controlled diversity allows us to examine how different perturbation types affect similarity in both the raw and calibrated embedding spaces.
-5.3 Metrics and Methodology
-For each pair $(a_{i},a_{i}^{\prime})$, we compute their normalized embeddings $e_{i}=\psi(a_{i})$ and $e_{i}^{\prime}=\psi(a_{i}^{\prime})$ within $\mathbb{S}^{d-1}$.
-We then measure stability using two complementary similarity functions:
-Base model similarity.
-We compute the raw cosine similarity:
-$s_{i}^{(m)}=s^{(m)}(e_{i},e_{i}^{\prime})=\langle e_{i},e_{i}^{\prime}\rangle,$ (13)
-which measures the uncalibrated similarity between perturbed pairs.
+(2023) demonstrate sophisticated behaviors without characterizing convergence or divergence patterns.
+Embedding geometry research Ethayarajh (2019); Wang and Isola (2020) analyzes static distributional properties but not how representations evolve under repeated transformations.
+Contributions.
+We formalize agentic loops as discrete dynamical systems in semantic space, drawing on classical theory Strogatz (2018). Our contributions are:
+  1. (i)
+    Definitions.
+Operational concepts of trajectory (the sequence of semantic embeddings as the loop evolves), attractor (stable regions detected via incremental clustering), and dynamical regime (contractive, oscillatory, or exploratory).
+These are tailored to LLM transformations, not generic dynamical systems theory.
+  2. (ii)
+    Measurement protocol.
+A concrete pipeline from text to calibrated embedding to geometric indicators.
+We define and compute local drift (similarity to previous iteration), global drift (similarity to initial state), dispersion, and cluster persistence—all using calibrated similarity Tacheny (2025) to ensure measurements align with human semantic judgments.
+  3. (iii)
+    Control insight.
+Empirical demonstration that prompt design determines the dynamical regime.
+On the same LLM, iterative paraphrasing produces contractive dynamics (similarity increases, dispersion decreases, stable attractors form), while iterative negation produces exploratory dynamics (similarity fluctuates, no stable structure emerges).
+The transformation operator, not the model, controls the regime.
+Why it matters.
+These contributions enable capabilities absent from prior work: stability analysis (detect convergence/divergence from early iterations), trajectory prediction (forecast future semantic states), and principled loop design (engineer composite loops with controlled dynamics).
+For computational creativity Boden (2009); Colton and Wiggins (2012), this provides infrastructure to balance exploration (novelty) and consolidation (value preservation).
+Scope.
+For improved interpretability, we use a calibrated semantic similarity Tacheny (2025) where measurements align with human judgments.
+However, our definitions and qualitative results (contraction vs.
+divergence, attractor detection) are order-based and thus robust to any monotonic transformation of the similarity metric, including raw cosine similarity.
+2 Related Work
+2.1 Iterative LLM Systems and Agentic Loops
+A growing body of work studies iterative and recursive LLM architectures.
+Self-Refine Madaan et al.
+(2023) demonstrated that LLMs can iteratively improve outputs through self-critique, measuring task accuracy gains across iterations.
+Chain-of-thought prompting Wei et al.
+(2022) decomposes reasoning into sequential steps, evaluated by final answer correctness.
+ReAct Yao et al.
+(2023) interleaves reasoning with actions; Reflexion Shinn et al.
+(2023) uses verbal reinforcement for multi-trial improvement; generative agents Park et al.
+(2023) simulate human behavior through memory and reflection.
+What these works measure: task performance, reward, success rate, output quality—all evaluated at the endpoint of the iterative process.
+What they do not measure: the trajectory itself.
+How does semantic content evolve across iterations?
+Is the loop converging, oscillating, or drifting?
+At what rate?
+These questions are not addressed because no geometric measurement framework exists for this purpose.
+2.2 Dynamical Systems and Stability Analysis
+Discrete dynamical systems theory Strogatz (2018) provides mature tools for analyzing iterations $x_{t+1}=F(x_{t})$: fixed points, periodic orbits, attractors, Lyapunov stability, contraction mappings.
+These concepts are well-studied in control theory, optimization, and numerical analysis, where operators act on well-defined vector spaces with known analytic properties.
+Why this does not directly transfer: LLM transformations differ fundamentally from classical settings.
+The operator $F$ is a deep neural network with no closed-form expression.
+The state space is not the embedding space itself but a discrete artifact space (text), which is only projected into embeddings post-hoc for measurement.
+The geometry is indirect, high-dimensional, and noisy.
+Classical stability theorems assume conditions (Lipschitz bounds, differentiability) that cannot be verified for prompt-driven text transformations.
+Consequently, while the vocabulary of dynamical systems (trajectory, attractor, regime) is applicable, the measurement problem is entirely new: how do we quantify these concepts when the operator is a black-box LLM and the state space is linguistic?
+2.3 Embedding Geometry
+Research on embedding geometry Ethayarajh (2019); Wang and Isola (2020) has characterized static properties of representation spaces: anisotropy, clustering, alignment-uniformity trade-offs.
+This work analyzes the distribution of embeddings but not how representations evolve under repeated transformations.
+The temporal dimension—trajectories through embedding space—remains unexplored.
+2.4 Gap and Positioning
+The gap is now clear:
+  • Iterative LLM systems exist and are widely deployed, but they are evaluated by outcomes, not by the geometry of the transformation process.
+  • Dynamical systems theory provides the conceptual vocabulary, but its tools assume conditions that do not hold for LLM transformations.
+  • Embedding geometry characterizes static distributions, not temporal evolution.
+This work bridges these lines by providing: (i) operational definitions of trajectory, attractor, and regime adapted to the LLM setting; (ii) a measurement protocol that maps text sequences to geometric indicators; (iii) empirical demonstration that these indicators reveal distinct, controllable dynamical behaviors.
+Importantly, the notions of trajectory, attractor, and regime used here are not inherited from classical dynamical systems theory, but redefined operationally through observable, embedding-based indicators derived from discrete linguistic transformations.
+The contribution is not the application of classical concepts to a new domain, but the construction of a measurement framework that makes trajectory-level analysis possible where it previously was not.
+3 Semantic Space and Calibrated Similarity
+This section describes the semantic measurement infrastructure used throughout this paper, following the calibration framework introduced in Tacheny (2025).
+Artifact and representation spaces.
+We define the artifact space $\mathcal{A}$ as the set of all textual outputs that can be generated by a language model.
+To enable geometric analysis, artifacts are projected into a continuous representation space $\mathcal{E}\subset\mathbb{S}^{d-1}$ via a normalized embedding function $\psi:\mathcal{A}\to\mathcal{E}$, where $\mathbb{S}^{d-1}$ is the unit hypersphere in $\mathbb{R}^{d}$.
+Each artifact $a\in\mathcal{A}$ is represented by a unit vector $e=\psi(a)\in\mathcal{E}$.
+The framework is embedding-model agnostic: any sentence encoder can instantiate $\psi$.
 Calibrated similarity.
-We also compute the calibrated similarity:
-$\tilde{s}_{i}=\tilde{s}(e_{i},e_{i}^{\prime})=f_{\text{isotonic}}(s_{i}^{(m)}),$ (14)
-where $f_{\text{isotonic}}$ denotes the isotonic regression mapping introduced in Section 4.
-For each perturbation type $k$ with $N_{k}$ pairs, we compute three stability indicators:
-$\mu_{k}=\frac{1}{N_{k}}\sum_{i\in\mathcal{I}_{k}}s_{i},\quad\sigma_{k}=\sqrt{\frac{1}{N_{k}}\sum_{i\in\mathcal{I}_{k}}(s_{i}-\mu_{k})^{2}},\quad r_{k}=\frac{1}{N_{k}}\sum_{i\in\mathcal{I}_{k}}\mathbf{1}_{\{s_{i}\geq\tau\}},$ (15)
-where $\mathcal{I}_{k}$ denotes the index set of pairs for perturbation type $k$ and $\tau$ is the stability threshold.
-Threshold selection.
-Rather than choosing arbitrary thresholds, we use the data-driven high-confidence similarity thresholds defined in Section 3.6:
-  • For base model similarity: $\tau=\tau_{\text{HCS}}^{(m)}\approx 0.72$
-  • For calibrated similarity: $\tau=\tilde{\tau}_{\text{HCS}}\approx 0.65$
-The stability rate $r_{k}$ thus measures the proportion of perturbed pairs that maintain similarity above the threshold derived from high-confidence human judgments, providing a principled criterion for local stability.
-5.4 Empirical Results
-Tables 2 and 3 report the local stability metrics for base model (raw cosine) and calibrated similarities, respectively.
-For each perturbation type, we provide the number of pairs, mean similarity, standard deviation and stability rate (proportion of pairs exceeding the corresponding threshold).
-Perturbation Type Pairs ($N_{i}$) Mean ($\mu_{\text{base},i}$) Std. Dev. ($\sigma_{\text{base},i}$) Stab. Rate ($r_{\text{base},i}$)
-DETERMINER VARIATION 100 0.932 0.017 1.00
-TENSE VARIATION 100 0.909 0.030 1.00
-SYNONYM SUBSTITUTION 100 0.853 0.060 0.95
-LOGICAL PARAPHRASE 100 0.891 0.023 1.00
-NOMINALIZATION 100 0.866 0.035 1.00
-COREFERENCE EXPANSION 100 0.882 0.022 1.00
-QUANTIFIER VARIATION 100 0.848 0.039 0.97
-All types (overall) 700 0.883 0.045 0.99
-Table 2: Local stability metrics using base model similarity $s^{(m)}$ (threshold $\tau_{\text{HCS}}^{(m)}=0.72$).
-Perturbation Type Pairs ($N_{i}$) Mean ($\mu_{\text{sim},i}$) Std. Dev. ($\sigma_{\text{sim},i}$) Stab. Rate ($r_{\text{sim},i}$)
-DETERMINER VARIATION 100 0.926 0.054 1.00
-TENSE VARIATION 100 0.868 0.069 1.00
-SYNONYM SUBSTITUTION 100 0.799 0.082 0.87
-LOGICAL PARAPHRASE 100 0.831 0.049 1.00
-NOMINALIZATION 100 0.818 0.078 1.00
-COREFERENCE EXPANSION 100 0.830 0.045 1.00
-QUANTIFIER VARIATION 100 0.798 0.050 0.97
-All types (overall) 700 0.839 0.075 0.98
-Table 3: Local stability metrics using calibrated similarity $\tilde{s}$ (threshold $\tilde{\tau}_{\text{HCS}}=0.65$).
-5.5 Discussion
-The comparison between Tables 2 and 3 reveals two critical findings regarding local stability.
-Embedding function exhibits strong local stability.
-The base model similarity analysis (Table 2) demonstrates that the embedding function $\psi$ exhibits excellent local stability, with an overall stability rate of $99\%$.
-All perturbation types maintain mean similarities well above the threshold $\tau_{\text{HCS}}^{(m)}=0.72$, ranging from $0.848$ (QUANTIFIER_VARIATION) to $0.932$ (DETERMINER_VARIATION).
-This confirms that the pretrained embedding model is robust to small linguistic perturbations, preserving high semantic similarity even under controlled transformations such as tense variation, synonym substitution and logical paraphrasing.
-Isotonic calibration introduces minimal stability degradation.
-Table 3 shows that isotonic regression calibration slightly reduces the overall stability rate from $99\%$ to $98\%$, a modest degradation of only 1 percentage point.
-This small decrease is attributable to the piecewise constant (discontinuous) nature of isotonic regression: when a perturbed pair $(a_{i},a_{i}^{\prime})$ has a raw cosine similarity $s^{(m)}$ near a discontinuity point of $f_{\text{isotonic}}$, the calibrated similarities $\tilde{s}(e_{i},e_{i})$ and $\tilde{s}(e_{i}^{\prime},e_{i}^{\prime})$ may fall into different constant segments, potentially crossing the threshold $\tilde{\tau}_{\text{HCS}}=0.65$.
-The most affected perturbation type is SYNONYM_SUBSTITUTION, whose stability rate drops from $0.95$ to $0.87$.
-This is expected: synonym substitutions produce moderate cosine similarities (mean $0.853$) that fall in a region where isotonic regression has more step discontinuities.
-In this moderate range, small linguistic perturbations can cause similarities to jump across step boundaries in $f_{\text{isotonic}}$, leading to larger changes in calibrated values than in the original cosine similarities.
-Conversely, perturbation types with very high base similarities (TENSE_VARIATION: $0.909$, DETERMINER_VARIATION: $0.932$) remain perfectly stable ($r=1.00$) after calibration, as their similarities fall in a region where isotonic regression is nearly constant and far from the threshold.
-Practical acceptability of calibrated stability.
-Despite the theoretical discontinuity of isotonic regression, the empirical stability rate of $98\%$ remains highly acceptable for practical use.
-This demonstrates that isotonic regression achieves its primary goal (eliminating bias and improving correlation with human judgments) without substantially compromising local stability.
-The 1% degradation is a negligible cost compared to the significant gains in calibration quality (ECE reduction from $0.0797$ to $0.0000$, as shown in Table 1).
-Implications for dynamic analysis.
-These results validate the use of calibrated similarity $\tilde{s}$ for studying agentic loop dynamics.
-The embedding function $\psi$ exhibits strong local stability and isotonic calibration preserves this property with only minimal degradation.
-This ensures that the semantic trajectories observed in agentic loops reflect genuine semantic evolution rather than instabilities introduced by either the embedding function or the calibration process.
-With local stability empirically established, we now turn to the formal analysis of agentic loop dynamics, confident that our similarity metric provides a stable and human-aligned foundation for measuring iterative transformations in generative systems.
+Raw cosine similarity between embeddings suffers from anisotropy: pretrained embedding spaces concentrate vectors in a narrow cone, causing similarities to cluster around high values even for semantically distant pairs Ethayarajh (2019).
+To address this, we employ the calibrated similarity function proposed in Tacheny (2025):
+$\tilde{s}(e_{1},e_{2})=f_{\mathrm{isotonic}}(\langle e_{1},e_{2}\rangle),$ (1)
+where $f_{\mathrm{isotonic}}$ is a monotonic mapping learned via isotonic regression on human-annotated sentence pairs.
+This calibration eliminates systematic bias, improves correlation with human judgments (Spearman $\rho\approx 0.86$) and preserves local stability under small linguistic perturbations ($\geq 98\%$).
+Throughout this paper, all geometric measurements use the calibrated similarity $\tilde{s}$.
+High-confidence similarity threshold.
+Following Tacheny (2025), we use the data-driven decision boundary $\tilde{\tau}_{\text{HCS}}$, the 5% quantile of calibrated similarity among pairs with high human similarity.
+This threshold guarantees that at least 95% of truly similar pairs exceed this value and is used for cluster detection in subsequent sections.
+Note that the specific value of $\tilde{\tau}_{\text{HCS}}$ depends on the embedding model; throughout this paper, we adopt the same embedding model as Tacheny (2025), yielding $\tilde{\tau}_{\text{HCS}}\approx 0.65$.
+Robustness to metric choice.
+While we use calibrated similarity for improved interpretability, the qualitative conclusions of this work—contractive vs.
+exploratory dynamics, attractor formation, regime classification—depend only on order preservation, not on the specific numeric values.
+Since isotonic calibration is a monotonic transformation, it preserves pairwise similarity rankings and nearest-neighbor relationships.
+Consequently, all definitions and empirical trends reported here would hold under raw cosine similarity or any other order-preserving metric.
+The calibration improves human interpretability of threshold values but is not required for the framework’s validity.
 4 Agentic Loop Dynamics: Theoretical Framework
 This section formalizes the theoretical framework underlying agentic loop dynamics, defining key concepts such as trajectory, cluster and attractor before turning to empirical experiments.
-6.1 Foundational Spaces and Operations
-We begin by establishing the mathematical spaces and primitive operations that underlie all agentic loop structures.
-Artifact and embedding spaces.
-Let $\mathcal{A}$ denote the artifact space, where each element $a\in\mathcal{A}$ represents a linguistic artifact (e.g., a sentence, paragraph or document).
-Let $\psi:\mathcal{A}\to\mathcal{E}$ be the embedding map introduced in Section 3, which projects artifacts into a semantic vector space where their meaning can be quantitatively compared.
-Primitive operations.
-We introduce two fundamental operations:
+4.1 Primitive Operations
+Building on the artifact space $\mathcal{A}$, representation space $\mathcal{E}$, and calibrated similarity $\tilde{s}$ defined in Section 3, we introduce two fundamental operations:
   • $\mathrm{LLM}:\mathcal{A}\to\mathcal{A}$ denotes a large language model’s generation operation, which takes a formatted input string (prompt) and produces a generated output string.
@@
 5.2.1 Global and Local Dynamics
-Figure 9 presents the evolution of both local and global geometric metrics along the trajectory of 50 iterations.
-These indicators provide a direct view of the loop’s dynamics before any segmentation into clusters.
-Figure 9: Local and global geometric evolution of the contractive loop trajectory.
-(Top) Local distances and similarities between consecutive embeddings $(e_{t},e_{t-1})$.
-(Bottom) Global distances and similarities relative to the initial embedding $e_{0}$.
 Local stability.
-The local Euclidean distance $\|e_{t}-e_{t-1}\|_{2}$ fluctuates between $0.45$ and $0.55$, while the corresponding calibrated similarity $\tilde{s}(e_{t},e_{t-1})$ remains consistently high in the range 0.820.820.950.95.
-This indicates that each rewriting step introduces only minimal changes to the representation, consistent with the model’s smooth behavior under incremental rewriting.
+The calibrated similarity $\tilde{s}(e_{t},e_{t-1})$ between consecutive embeddings remains consistently high, in the range 0.820.820.950.95 (Figure 1).
+Each rewriting step introduces only minimal semantic changes.
+Figure 1: Local similarity $\tilde{s}(e_{t},e_{t-1})$ for the contractive loop. Values remain consistently high (>0.8>0.8), indicating small semantic displacements between consecutive iterations.
 Global convergence.
-When compared to the initial state $e_{0}$, the global distance stabilizes around $0.6$ and the calibrated similarity remains near $0.75$.
-After a few initial iterations, the trajectory stops drifting and remains confined within a bounded region of $\mathcal{E}$, consistent with the expected contractive regime.
-This global stability validates the use of local geometric measures as reliable indicators of semantic convergence.
+The similarity to the initial embedding $\tilde{s}(e_{t},e_{0})$ stabilizes near $0.75$ after initial iterations (Figure 2).
+The trajectory stops drifting and remains confined within a bounded region of $\mathcal{E}$.
+Figure 2: Global similarity $\tilde{s}(e_{t},e_{0})$ for the contractive loop. The trajectory stabilizes at moderate similarity to the initial state, indicating bounded drift.
 Interpretation.
@@
 5.3.1 Global and Local Dynamics
-Figure 13 presents the evolution of both local and global geometric metrics over 50 iterations.
-Unlike the contractive regime, these indicators show strong irregularity and the absence of any stabilization phase.
-Figure 13: Local and global geometric evolution of the exploratory loop trajectory.
-(Top) Local distances and similarities between consecutive embeddings $(e_{t},e_{t-1})$.
-(Bottom) Global distances and similarities relative to the initial embedding $e_{0}$.
 Local instability.
-The local Euclidean distance $\|e_{t}-e_{t-1}\|_{2}$ remains high across iterations, oscillating between $0.8$ and $1.3$ with no decreasing trend.
-Similarly, the local similarity $\tilde{s}(e_{t},e_{t-1})$ fluctuates widely between $0.2$ and $0.6$, occasionally dropping close to zero.
-These erratic variations indicate that each step produces a major semantic reorientation rather than a refinement of the previous state.
+The local similarity $\tilde{s}(e_{t},e_{t-1})$ fluctuates widely between $0.2$ and $0.6$, occasionally dropping close to zero (Figure 6).
+Each step produces a major semantic reorientation rather than incremental refinement.
+Figure 6: Local similarity $\tilde{s}(e_{t},e_{t-1})$ for the exploratory loop. Values fluctuate erratically, indicating large semantic displacements at each iteration.
 Global divergence.
-The global distance to the initial embedding $e_{0}$ rapidly increases during the first iterations and stabilizes around $\|e_{t}-e_{0}\|_{2}\approx 1.3$ (see Figure 13, bottom left).
-The corresponding global similarity $\tilde{s}(e_{t},e_{0})$ remains very low (mostly below $0.2$), consistent with a loop that continuously generates artifacts semantically unrelated to the starting point.
-This behavior reveals the absence of any long-term memory or attractor basin in the semantic manifold.
+The similarity to the initial embedding $\tilde{s}(e_{t},e_{0})$ remains very low throughout, mostly below $0.2$ (Figure 7), indicating continuous generation of artifacts semantically unrelated to the starting point.
+Figure 7: Global similarity $\tilde{s}(e_{t},e_{0})$ for the exploratory loop. The trajectory rapidly diverges from the initial state and remains semantically distant throughout.
 Interpretation.
@@
 Summary.
-The exploratory loop displays:
-  • persistent high local distances ($\|e_{t}-e_{t-1}\|_{2}\approx 1.0$) and low similarities ($\tilde{s}<0.5$);
-  • no geometric stabilization, either locally or globally;
-  • and the complete absence of clusters or attractor basins for all tested parameters.
-In semantic terms, this loop corresponds to a continuous destruction and regeneration of meaning: each iteration rewrites the narrative space from scratch, producing an intrinsically non-convergent agentic dynamic.
+The exploratory loop displays persistently low local similarities ($\tilde{s}(e_{t},e_{t-1})<0.5$), no geometric stabilization either locally or globally, and complete absence of clusters or attractor basins for all tested parameters.
+In semantic terms, this loop corresponds to continuous destruction and regeneration of meaning: each iteration rewrites the narrative space from scratch.
 5.4 Comparative Summary
 The empirical analysis of contractive and exploratory loops reveals fundamentally distinct dynamical regimes, following the formal definitions established in Section 4.5.
-Table 9 summarizes the key quantitative and qualitative differences between these two archetypal behaviors.
+Table 6 summarizes the key differences.
 Metric Contractive Loop Exploratory Loop
-Mean local distance $\sim$0.5 $\sim$1.0
-Mean local similarity >>0.85 <<0.5
+Mean local similarity $\tilde{s}(e_{t},e_{t-1})$ >>0.85 <<0.5
+Global similarity $\tilde{s}(e_{t},e_{0})$ $\sim$0.75 <<0.2
 Clusters detected 5 $\to$ 1 0
-Global drift Bounded Unbounded
-Regime type Convergent Divergent
+Regime type Contractive Exploratory
 Table 6: Comparative summary of contractive and exploratory agentic loop dynamics.
@@
 The geometric framework established in this paper provides the measurement infrastructure required to empirically investigate these questions, enabling rigorous quantification of whether and how agentic systems can achieve geometric self-regulation.
-8.4 Application: Creativity as Objective
-The agentic geometric cognition framework introduced in Section 8.3 raises a natural application question: what if the objective that the agent targets is to maximize creativity?
-This section explores how computational creativity, formalized by Boden Boden (2009) as exploration-transformation-combination processes and by Colton and Wiggins Colton and Wiggins (2012) as novelty-value-surprise dimensions, can be operationalized as a specific objective within the geometry of agentic cognition framework.
-Creativity metrics as geometric observables.
-To target creativity as an objective, an agent with geometric self-awareness must first be able to measure creativity dimensions along its trajectory.
-Our geometric framework provides measurable indicators that correspond to Colton & Wiggins’ creativity dimensions:
-  • Novelty corresponds to exploratory dynamics, that is, trajectories that visit diverse regions of semantic space without settling into stable attractors.
-The absence of persistent clusters and low calibrated similarity between consecutive states ($\tilde{s}(e_{t},e_{t-1})$) indicate ongoing exploration that generates new ideas rather than refining existing ones.
-  • Value corresponds to convergence toward stable attractors with low dispersion.
-Trajectories that stabilize around a fixed point produce semantically coherent, trustworthy outputs.
-Conversely, divergent trajectories exhibit semantic drift, where outputs lose coherence and become unreliable, yielding low value.
-  • Surprise corresponds to sudden, unpredictable semantic transitions, specifically trajectory discontinuities where the semantic state jumps sharply after periods of stability.
-Contractive loops exhibit low surprise (smooth, predictable convergence), while exploratory loops may produce high surprise through unpredictable jumps between distant semantic regions.
-This operationalization enables measurable creativity control.
-Boden’s exploratory creativity can be triggered within agentic loop iterations by introducing high-temperature sampling and role-based divergent prompts Zhao et al.
-(2024) (for instance, adopting perspectives such as an artist, a child or a poet), which induce high dispersion and novelty generation.
-However, unconstrained exploration risks semantic drift and loss of value.
-To balance creativity, subsequent iterations can act as consolidators, employing role-based contractive prompts (e.g., scientist, engineer, critic) that reduce dispersion and restore coherence, maximizing value while preserving novel content.
-Composite loops as creativity engines.
-This dual-phase architecture naturally maps to composite loops as defined in Section 6.2. A composite loop with two phases per iteration would take the form:
-$a_{t+1}=\mathrm{LLM}_{\text{contract}}(P_{\text{contract}}(\mathrm{LLM}_{\text{explore}}(P_{\text{explore}}(a_{t})))),$ (27)
-where $P_{\text{explore}}$ uses role-based divergent prompts (e.g., "adopt the perspective of an artist") to induce exploratory divergence and $P_{\text{contract}}$ employs precision-oriented prompts (e.g., "critically evaluate and consolidate") to enforce contraction.
-The generation parameters can be tuned accordingly: $\mathrm{LLM}_{\text{explore}}$ could use high temperature (e.g., $T=1.2$) for creative diversity, while $\mathrm{LLM}_{\text{contract}}$ could use low temperature (e.g., $T=0.3$) for deterministic refinement.
-The singular loop experiments presented in Section 7 establish that these individual $(P,\mathrm{LLM})$ pairs can reliably induce their respective regimes.
-Composing them sequentially would create an oscillatory dynamic where each iteration explores the semantic manifold then contracts toward coherent outputs, effectively implementing Boden’s explore-transform-combine cycle through geometric dynamics.
-The interplay between these regimes may enable a new kind of controlled computational creativity, where the balance between exploration and consolidation is explicitly managed through prompt design, role assignment and iterative feedback.
-Future work could formalize this balance through composite loop architectures that adaptively modulate their dynamics based on creativity objectives, bridging Boden’s theoretical framework with the geometric formalism introduced here.
+6.4 Speculative Application: Creativity
+The geometric indicators developed here—drift, dispersion, attractor formation—bear a suggestive correspondence to dimensions of computational creativity as formalized by Boden Boden (2009) and Colton & Wiggins Colton and Wiggins (2012).
+Exploratory dynamics (high dispersion, no stable attractors) may relate to novelty generation, while contractive dynamics (convergence, low dispersion) may relate to value consolidation.
+This suggests a speculative direction: composite loops that alternate exploratory and contractive phases could potentially balance novelty and coherence, with prompt design and temperature modulating the regime at each phase Zhao et al.
+(2024).
+We emphasize that this remains a hypothesis for future investigation; the present work provides only the measurement infrastructure, not empirical validation of creativity-specific claims.
 6.5 Extending Toward Complex Agentic Architectures
@@
 This work introduced a geometric framework for analyzing the dynamics of agentic loops: recursive processes in which large language models iteratively transform their own outputs.
-By grounding these dynamics in a calibrated semantic embedding space (,s~)($\mathcal{E}$$\tilde${s}) that enables human-aligned measurements, we made it possible to quantify how iterative LLM processes unfold, stabilize or diverge across iterations.
+By grounding these dynamics in a calibrated semantic embedding space (,s~)($\mathcal{E}$$\tilde${s}), we made it possible to quantify how iterative LLM processes unfold, stabilize or diverge.
 Summary of contributions.
-The study provides three main contributions:
   1. 1.
-    A formal distinction between the artifact space $(\mathcal{A},F)$, where linguistic transformations occur, and its continuous representation space $(\mathcal{E},\|\cdot\|)$, enabling quantitative measurement of agentic dynamics through geometric trajectory analysis.
+    A formal theoretical framework distinguishing artifact space $(\mathcal{A},F)$ from representation space $(\mathcal{E},\|\cdot\|)$, with precise definitions of trajectories, clusters, attractors and dynamical regimes.
   2. 2.
-    A calibrated similarity metric $\tilde{s}=f_{\mathrm{isotonic}}\circ s^{(m)}$ that corrects the anisotropy of embedding spaces and provides human-aligned geometric indicators: dispersion, displacement, cluster coherence and trajectory discontinuities.
+    An incremental cluster detection algorithm that operationalizes these concepts for empirical analysis.
   3. 3.
-    An empirical demonstration that agentic loops exhibit distinct dynamical regimes (contractive or exploratory) depending solely on their prompt formulation, with measurable geometric signatures that characterize each regime.
+    Empirical demonstration that agentic loops exhibit distinct dynamical regimes (contractive or exploratory) depending on prompt formulation, with measurable geometric signatures.
 Empirical findings.
-Using a controlled experimental setup, we showed that:
-  • The contractive loop, driven by incremental rewriting, produces a smooth and bounded trajectory in $\mathcal{E}$ that converges toward a dense attractor region.
-Its dynamics exhibit local contractivity: successive embeddings remain close, dispersion decreases over time, and the trajectory stabilizes within persistent clusters.
-  • The exploratory loop, built on alternating summarization and negation, generates an unbounded trajectory with no persistent clusters or attractors.
-Its behavior is characterized by high local displacements and low inter-iteration similarity, with dispersion remaining elevated throughout the process.
-Together, these two regimes define a spectrum of agentic dynamics, from convergent stabilization to divergent exploration, each with distinct geometric signatures.
+  • The contractive loop produces a bounded trajectory converging toward a dense attractor region, with local contractivity and persistent clustering.
+  • The exploratory loop generates an unbounded trajectory with no persistent clusters, characterized by high local displacements and continuous semantic drift.
+These regimes define a spectrum from convergent stabilization to divergent exploration.
 Theoretical implications.
-From a dynamical systems perspective, these observations are consistent with a view of LLM-based transformations $F$ as mappings whose dynamics resemble those governed by an effective contraction rate:
-$\|F(a_{1})-F(a_{2})\|\leq L\,\|a_{1}-a_{2}\|,\qquad L<1\Rightarrow\text{contraction},\quad L>1\Rightarrow\text{divergence}.$ (28)
-Since $F$ operates in $\mathcal{A}$ but all geometric measurements occur in $\mathcal{E}$ through $\psi$, the contraction rate is understood with respect to the induced operator $\hat{F}=\psi\circ F$.
-While such bounds cannot be measured directly for deep generative operators, the observed trajectory patterns suggest a regime-dependent effective contraction rate that could be estimated through embedding-space analysis.
-This opens the door to a formal theory of agentic dynamics, linking prompt design, model sampling and geometric regularity.
+The observations are consistent with a view of LLM transformations as mappings whose dynamics resemble those governed by an effective contraction rate.
+While such bounds cannot be measured directly for deep generative operators, observed trajectory patterns suggest a regime-dependent effective contraction rate that could be estimated through embedding-space analysis.
 Broader outlook.
-The contrast between contractive and exploratory regimes highlights a key tension in autonomous generative systems: the need to balance semantic exploration and stability.
-While this paper focused on singular loops to establish foundational dynamical phenomena, the natural next step is the study of composite loops, structures where multiple LLM calls with different prompts compose sequentially within each iteration:
-$a_{t+1}=\mathrm{LLM}_{2}(P_{2}(\mathrm{LLM}_{1}(P_{1}(a_{t})))).$ (29)
-For example, a generate-critique-revise cycle would use $P_{1}$ for generation, $P_{2}$ for critique and $P_{3}$ for revision.
-The singular loop experiments demonstrated that individual $(P,\mathrm{LLM})$ pairs can reliably induce contractive or exploratory dynamics.
-Composing these building blocks offers a systematic path toward controlled agentic behavior: alternating exploratory (divergent) and consolidating (contractive) phases to enable agents that explore the semantic space while maintaining coherence.
-This principle extends naturally to complex agentic architectures, where multiple loops interact, synchronize or self-regulate within shared semantic manifolds.
-Hierarchical, cooperative and adversarial multi-loop systems could exhibit emergent behaviors (collective attractors, dynamic equilibria or oscillatory states) that go beyond the capabilities of individual loops.
-Understanding these phenomena requires extending the geometric framework presented here to handle compositions, feedback and synchronization across interacting semantic trajectories.
-Final remark.
-Ultimately, this work provides both a language and a methodology for describing the internal geometry of agentic loops in LLMs.
-By establishing these foundations through singular loop analysis, it opens the path toward understanding composite and complex agentic systems that balance exploration and consolidation.
-The framework connects naturally to computational creativity research, where the geometric indicators developed here (dispersion, cluster coherence, trajectory discontinuities) can operationalize theoretical notions of novelty, value and surprise.
-More broadly, it establishes a bridge between agentic AI and dynamical systems theory, encouraging future study of how autonomous generative processes evolve within semantic embedding spaces.
+The contrast between contractive and exploratory regimes highlights a key tension: the need to balance semantic exploration and stability.
+Composite loops that alternate exploratory and contractive phases offer a systematic path toward controlled agentic behavior: exploring semantic space while maintaining coherence.
+This connects naturally to computational creativity research, where the geometric indicators developed here can operationalize theoretical notions of novelty, value and surprise.
+More broadly, this work establishes a bridge between agentic AI and dynamical systems theory, encouraging future study of how autonomous generative processes evolve within semantic embedding spaces.
 Acknowledgments
@@
 Appendix A Contractive Loop: Complete Trajectory
-This appendix provides the complete sequence of artifacts generated by the contractive loop experiment described in Section 7.
-The initial artifact $a_{0}$ undergoes 50 iterations of the rewriting transformation, producing the trajectory $\{a_{t}\}_{t=0}^{50}$ shown below.
-The progressive evolution of the text demonstrates the contractive dynamics analyzed in Subsection 7.2: initial rapid refinement (iterations 0–20), followed by stabilization into a dense semantic attractor (iterations 20–50) where only minor stylistic variations occur while the core message remains fixed.
+This appendix provides the complete sequence of artifacts generated by the contractive loop experiment. The initial artifact $a_{0}$ undergoes 50 iterations of the rewriting transformation.
+The progressive evolution demonstrates contractive dynamics: initial rapid refinement (iterations 0–20), followed by stabilization into a dense semantic attractor (iterations 20–50) where only minor stylistic variations occur while the core message remains fixed.
 t=0:  Music has the power to connect people across cultures and
@@
 Appendix B Exploratory Loop: Complete Trajectory
-This appendix provides the complete sequence of artifacts generated by the exploratory loop experiment described in Section 7.
-The initial artifact $a_{0}$ undergoes 50 iterations of a summarize-then-negate transformation, producing the trajectory $\{a_{t}\}_{t=0}^{50}$ shown below.
-The trajectory demonstrates the exploratory dynamics analyzed in Subsection 7.3: persistent semantic instability with no convergence, exhibiting large inter-iteration distances and continuous oscillation between opposing philosophical positions without settling into any attractor basin.
+This appendix provides the complete sequence of artifacts generated by the exploratory loop experiment. The initial artifact $a_{0}$ undergoes 50 iterations of a summarize-then-negate transformation.
+The trajectory demonstrates exploratory dynamics: persistent semantic instability with no convergence, exhibiting large inter-iteration distances and continuous oscillation between opposing philosophical positions without settling into any attractor basin.
 t=0:  Music has the power to connect people across cultures and

### v5

@@
 divergence, attractor detection) are order-based and thus robust to any monotonic transformation of the similarity metric, including raw cosine similarity.
+Claims investigated.
+  • C1: Singular loops induced by prompt P can fall into regimes measurable in the embedding space.
+  • C2: Under a paraphrase prompt, trajectories form persistent clusters with identifiable attractors and decreasing dispersion over time.
+  • C3: Under a negation prompt, no persistent cluster emerges under the tested similarity and dispersion thresholds.
 2 Related Work
@@
 Exploring conditions that produce oscillatory dynamics, such as prompts with explicit alternating instructions or adversarial constraints, remains an avenue for future work.
+4.6 Assumptions and Scope
+The framework introduced in this work relies on a number of explicit assumptions and scope restrictions, which we state here to clarify the validity and intended interpretation of the proposed definitions and empirical observations.
+A1. Black-box transformation operator.
+We assume that the transformation operator $F:\mathcal{A}\rightarrow\mathcal{A}$, implemented via a prompt and a language model invocation, is treated as a black box.
+No assumptions are made about its internal architecture, differentiability, Lipschitz continuity, or contractive properties in the classical analytical sense.
+All characterizations of dynamics are therefore observational rather than analytical.
+A2. Finite-horizon trajectories.
+Agentic trajectories are observed over a finite number of iterations $T$.
+All notions of convergence, divergence, and stability are defined operationally with respect to finite trajectories, rather than as asymptotic guarantees.
+The framework does not claim formal convergence in the limit $T\rightarrow\infty$.
+A3. Representation-based observation.
+While agentic loops operate in the discrete artifact space $\mathcal{A}$, their dynamics are observed indirectly through a projection into a continuous representation space $\mathcal{E}$ via an embedding function $\psi$.
+We assume that relative semantic proximity in $\mathcal{E}$ reflects meaningful semantic relationships between artifacts, but make no claim that $\mathcal{E}$ constitutes the true state space of the generative process.
+A4. Order-preserving similarity measures.
+The proposed definitions of clusters, attractors, and dynamical regimes rely on relative similarity comparisons rather than absolute metric values.
+Consequently, the framework assumes only that the similarity function preserves the ordering of semantic proximity.
+While calibrated similarity is used for interpretability, the qualitative classification of regimes is invariant under any monotonic transformation of the similarity measure.
+A5. Prompt-defined operators and stochasticity.
+The effective transformation operator $F$ is defined jointly by the prompt template and the language model decoding configuration.
+Sampling stochasticity is therefore an inherent component of the dynamics.
+Observed regimes characterize the statistical behavior induced by a given prompt-model configuration, rather than deterministic properties of the model alone.
+These assumptions delineate the scope of the framework.
+The objective is not to derive formal guarantees about language model dynamics, but to provide an operational and interpretable methodology for characterizing and comparing the observable behavior of iterative agentic loops.
 4.7 Cluster Detection Algorithm
@@
 All experiments are performed locally using the Ollama inference framework (version 0.12.9) with the deepseek-r1:8b model.
-This model provides a lightweight yet expressive transformer architecture, suitable for studying fine-grained textual transformations over multiple recursive steps.
+This model was selected for three reasons: (i) it is small enough to run locally without GPU clusters, enabling full reproducibility; (ii) despite its size, it demonstrates strong instruction-following capabilities on rewriting and reasoning tasks, ensuring that prompt-induced behaviors are faithfully executed; and (iii) local execution eliminates API rate limits and guarantees deterministic infrastructure, which is essential when running 50-iteration loops repeatedly.
 We use a sampling temperature of $0.8$, which introduces mild stochasticity while preserving semantic coherence, ensuring that the observed trajectories reflect both the model’s internal tendencies and the intrinsic stability of the prompt formulation.
@@
 The process is repeated for $T=50$ iterations, generating a trajectory $\{a_{t}\}_{t=0}^{T}$ and its projection $\{e_{t}=\psi(a_{t})\}_{t=0}^{T}$ in the representation space $\mathcal{E}$.
+This horizon balances computational cost against sufficient observation length: preliminary runs indicated that contractive loops stabilize within 20–30 iterations, while exploratory loops show no sign of convergence; $T=50$ thus provides adequate margin to distinguish these regimes.
 All trajectories are embedded using the same function $\psi$ and analyzed in the calibrated similarity space defined in Section 3.
@@
 […]
+The similarity threshold $\lambda=0.8$ enforces high local coherence, admitting only consecutive pairs that are clearly semantically related.
+The dispersion radii $\rho\in\{0.1,0.2,0.3\}$ provide multi-resolution analysis: $\rho=0.1$ detects only tightly packed micro-clusters, while $\rho=0.3$ captures broader attractor basins; the increments are chosen to span interpretable fractions of the calibrated similarity scale.
+The patience parameter $\kappa=2$ tolerates brief noise (up to two consecutive violations) without prematurely terminating a cluster, while remaining strict enough to separate genuine regime transitions.
 This grid allows both fine-grained and coarse-grained analyses of the clustering structure, ensuring that attractor detection is not an artifact of a single threshold choice.
@@
   1. 1.
-    The sequence of pairwise displacements $\|e_{t}-e_{t-1}\|_{2}$ and calibrated similarities $\tilde{s}(e_{t},e_{t-1})$, quantifying instantaneous dynamics.
+    The sequence of calibrated similarities $\tilde{s}(e_{t},e_{t-1})$, quantifying instantaneous dynamics.
   2. 2.
-    The evolution of the distance $\|e_{t}-e_{0}\|_{2}$ and similarity $\tilde{s}(e_{t},e_{0})$ to the initial state, capturing long-term semantic drift or contraction.
+    The evolution of the similarity $\tilde{s}(e_{t},e_{0})$ to the initial state, capturing long-term semantic drift or contraction.
   3. 3.
@@
 We now apply the density-based clustering framework defined in Section 4.2 to characterize how the trajectory organizes into stable subregions or attractors.
-Three clustering configurations were tested, varying only the spatial radius $\rho$ while keeping $\lambda=0.8$ and $\kappa=2$ constant.
-Configuration $(\lambda,\rho,\kappa)=(0.8,0.1,2)$.
-This most restrictive configuration identifies only tightly packed regions in the embedding space. Five clusters were detected, corresponding to distinct stabilization phases during the trajectory.
-Cluster Iteration Range Dispersion Outliers
-Cluster 1 0–3 0.0648 2
-Cluster 2 6–12 0.0400 2
-Cluster 3 15–17 0.0400 2
-Cluster 4 20–27 0.0967 4
-Cluster 5 30–47 0.0800 8
-Table 1: Cluster configuration for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.1,2)$.
-Figure 3 shows the temporal progression of cluster memberships. The transition between clusters follows a smooth temporal sequence rather than abrupt shifts, reflecting progressive refinement.
-Figure 3: Cluster membership timeline for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.1,2)$.
-Outliers are marked in red.
-The sequence shows smooth temporal progression through five stabilization phases.
-Attractor similarity matrix.
-The inter-cluster similarity matrix (Table 2) reveals strong geometric coherence among all cluster centroids.
-C1 C2 C3 C4 C5
-C1 1.0000 0.9800 0.8060 0.8580 0.8500
-C2 0.9800 1.0000 0.8326 0.7956 0.7804
-C3 0.8060 0.8326 1.0000 0.7937 0.8600
-C4 0.8580 0.7956 0.7937 1.0000 0.9800
-C5 0.8500 0.7804 0.8600 0.9800 1.0000
-Table 2: Attractor Similarity Matrix for the contractive loop $(\lambda,\rho,\kappa)=(0.8,0.1,2)$.
-All off-diagonal similarities remain above $0.78$, consistent with the observation that even when the trajectory transiently visits different micro-regions, these remain tightly aligned in the semantic manifold.
-Clusters 4 and 5 exhibit the highest cohesion and persistence, suggesting convergence toward a single dominant attractor.
-Interpretation.
-The contractive loop thus exhibits a robust convergence regime characterized by:
-  • bounded local and global geometric variation;
-  • smooth temporal progression across increasingly stable clusters;
-  • and a high-density attractor region exhibiting inter-cluster similarity above $0.8$.
-From a linguistic perspective, this behavior reflects progressive syntactic smoothing and stylistic normalization: the LLM converges toward a canonical phrasing that represents an equilibrium point in the semantic manifold.
-Configuration $(\lambda,\rho,\kappa)=(0.8,0.2,2)$.
-When the clustering radius is increased to $\rho=0.2$, smaller micro-clusters merge into broader regions, revealing a simplified attractor landscape composed of only two persistent clusters.
-These clusters correspond to two major stabilization phases within the contractive trajectory: an initial transient phase and a final convergence basin.
+Three clustering configurations were tested, varying only the spatial radius $\rho\in\{0.1,0.2,0.3\}$ while keeping $\lambda=0.8$ and $\kappa=2$ constant.
+We present the analysis for $\rho=0.2$, which provides an intermediate resolution that clearly reveals the attractor structure; complete results for all configurations are provided in Appendix B.
+Cluster structure at $\rho=0.2$.
+At this resolution, the clustering algorithm identifies two persistent clusters corresponding to distinct stabilization phases: an initial transient phase (iterations 0–18) and a final convergence basin (iterations 21–50).
 Cluster Iteration Range Dispersion Outliers
@@
 Table 1: Cluster configuration for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.2,2)$.
-The temporal organization of the trajectory, shown in Figure 4, confirms a single transition between two stable basins.
-The early iterations (Cluster 1) exhibit slightly higher dispersion, while the second phase (Cluster 2) displays long persistence and near-perfect internal similarity.
+The temporal organization of the trajectory, shown in Figure 3, confirms a single transition between two stable basins around iteration 20.
+The early iterations (Cluster 1) exhibit slightly higher dispersion reflecting initial semantic exploration, while the second phase (Cluster 2) displays long persistence and decreasing internal dispersion—hallmarks of convergence toward a stable attractor.
 Figure 3: Cluster membership timeline for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.2,2)$.
@@
 Outliers (in red) correspond to transient points at cluster boundaries.
-Attractor similarity matrix.
-The inter-cluster similarity matrix in Table 4 shows that both clusters are strongly aligned in the embedding space, with a centroid similarity of $0.89$.
-C1 C2
-C1 1.0000 0.8900
-C2 0.8900 1.0000
-Table 4: Attractor Similarity Matrix for the contractive loop $(\lambda,\rho,\kappa)=(0.8,0.2,2)$.
-The high centroid similarity confirms that the two attractors occupy closely neighboring regions in $\mathcal{E}$, consistent with a smooth convergence process.
-The transition between them occurs around iteration 20 and marks the point where the rewriting process reaches its asymptotic stability.
-Interpretation.
-Compared to the previous configuration ($\rho=0.1$), this setting highlights the macro-scale structure of the trajectory: the loop first explores a transient manifold (Cluster 1) before collapsing into a single dense attractor (Cluster 2).
-The overall trajectory thus reflects a monotonic contraction process toward a stable equilibrium region, with no evidence of oscillation or divergence.
-Configuration $(\lambda,\rho,\kappa)=(0.8,0.3,2)$.
-At a larger radius $\rho=0.3$, the clustering process identifies a single dense region encompassing the entire trajectory.
-This configuration effectively merges all transient substructures into one unified attractor, consistent with the strongly contractive nature of the rewriting loop.
-Cluster Iteration Range Dispersion Outliers
-Cluster 1 0–50 0.2100 13
-Table 5: Cluster configuration for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.3,2)$.
-The corresponding temporal distribution, shown in Figure 5, confirms the absence of any segmentation: all iterations are assigned to the same cluster, with a few boundary outliers reflecting minimal fluctuations near the threshold $\rho=0.3$.
-Figure 5: Cluster membership timeline for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.3,2)$.
-All iterations belong to a single dense cluster, with minor outliers (in red) near the stability boundary.
-Interpretation.
-The detection of a single persistent cluster across all iterations confirms that the rewriting operator $F$ acts as a strong contraction mapping in (,s~)($\mathcal{E}$$\tilde${s}).
-Even under relaxed density constraints, no internal bifurcation or oscillatory pattern emerges.
-This configuration thus captures the fixed-point regime of the loop, where the generated sentences become semantically and stylistically indistinguishable over time.
-Summary of clustering outcomes.
-When the clustering radius $\rho$ increases to $0.2$ or above, micro-clusters merge into a single dense region, effectively collapsing the entire trajectory into one stable attractor.
-This collapse into a single dense attractor empirically validates the $\mathcal{R}_{\mathrm{ctr}}$ definition (Section 4.5), suggesting convergence toward a stable fixed point in the embedding space.
-No secondary structure is detected for $\rho\geq 0.2$, consistent with a globally contractive process with negligible oscillation.
-However, these observations are based on $T=50$ iterations.
-While the trajectory shows no signs of leaving the attractor within this window, we cannot formally guarantee that oscillations or cluster exits would not occur over longer horizons.
-The empirical stability observed here suggests such escapes are unlikely, but proving this rigorously would require a more advanced mathematical framework, for instance analyzing the probability of entering and remaining within a cluster as a function of the transformation operator $F$ and the embedding geometry.
+Multi-resolution robustness.
+Identical qualitative behavior is observed across all tested resolutions: at $\rho=0.1$ (finer resolution), five micro-clusters emerge that merge smoothly over time; at $\rho=0.3$ (coarser resolution), a single cluster encompasses the entire trajectory.
+In all cases, clusters exhibit temporal persistence, high inter-cluster similarity (>0.78>0.78), and monotonically decreasing dispersion.
+See Appendix B for complete tables, figures, and similarity matrices.
+Summary.
+The cluster analysis confirms that the contractive loop satisfies the empirical criteria for $\mathcal{R}_{\mathrm{ctr}}$ (Section 4.5):
+  • persistent clusters with identifiable attractors;
+  • decreasing dispersion indicating convergence;
+  • no oscillation or divergence across 50 iterations.
+The consistency of these findings across three clustering resolutions—without parameter tuning—strengthens the conclusion that the observed convergence reflects genuine geometric properties of the trajectory rather than artifacts of a particular configuration.
+However, these observations are based on $T=50$ iterations; while the trajectory shows no signs of leaving the attractor within this window, proving long-term stability rigorously would require analyzing the probability of entering and remaining within a cluster as a function of the transformation operator $F$ and the embedding geometry.
 5.3 Exploratory Loop Dynamics
@@
 The resulting trajectory exhibits high variability and recurrent semantic inversions, typical of a divergent or oscillatory dynamic.
-The complete sequence of generated artifacts is provided in Appendix B for reference.
+The complete sequence of generated artifacts is provided in Appendix C for reference.
 5.3.1 Global and Local Dynamics
@@
 In contrast, the exploratory loop exhibits persistent instability, with each iteration producing semantically distant outputs ($\sim$1.0 Euclidean distance) and no detectable cluster structure.
+These findings substantiate Claims C1–C3.
 This dichotomy confirms that the choice of prompt and transformation operator fundamentally determines the long-term behavior of agentic trajectories, ranging from stable convergence to unbounded semantic wandering.
@@
 These two archetypal behaviors constitute the endpoints of a broader spectrum of agentic dynamics that could potentially be modulated, combined or even harnessed for creative control.
+6.1 Limitations
+The present study has several limitations that scope its conclusions:
+  • All experiments use a single LLM (deepseek-r1:8b) with fixed decoding settings (temperature $0.8$); different architectures or sampling strategies may yield different dynamical signatures.
+  • Trajectories are observed through a single embedding function $\psi$ and calibrated similarity $\tilde{s}$; alternative embeddings or metrics could reveal different geometric structures.
+  • The observation horizon is limited to $T=50$ iterations, which may not capture slower convergence phenomena or late-emerging instabilities.
+  • The two prompts were chosen as archetypal extremes (pure paraphrase, pure negation) to cleanly separate regimes; they do not constitute an exhaustive map of the space of possible dynamical behaviors.
 6.2 Model Dependence and Generalization
@@
 6.4 Toward Geometry of Agentic Cognition
-The trajectory analysis framework developed in this paper enables a natural question: can an agentic loop reason about its own geometric dynamics and adapt its behavior to achieve specific objectives?
-This motivates a future research direction we term Geometry of Agentic Cognition: systems where the agent possesses geometric self-awareness and can metacognitively control its trajectory evolution.
-Current state: external control.
-Presently, trajectory control remains external to the loop itself. The intensity of contraction or exploration can be modulated through:
-  • Prompt engineering: adjusting the degree of rewriting, abstraction or negation in the instruction layer;
-  • Sampling parameters: manipulating temperature, nucleus probability (top-p), top-k or repetition penalties to alter stochasticity.
-These mechanisms require external intervention: a human engineer or outer control system must specify parameters a priori.
-Future direction: agentic geometric cognition.
-A fundamentally different paradigm would enable the agentic loop itself to observe, reason about and adapt its trajectory dynamics based on geometric awareness.
-In this vision, at each iteration $t$, the agent would:
-  1. 1.
-    Observe trajectory geometry: Compute calibrated similarities $\tilde{s}(e_{t},e_{t-1})$ (local dynamics), $\tilde{s}(e_{t},e_{0})$ (global drift) and trajectory dispersion to assess its current dynamical regime.
-  2. 2.
-    Reason about objective alignment: Evaluate whether its trajectory is progressing toward the desired objective, whether convergence toward a solution (requires contraction), exploration of alternatives (requires divergence) or escape from a suboptimal attractor (requires controlled instability).
-  3. 3.
-    Adapt iteration strategy: Based on this geometric self-assessment, metacognitively adjust the next transformation by increasing exploration if trapped in local minima, inducing contraction if approaching the target region or maintaining the current regime if progressing optimally.
-This would transform agentic loops from passive trajectory executors into geometric navigators: systems that understand their position in semantic space and actively steer toward objectives through informed dynamical control.
-Enabling mechanisms.
-Implementing agentic geometric cognition requires:
-  • Trajectory introspection: Embedding the agent’s own outputs $\{a_{0},a_{1},\ldots,a_{t}\}$ and computing geometric indicators (similarity, dispersion, clustering metrics) as part of the loop context.
-  • Geometric prompt characterization: Pre-characterizing prompts by their dynamical properties (contractive, exploratory) and selecting from a prompt library based on desired trajectory behavior.
-This requires a mathematical framework capable of predicting a prompt’s dynamical regime a priori, which we address in forthcoming work.
-  • Adaptive parameter control: Allowing the agent to specify its own generation parameters or select from predefined strategies (e.g., "use high-temperature exploration" vs.
-"use low-temperature refinement") based on trajectory analysis.
-This research direction opens fundamental questions at the intersection of dynamical systems, metacognition and autonomous reasoning: Can language models learn to recognize geometric patterns in their own trajectories?
-Can they develop heuristics for when to contract versus explore based on objective alignment?
-Can they generalize trajectory control strategies across different tasks?
-The geometric framework established in this paper provides the measurement infrastructure required to empirically investigate these questions, enabling rigorous quantification of whether and how agentic systems can achieve geometric self-regulation.
+The trajectory analysis framework developed in this paper enables a natural question: can an agentic loop reason about its own geometric dynamics and adapt its behavior accordingly?
+Currently, trajectory control remains external—modulated through prompt engineering or sampling parameters (temperature, top-p) specified a priori by a human engineer.
+A fundamentally different paradigm, which we term Geometry of Agentic Cognition, would enable the loop itself to observe its trajectory geometry (local similarity, global drift, dispersion), reason about whether its current regime aligns with the objective (convergence toward a solution vs.
+exploration of alternatives), and adapt its strategy accordingly.
+This would require trajectory introspection (embedding recent outputs and computing geometric indicators within the loop context) and geometric prompt characterization (pre-classifying prompts by their dynamical properties to enable informed selection).
+Such systems would transform from passive trajectory executors into geometric navigators that actively steer through semantic space.
+The framework established here provides the measurement infrastructure for investigating whether agentic systems can achieve such geometric self-regulation.
 6.5 Speculative Application: Creativity
@@
       geographical boundaries.
-Appendix B Exploratory Loop: Complete Trajectory
+Appendix B Robustness of Contractive Dynamics to Clustering Parameters
+This appendix provides complete clustering results for the contractive loop across all tested configurations, demonstrating that the qualitative conclusions of Section 5.2 are robust to the choice of clustering resolution $\rho$.
+Configuration $(\lambda,\rho,\kappa)=(0.8,0.1,2)$
+This most restrictive configuration identifies only tightly packed regions in the embedding space. Five clusters were detected, corresponding to distinct stabilization phases during the trajectory.
+Cluster Iteration Range Dispersion Outliers
+Cluster 1 0–3 0.0648 2
+Cluster 2 6–12 0.0400 2
+Cluster 3 15–17 0.0400 2
+Cluster 4 20–27 0.0967 4
+Cluster 5 30–47 0.0800 8
+Table 3: Cluster configuration for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.1,2)$.
+Figure 6: Cluster membership timeline for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.1,2)$.
+Outliers are marked in red.
+The sequence shows smooth temporal progression through five stabilization phases.
+The inter-cluster similarity matrix (Table 4) reveals strong geometric coherence among all cluster centroids, with all off-diagonal similarities above $0.78$.
+C1 C2 C3 C4 C5
+C1 1.0000 0.9800 0.8060 0.8580 0.8500
+C2 0.9800 1.0000 0.8326 0.7956 0.7804
+C3 0.8060 0.8326 1.0000 0.7937 0.8600
+C4 0.8580 0.7956 0.7937 1.0000 0.9800
+C5 0.8500 0.7804 0.8600 0.9800 1.0000
+Table 4: Attractor Similarity Matrix for the contractive loop $(\lambda,\rho,\kappa)=(0.8,0.1,2)$.
+Configuration $(\lambda,\rho,\kappa)=(0.8,0.2,2)$
+The inter-cluster similarity matrix for the intermediate resolution (main text, Table 1) shows that both clusters are strongly aligned in the embedding space, with a centroid similarity of $0.89$.
+C1 C2
+C1 1.0000 0.8900
+C2 0.8900 1.0000
+Table 5: Attractor Similarity Matrix for the contractive loop $(\lambda,\rho,\kappa)=(0.8,0.2,2)$.
+Configuration $(\lambda,\rho,\kappa)=(0.8,0.3,2)$
+At a larger radius $\rho=0.3$, the clustering process identifies a single dense region encompassing the entire trajectory, effectively merging all transient substructures into one unified attractor.
+Cluster Iteration Range Dispersion Outliers
+Cluster 1 0–50 0.2100 13
+Table 6: Cluster configuration for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.3,2)$.
+Figure 7: Cluster membership timeline for the contractive loop under $(\lambda,\rho,\kappa)=(0.8,0.3,2)$.
+All iterations belong to a single dense cluster, with minor outliers (in red) near the stability boundary.
+Summary Across Configurations
+Table 7 consolidates the clustering outcomes across all three resolutions.
+Resolution $\rho$ Clusters Min Similarity Interpretation
+0.1 (fine) 5 0.78 Micro-phases of refinement
+0.2 (medium) 2 0.89 Transient + convergence basin
+0.3 (coarse) 1 Single unified attractor
+Table 7: Summary of clustering outcomes for the contractive loop across three resolution levels.
+All configurations reveal the same qualitative pattern: temporally persistent clusters with high inter-cluster similarity and no oscillation or divergence.
+The number of detected clusters decreases monotonically with increasing $\rho$, as expected from hierarchical structure in the embedding space.
+This consistency across resolutions confirms that the observed convergence is a genuine property of the trajectory rather than an artifact of parameter choice.
+Appendix C Exploratory Loop: Complete Trajectory
 This appendix provides the complete sequence of artifacts generated by the exploratory loop experiment. The initial artifact $a_{0}$ undergoes 50 iterations of a summarize-then-negate transformation.